### v1

# Imbalanced Data Clustering using Equilibrium K-Means

Abstract
Imbalanced data, characterized by an unequal distribution of data points across different clusters, poses a challenge for traditional hard and fuzzy clustering algorithms, such as hard K-means (HKM, or Lloyd’s algorithm) and fuzzy K-means (FKM, or Bezdek’s algorithm). This paper introduces equilibrium K-means (EKM), a novel and simple K-means-type algorithm that alternates between just two steps, yielding significantly improved clustering results for imbalanced data by reducing the tendency of centroids to crowd together in the center of large clusters. We also present a unifying perspective for HKM, FKM, and EKM, showing they are essentially gradient descent algorithms with an explicit relationship to Newton’s method. EKM has the same time and space complexity as FKM but offers a clearer physical meaning for its membership definition. We illustrate the performance of EKM on two synthetic and ten real datasets, comparing it to various clustering algorithms, including HKM, FKM, maximum-entropy fuzzy clustering, two FKM variations designed for imbalanced data, and the Gaussian mixture model. The results demonstrate that EKM performs competitively on balanced data while significantly outperforming other techniques on imbalanced data. For high-dimensional data clustering, we demonstrate that a more discriminative representation can be obtained by mapping high-dimensional data via deep neural networks into a low-dimensional, EKM-friendly space. Deep clustering with EKM improves clustering accuracy by 35% on an imbalanced dataset derived from MNIST compared to deep clustering based on HKM.
Index Terms:
K-means, fuzzy K-means, equilibrium K-means, imbalanced clustering, deep clustering
I Introduction
Imbalanced data refers to data that is unevenly distributed across different clusters, which is prevalent in medical diagnosis, fraud detection, and anomaly detection. Learning from imbalanced data has been considered a challenge and a valuable direction [1]. While there is a considerable amount of research on supervised learning from imbalanced data, e.g., [2, 3, 4], unsupervised learning has not been as thoroughly explored, due to the unknown imbalance ratio between clusters making the task more challenging [5].
Clustering is an important unsupervised learning task involving grouping data into clusters based on similarity. K-means (KM) is the most popular clustering technique, valued for its simplicity, scalability, and effectiveness with real datasets. It is also used as an initialization method for more advanced clustering techniques, such as the Gaussian mixture model [6, 7]. KM starts with an initial set of centroids (cluster centers) and iteratively refines them to increase cluster compactness. The hard KM (HKM, or Lloyd’s algorithm) [8, 9] and fuzzy KM (FKM, or Bezdek’s algorithm) [10] are the two most representative KM algorithms. HKM assigns a data point to only one cluster, while FKM allows a data point to belong to multiple clusters with varying degrees of membership. In recent, they are also widely employed in deep learning, jointly trained with deep neural networks (DNNs) for clustering high-dimensional data, such as images [11, 12, 13, 14].
Many FKM variations have been developed to enhance its performance in different aspects. For example, the possibilistic FKM (PFKM) [15] was proposed to address the sensitivity of FKM to noise and outliers. FKM-$\sigma$ [16] was proposed to improve FKM’s performance on data points with uneven densities or non-spherical shapes in individual clusters. Fuzzy local information K-means (FLICM) [17] was designed to promote FKM’s performance in image segmentation. For more information on other FKM variants, please refer to the review paper [18].
Figure 1: Clustering results of a highly imbalanced dataset. (a) Ground truth. The colors represent the labels of the data points. (b) Clustering by hard K-means. (c) Clustering by fuzzy K-means. (d) Clustering by the proposed equilibrium K-means. They are all two-step alternating algorithms that iteratively update centroids and clusters.
Clustering imbalanced data also poses a challenge for both HKM and FKM. This is due to an undesirable property known as the “uniform effect”, which causes centroids to crowd together in large clusters, resulting in clusters of equal size even when input data has varied cluster sizes. An example illustrating the uniform effect of HKM and FKM is given in Fig. 1.
I-A Existing Efforts to Overcome Uniform Effect
To improve the poor performance of FKM on imbalanced data, Noordam et al. [19] developed a modified FKM called cluster-size insensitive FKM (csiFKM). This method assigns weight to adjust the contribution of data points according to the size of the clusters they belong to. Lin et al. [20] proposed a size-insensitive integrity-based FKM (siibFKM) to reduce the sensitivity of csiFKM to the distance between adjacent clusters by increasing the cluster’s integrity. However, weighting based on cluster size inadvertently increases the influence of outliers, making these algorithms sensitive to noise [18].
Another approach to tackle the challenge of imbalanced data clustering is to use a multiprototype mechanism. This method first groups data into multiple subclusters, ensuring that the data points in each subcluster are relatively balanced. The final clusters are obtained by merging adjacent subclusters. Liang et al. [21] introduced a multiprototype clustering algorithm that employs FKM to generate subclusters. Later, Lu et al. [5] proposed a self-adaptive multiprototype clustering algorithm that can automatically adjust the number of subclusters. However, multiprototype clustering algorithms involve a complex process and high complexity, with a time complexity of $O(N^{2})$, where $N$ is the number of data points in the dataset, making them computationally expensive for large datasets. We should additionally mention that Zeng et al. [22] recently proposed a soft multiprototype clustering algorithm with time complexity linear to $N$. However, the clustering process of this algorithm remains complex and is aimed at clustering high-dimensional and complex-structured data rather than imbalanced data.
I-B Our Contributions
In this paper, we propose equilibrium K-means (EKM) for imbalanced data clustering, which avoids the limitations of the aforementioned techniques. Our contributions can be summarized in three main aspects.
First, we offer a new framework that unifies HKM, FKM, and EKM with three key advantages. 1. This framework can directly re-derive HKM and FKM from unconstrained optimization problems with differentiable objectives without applying Lagrangian multipliers, simplifying the previous derivation from constrained optimization problems and non-differentiable objectives. 2. The new objective functions of HKM and FKM can be compared in the same plot, which used to be unfeasible due to the old objective function of FKM having more optimization variables than that of HKM. Such comparison can facilitate a more comprehensive and fundamental view than existing comparisons (such as [23, 24]) based on clustering outcomes. 3. This new framework allows for the uniform study of KM-type algorithms, such as the convergence conditions provided in this paper. We also reveal an explicit relationship between these KM algorithms and Newton’s method.
Second, we present EKM, derived from the new framework. Similar to FKM, EKM is a two-step alternating algorithm between updating the weighted centroid and calculating the weight, but the membership defined in EKM has a clearer physical meaning than that defined in FKM. Besides, EKM prevents centroids crowd together by introducing appropriate repulsive forces between centroids, making it more effective than HKM and FKM on imbalanced data. In addition to having the same time and space complexity as FKM, EKM also supports batch learning, making it scalable for large datasets. We demonstrate EKM’s performance through numerical experiments on two synthetic and ten real datasets from various domains. Because of space limitations, quantified metrics of performance are presented in the supplemental material, but the scatter graphs of partial clustering outcomes that we present in the paper are sufficient to demonstrate the superiority of EKM.
Finally, we investigate the cooperation between EKM and DNNs for high-dimensional data clustering. We find that mapping high-dimensional data via DNNs to a low-dimensional EKM-friendly space results in a more discriminative representation than mapping to an HKM-friendly space. When applied to an imbalanced dataset derived from MNIST, joint learning of DNNs and EKM improves clustering accuracy by 35% compared to joint learning of DNNs and HKM.
I-C Organization
We introduce the HKM, FKM, and a variation of FKM, the maximum-entropy fuzzy clustering (MEFC) in Section II. In Section III and Section IV, we introduce the proposed clustering framework and the proposed EKM. We demonstrate the performance of EKM on classic clustering tasks in Section V and deep clustering in Section VI. Finally, we conclude in Section VII.
II K-Means and Its Variations
II-A The Hard K-Means Algorithm (Lloyd’s Algorithm)
KM clustering aims to partition $N$ data points into $K$ clusters, minimizing the sum of the variances within each cluster. Mathematically, the objective can be expressed as:
$\underset{\mathbb{S}_{1},\cdots,\mathbb{S}_{K}}{\arg\min}\sum_{k=1}^{K}\sum_{\mathbf{x}\in\mathbb{S}_{k}}\|\mathbf{x}-\bm{\mu}_{k}\|_{2}^{2},$ (1)
where $\mathbb{S}_{k}$ represents the set of data points in the $k$-th cluster, $\mathbf{x}\in\mathbb{S}_{k}$ denotes a data point belonging to $\mathbb{S}_{k}$, $\bm{\mu}_{k}$ is the centroid of points in $\mathbb{S}_{k}$, expressed by
$\bm{\mu}_{k}=\frac{1}{|\mathbb{S}_{k}|}\sum_{\mathbf{x}\in\mathbb{S}_{k}}\mathbf{x},$ (2)
$|\mathbb{S}_{k}|$ signifies the size of $\mathbb{S}_{k}$, and $\|\cdot\|_{2}$ is the $l_{2}$ norm. The optimization problem (1) is NP-hard [25], and Lloyd‘s algorithm is a popular heuristic algorithm for solving this problem. Given $K$ points ($\mathbf{c}_{1}^{(1)},\cdots,\mathbf{c}_{K}^{(1)})$ as initial centroids, Lloyd’s algorithm alternates between two steps:
  1. 1.
    Assign each data point to the nearest centroid, forming $K$ clusters:
[…]
  2. 2.
    Recalculate the centroid of each cluster by taking the mean of all data points assigned to it:
    $\mathbf{c}_{k}^{(\tau+1)}=\frac{1}{|\mathbb{S}_{k}^{(\tau)}|}\sum_{\mathbf{x}\in\mathbb{S}_{k}^{(\tau)}}\mathbf{x}.$ (4)
Lloyd’s algorithm converges when the assignment of data points to clusters ceases to change, or when a maximum number of iterations is reached. The time complexity of one iteration of the above two steps is $O(NK)$. The HKM algorithm mentioned in this paper refers to Lloyd’s algorithm.
II-B The Uniform Effect of K-Means
The uniform effect refers to the propensity to generate clusters of equal size, which is implicitly implied in the objective (1) of KM. For simplicity, kindly consider a case with two clusters (i.e., $K=2$). Minimizing the objective function (1) is equivalent to maximizing the following objective function [26]:
$\max_{\mathbb{S}_{1},\,\mathbb{S}_{2}}N_{1}N_{2}\|\bm{\mu}_{1}-\bm{\mu}_{2}\|_{2}^{2},$ (5)
where $N_{1}$ and $N_{2}$ denote the sizes of $\mathbb{S}_{1}$ and $\mathbb{S}_{2}$, respectively. By isolating the effect of $\|\bm{\mu}_{1}-\bm{\mu}_{2}\|_{2}^{2}$, maximizing the above objective leads to $N_{1}=N_{2}=N/2$, indicating KM tends to produce equally sized clusters.
II-C The Fuzzy K-Means Algorithm (Bezdek’s Algorithm)
FKM attempts to minimize the sum of weighted distances between data points and centroids, with the following objective and constraints [10]:
$\min_{\mathbf{c}_{1},\cdots,\mathbf{c}_{K},\{u_{kn}\}}\quad$ $\sum_{n=1}^{N}\sum_{k=1}^{K}(u_{kn})^{m}\|\mathbf{x}_{n}-\mathbf{c}_{k}\|_{2}^{2}$ (6)
[…]
$\sum_{k}u_{kn}=1\,\forall n,$
$0<\sum_{n}u_{kn}<N\,\forall k,$
where $\mathbf{x}_{n}$ represents the $n$-th data point, $\mathbf{c}_{k}$ denotes the $k$-th centroid, $u_{kn}$ is a coefficient called membership that indicates the degree of $\mathbf{x}_{n}$ belonging to the $k$-th cluster, and $m\in(1,+\infty)$ is a hyperparameter controlling the degree of fuzziness level. Similar to the HKM algorithm, the FKM algorithm operates by alternating two steps:
  1. 1.
    Calculate the membership values of data points of being in the cluster:
    $u_{kn}^{(\tau)}=\frac{1}{\sum_{i=1}^{K}\bigg{(}\frac{\|\mathbf{x}_{n}-\mathbf{c}_{k}^{(\tau)}\|_{2}}{\|\mathbf{x}_{n}-\mathbf{c}_{i}^{(\tau)}\|_{2}}\bigg{)}^{\frac{2}{m-1}}}.$ (7)
  2. 2.
    Recalculate the centroid of each cluster by taking the mean of all data points, weighted by their membership to the cluster:
    $\mathbf{c}_{k}^{(\tau+1)}=\frac{\sum_{n}(u_{kn}^{(\tau)})^{m}\mathbf{x}_{n}}{\sum_{n}(u_{kn}^{(\tau)})^{m}}.$ (8)
The time complexity of one iteration of the above two steps is $O(NK^{2})$. The higher time complexity of FKM than that of HKM is due to the extra calculation of membership. A practical time consumption comparison can be found in [27].
II-D Maximum-Entropy Fuzzy Clustering
Karayiannis [28] added an entropy term to the objective function of FKM, resulting in MEFC. The new objective function is given as follows:
$\min_{\mathbf{c}_{1},\cdots,\mathbf{c}_{K},\{u_{kn}\}}\quad$ $\eta\sum_{n=1}^{N}\sum_{k=1}^{K}u_{kn}\ln{u_{kn}}$ (9)
$+(1-\eta)\frac{1}{N}\sum_{n=1}^{N}\sum_{k=1}^{K}u_{kn}\|\mathbf{x}_{n}-\mathbf{c}_{k}\|_{2}^{2}$
[…]
$\sum_{k}u_{kn}=1\,\forall n,$
where $\eta\in(0,1)$ is a hyperparameter, controlling the transition from maximization of the entropy to the minimization of centroid-data distances. MEFC is similar to FKM but with a different membership:
$u^{(\tau)}_{kn}=\frac{\exp{(-\lambda\|\mathbf{x}_{n}-\mathbf{c}_{k}\|_{2}^{2})}}{\sum_{i=1}^{K}\exp{(-\lambda\|\mathbf{x}_{n}-\mathbf{c}_{i}\|_{2}^{2})}},$ (10)
where $\lambda=\frac{1}{N}\frac{1-\eta}{\eta}$. The time complexity of MEFC is the same as FKM, which is $O(NK^{2})$ for one iteration.
III Smooth K-Means - A Unified Framework
III-A Objective of Smooth K-Means
Here, we propose a novel framework called smooth K-means (SKM) and demonstrate that the three KM algorithms introduced in the previous section are special cases of SKM. Based on this framework, we further develop a new clustering algorithm, aiming to improve the degraded performance of KM algorithms on imbalanced data. Denote the squared Euclidean distance [^1] Other distance metrics, such as the absolute difference and the angle between points, can also be used to define $d_{kn}$. Considering the derivation process remains the same, we use the Euclidean distance in the paper for simplicity. between the $k$-th centroid and the $n$-th data point as
$d_{kn}=\frac{1}{2}\|\mathbf{x}_{n}-\mathbf{c}_{k}\|_{2}^{2},$ (11)
and define the within-cluster sum of squares (WCSS) as the sum of squared Euclidean distances between data points and their nearest centroid, i.e.,
$\text{WCSS}:=\sum_{n=1}^{N}\min(d_{1n},\cdots,d_{Kn}).$ (12)
The goal of SKM is to find $K$ centroids that minimize an approximated WCSS, resulting in the following optimization problem:
$\underset{\mathbf{c}_{1},\cdots,\mathbf{c}_{K}}{\arg}\min\sum_{n=1}^{N}h(d_{1n},\cdots,d_{Kn}),$ (13)
where $h(d_{1n},\cdots,d_{Kn})$ is a smooth approximation to $\min(d_{1n},\cdots,d_{Kn})$ and is referred to as the smooth minimum function. Below we explain the significance of replacing the minimum function with a smooth minimum function from a statistical point of view.
III-B Statistical Interpretation of Smooth K-Means
Statistically, the centroids obtained by minimizing WCSS are maximum likelihood estimators (MLEs) of parameters in a specific type of Gaussian mixture model (GMM) referred to as “hard” GMM. Assuming that the dataset $(\mathbf{x}_{1},\cdots,\mathbf{x}_{N})$ is generated by $K$ independent multivariate normal distributions, we denote the mean vector and covariance matrix of the $k$-th normal distribution as $\bm{\mu}_{k}$ and $\mathbf{\Sigma}_{k}$, respectively. The standard GMM assumes that data points are generated from multiple normal distributions based on a certain probability distribution, with the following likelihood
$L(\bm{\mu},\mathbf{\Sigma},p\,|\,\mathbf{x})\propto$ $\prod_{n=1}^{N}\sum_{k=1}^{K}p_{kn}\det(\mathbf{\Sigma}_{k})^{-1/2}$ (14)
$\exp{\bigl{(}-\frac{1}{2}(\mathbf{x}_{n}-\bm{\mu}_{k})^{\mathsf{T}}\mathbf{\Sigma}_{k}^{-1}(\mathbf{x}_{n}-\bm{\mu}_{k})\bigl{)}},$
where $p_{kn}$ is the probability that the $n$-th data point is generated from the $k$-th normal distribution. The hard GMM, on the other hand, assumes that data points can only be generated from a single normal distribution. Accordingly, for all $n$, $p_{kn}=1$ for one $k\in\{1,\cdots,N\}$ and $p_{in}=0$ if $i\neq k$. We further assume that all normal distributions have the same covariance matrix, specifically an identity matrix. Under these assumptions, the MLEs of $\{p_{kn}\}$ are $\hat{p}_{kn}=1$ if $k=\arg\min_{i\in\{1,\cdots,K\}}\|\mathbf{x}_{n}-\bm{\mu}_{i}\|_{2}^{2}$, and $\hat{p}_{kn}=0$ otherwise. By substituting $\mathbf{\Sigma}_{k}=\mathbf{I}$ and $\hat{p}_{kn}$ into (14) and taking the logarithmic value, we obtain the log-likelihood:
$l(\bm{\mu}\,|\,\mathbf{x})\propto-\sum_{n=1}^{N}\min(\|\mathbf{x}_{n}-\bm{\mu}_{1}\|_{2}^{2},\cdots,\|\mathbf{x}_{n}-\bm{\mu}_{K}\|_{2}^{2}).$ (15)
Now it is clear that the MLEs of $\{\bm{\mu}_{1},\cdots,\bm{\mu}_{K}\}$ are the centroids minimizing WCSS.
While the hard GMM model is more computationally efficient than the standard GMM, it only considers the impact of a data point on its closest centroid. However, as seen in the success of FKM, it can be advantageous to consider the effect of data points on all centroids for certain real datasets. This can be accomplished by SKM in a parameter-free manner, which applies a smooth function to WCSS. Different smoothing methods result in unique impact patterns, allowing for the capture of various data structures. In the following sections, we will explore three popular smooth minimum functions and their impact on clustering.
III-C Three Common Smooth Minimum Functions
Assume a monotonically increasing and differentiable function $f:[0,\,+\infty)\mapsto[0,\,+\infty)$ satisfies
$\lim_{x\to+\infty}\frac{x}{f(x)}\to 0,$ (16)
or equivalently
$\lim_{x\to+\infty}\frac{1}{f^{\prime}(x)}\to 0,$ (17)
where $f^{\prime}$ is the first derivative of $f$. Let $g(x)=1/f(x)$, a smooth minimum function $h_{1}:[0,\,+\infty)^{K}\mapsto[0,+\infty)$ can be constructed by
$h_{1}(x_{1},\cdots,x_{K})=g^{-1}(g(x_{1})+\cdots+g(x_{K}))$ (18)
Let $f(x)$ be defined as $e^{\lambda x}$, then the function $h_{1}$ takes on a specific form known as LogSumExp:
$h_{1}(x_{1},\cdots,x_{K})$ $=\text{LSE}_{\lambda}(x_{1},\cdots,x_{K})$ (19)
$=-\frac{1}{\lambda}\ln(e^{-\lambda x_{1}}+\cdots+e^{-\lambda x_{K}}),$
where $\lambda$ serves as a parameter that controls the degree of approximation, with $\text{LSE}_{\lambda}\to\min$ as $\lambda\to+\infty$. Another common smooth minimum function, called p-Norm, can be defined as $f(x)=x^{p}$, which has the following specific form
$h_{1}(x_{1},\cdots,x_{K})$ $=\text{PN}_{p}(x_{1},\cdots,x_{K})$ (20)
$=(x_{1}^{-p}+\cdots+x_{K}^{-p})^{-1/p},$
and converges to $\min(x_{1},\cdots,x_{K})$ as $p\to+\infty$.
Smooth minimum functions can also be constructed using
$h_{2}(x_{1},\cdots,x_{K})=\frac{x_{1}g(x_{1})+\cdots+x_{K}g(x_{K})}{g(x_{1})+\cdots+g(x_{K})}.$ (21)
A specific example is the Boltzmann operator, which is defined as $g(x)=e^{-\alpha x}$. The Boltzmann operator takes on the form of
$h_{2}(x_{1},\cdots,x_{K})$ $=\text{boltz}_{\alpha}(x_{1},\cdots,x_{K})$ (22)
$=\frac{\sum_{i=1}^{K}x_{i}e^{-\alpha x_{i}}}{\sum_{i=1}^{K}e^{-\alpha x_{i}}},$
and converges to the minimum function as $\alpha\to+\infty$.
III-D Clustering by Minimizing Smoothed WCSS
III-D1 The relationship between LogSumExp and Maximum-Entropy Fuzzy Clustering
Approximating WCSS (12) by LogSumExp (19), we have the following objective:
$\min_{\mathbf{c}_{1},\cdots,\mathbf{c}_{K}}J_{\text{LSE}}(\mathbf{c}_{1},\cdots,\mathbf{c}_{K})=\sum_{n=1}^{N}-\frac{1}{\lambda}\ln{(\sum_{k=1}^{K}e^{-\lambda d_{kn}})}.$ (23)
The differentiable objective function $J_{\text{LSE}}$ has the first-order partial derivative:
$\partial_{\mathbf{c}_{k}}{J_{\text{LSE}}}=\frac{\partial{J_{\text{LSE}}}}{\partial{\mathbf{c}_{k}}}$ $=\sum_{n=1}^{N}\frac{e^{-\lambda d_{kn}}}{\sum_{i=1}^{K}e^{-\lambda d_{in}}}\frac{\partial{d_{kn}}}{\partial{\mathbf{c}_{k}}}$ (24)
$=-\sum_{n=1}^{N}\frac{e^{-\lambda d_{kn}}}{\sum_{i=1}^{K}e^{-\lambda d_{in}}}(\mathbf{x}_{n}-\mathbf{c}_{k}).$
The minimizer of $J_{\text{LSE}}$ can be found by gradient descent iteration:
$\mathbf{c}_{k}^{(\tau+1)}=\mathbf{c}_{k}^{(\tau)}-\gamma^{(\tau)}_{k}\partial_{\mathbf{c}_{k}}{J_{\text{LSE}}}(\mathbf{c}_{k}^{(\tau)}),$ (25)
where $\gamma^{(\tau)}_{k}$ is the learning rate at the $\tau$-th iteration. This updating procedure (25) is equivalent to the MEFC algorithm if one set
$\gamma^{(\tau)}_{k}=1/\sum_{n=1}^{N}\bigg{(}\frac{e^{-\lambda d_{kn}^{(\tau)}}}{\sum_{i=1}^{K}e^{-\lambda d_{in}^{(\tau)}}}\bigg{)}.$ (26)
Such learning rate value is related to the second-order partial derivative of $J_{\text{LSE}}$, which we will discuss later.
III-D2 Towards Lloyd’s algorithm
In the limit $\lambda\to\infty$, $\frac{e^{-\lambda d_{kn}}}{\sum_{i}e^{-\lambda d_{in}}}\to 1$ if $d_{kn}\leq d_{in}\,\forall i\in\{1,\cdots,K\}$, and $\frac{e^{-\lambda d_{kn}}}{\sum_{i}e^{-\lambda d_{in}}}\to 0$ otherwise [^2] We assume $\forall n$, if $k\neq i$, then $d_{kn}\neq d_{in}$.. In this case, the learning rate $\gamma_{k}^{(\tau)}\to 1/N_{k}^{(\tau)}$ where $N_{k}^{(\tau)}$ is the number of data points closest to $\mathbf{c}_{k}^{(\tau)}$, and the updating procedure (25) approaches to Lloyd’s algorithm.
III-D3 Lloyd’s algorithm as a Newton’s method
The smooth objective function $J_{\text{LSE}}$ possesses the second-order partial derivative given by
$\partial_{\mathbf{c}_{k}}^{2}{J_{\text{LSE}}}=\frac{\partial^{2}{J_{\text{LSE}}}}{\partial{\mathbf{c}_{k}}^{2}}$ $=-\sum_{n=1}^{N}\bigg{(}\frac{e^{-\lambda d_{kn}}}{\sum_{i}e^{-\lambda d_{in}}}\frac{\partial^{2}d_{kn}}{\partial\mathbf{c}_{k}^{2}}$ (27)
$+\frac{-\lambda e^{-\lambda d_{kq}}\sum_{i\neq k}e^{-\lambda d_{in}}}{(\sum_{i}e^{-\lambda d_{in}})^{2}}\frac{\partial d_{kn}}{\partial\mathbf{c}_{k}}(\frac{\partial d_{kn}}{\partial\mathbf{c}_{k}})^{\mathsf{T}}\bigg{)}$
$=\sum_{n=1}^{N}\frac{e^{-\lambda d_{kn}}}{\sum_{i}e^{-\lambda d_{in}}}(\mathbf{I}+\xi_{kn}\mathbf{D}_{kn})$
where $\xi_{kn}=\frac{\lambda\sum_{i\neq k}e^{-\lambda d_{in}}}{\sum_{i}e^{-\lambda d_{in}}}$ and $\mathbf{D}_{kn}=(\mathbf{x}_{n}-\mathbf{c}_{k})(\mathbf{x}_{n}-\mathbf{c}_{k})^{\mathsf{T}}$. By employing Newton’s method, the minimizer of $J_{\text{LSE}}$ can be iteratively found using
$\mathbf{c}_{k}^{(\tau+1)}=\mathbf{c}_{k}^{(\tau)}-[\partial_{\mathbf{c}_{k}}^{2}{J_{\text{LSE}}}(\mathbf{c}_{k}^{(\tau)})]^{-1}\partial_{\mathbf{c}_{k}}{J_{\text{LSE}}}(\mathbf{c}_{k}^{(\tau)}).$ (28)
As $\lambda\to+\infty$, $\partial_{\mathbf{c}_{k}}^{2}{J_{\text{LSE}}(\mathbf{c}_{k}^{(\tau)})}\to N_{k}^{(\tau)}\mathbf{I}$, and the updating procedure (28) asymptotically approaches Lloyd’s algorithm. This indicates that Lloyd’s algorithm is essentially Newton’s method, which aligns with the perspective presented in [29]. There is a close relationship between the learning rate value (26) and the second-order partial derivative of $J_{\text{LSE}}$ (27): the gradient descent with the learning rate of  (26) is equivalent to Newton’s method considering only a portion of the Hessian matrix (ignoring a rank-1 matrix $\mathbf{D}_{kn}$). Consequently, MEFC can be viewed as an approximated Newton’s method. A discussion of the pros and cons of optimizing $J_{\text{LSE}}$ using gradient descent or Newton’s method would be valuable, but it is beyond the scope of this paper. In the following section, we will observe a similar connection between Newton’s method and FKM.
III-D4 The relationship between p-Norm and fuzzy K-means
By substituting the minimum function in WCSS with the $p$-Norm function (20), we have
$\min_{\mathbf{c}_{1},\cdots,\mathbf{c}_{K}}J_{\text{PN}}(\mathbf{c}_{1},\cdots,\mathbf{c}_{K})=\sum_{n=1}^{N}(d_{1n}^{-p}+\cdots+d_{Kn}^{-p})^{-1/p}.$ (29)
The objective $J_{\text{PN}}$ has the first partial derivative as
$\partial_{\mathbf{c}_{k}}{J_{\text{PN}}}=\frac{\partial{J_{\text{PN}}}}{\partial{\mathbf{c}_{k}}}=-\sum_{n=1}^{N}\frac{d_{kn}^{-p-1}}{(\sum_{i=1}^{K}d_{in}^{-p})^{1/p+1}}(\mathbf{x}_{n}-\mathbf{c}_{k}).$ (30)
The minimizer of $J_{\text{PN}}$ can be located using gradient descent iteration
$\mathbf{c}_{k}^{(\tau+1)}=\mathbf{c}_{k}^{(\tau)}-\gamma^{(\tau)}_{k}\partial_{\mathbf{c}_{k}}{J_{\text{PN}}}(\mathbf{c}_{k}^{(\tau)}).$ (31)
When setting $p=1/(m-1)$ and
$\gamma^{(\tau)}_{k}=1/\sum_{n=1}^{N}\frac{(d_{kn}^{(\tau)})^{-p-1}}{(\sum_{i=1}^{K}(d_{in}^{(\tau)})^{-p})^{1/p+1}},$ (32)
this gradient descent optimization is identical to the FKM algorithm.
The connection between FKM and Newton’s method becomes evident through the second-order partial derivative of $J_{\text{PN}}$, which is
$\partial^{2}_{\mathbf{c}_{k}}{J_{\text{PN}}}=\frac{\partial^{2}{J_{\text{PN}}}}{\partial{\mathbf{c}_{k}}^{2}}=\sum_{n=1}^{N}\frac{d_{kn}^{-p-1}}{(\sum_{i}d_{in}^{-p})^{1/p+1}}(\mathbf{I}+\zeta_{kn}\mathbf{D}_{kn}),$ (33)
where $\zeta_{kn}=(p+1)\frac{\sum_{i\neq k}d_{in}^{-p}}{d_{kn}\sum_{i}d_{in}^{-p}}$. Hence, FKM can also be viewed as an approximated Newton’s method that ignores the rank-1 matrix $\mathbf{D}_{kn}$.
III-E From Boltzmann Operator to A Novel Clustering Algorithm
Employing the Boltzmann operator to smooth WCSS results in a new objective function, which is
$\min_{\mathbf{c}_{1},\cdots,\mathbf{c}_{K}}J_{\text{B}}(\mathbf{c}_{1},\cdots,\mathbf{c}_{K})=\sum_{n=1}^{N}\frac{\sum_{i}d_{in}e^{-\alpha d_{in}}}{\sum_{i}e^{-\alpha d_{in}}}.$ (34)
The objective $J_{B}$ possesses the first-order partial derivative of
$\partial_{\mathbf{c}_{k}}J_{\text{B}}=\frac{\partial J_{\text{B}}}{\partial\mathbf{c}_{k}}=-\sum_{n=1}^{N}$ $\frac{e^{-\alpha d_{kn}}}{\sum_{i}e^{-\alpha d_{in}}}\big{[}1-\alpha(d_{kn}$ (35)
$-\frac{\sum_{i}d_{in}e^{-\alpha d_{in}}}{\sum_{i}e^{-\alpha d_{in}}})\big{]}(\mathbf{x}_{n}-\mathbf{c}_{k}).$
The minimizer of $J_{B}$ can be found using gradient descent iteration
$\mathbf{c}_{k}^{(\tau+1)}=\mathbf{c}_{k}^{(\tau)}-\gamma^{(\tau)}_{k}\partial_{\mathbf{c}_{k}}J_{\text{B}}(\mathbf{c}_{1}^{(\tau)},\cdots,\mathbf{c}_{K}^{(\tau)}),$ (36)
where
$\gamma^{(\tau)}_{k}=1/\bigl{(}\sum_{n=1}^{N}\frac{e^{-\alpha d_{kn}^{(\tau)}}}{\sum_{i}e^{-\alpha d_{in}^{(\tau)}}}[1-\alpha(d_{kn}^{(\tau)}-\frac{\sum_{i}d_{in}^{(\tau)}e^{-\alpha d_{in}^{(\tau)}}}{\sum_{i}e^{-\alpha d_{in}^{(\tau)}}})]\bigl{)}.$ (37)
This updating procedure can be reformulated into a two-step iteration procedure, akin to FKM, which entails finding centroids using
$\mathbf{c}_{k}^{(\tau+1)}=\frac{\sum_{n}w_{kn}^{(\tau)}\mathbf{x}_{n}}{\sum_{n}w_{kn}^{(\tau)}},$ (38)
where weights $w_{kn}^{(\tau)}$ are given by
$w_{kn}^{(\tau)}=\frac{e^{-\alpha d_{kn}^{(\tau)}}}{\sum_{i}e^{-\alpha d_{in}^{(\tau)}}}[1-\alpha(d_{kn}^{(\tau)}-\frac{\sum_{i}d_{in}^{(\tau)}e^{-\alpha d_{in}^{(\tau)}}}{\sum_{i}e^{-\alpha d_{in}^{(\tau)}}})].$ (39)
The time complexity of one iteration is $O(NK^{2})$. Although the sum of weights for each instance equals one:
$\sum_{k=1}^{K}w_{kn}^{(\tau)}=1,$ (40)
it is worth noting that these weights cannot be interpreted as probabilities nor memberships, since some values are negative. Section III-G will discuss how to define membership in EKM.
III-F Physical Interpretation of Equilibrium K-Means
The second law of thermodynamics asserts that in a closed system with constant external parameters (e.g., volume) and fixed entropy, the internal energy will reach its minimum value at the state of thermal equilibrium. The objective (34) of EKM adheres to this minimum energy principle. This connection can be established by envisioning data points as particles with discrete/quantized energy levels, where the energy value corresponds to the squared Euclidean distance between a data point and a centroid, and the number of energy levels is equivalent to the number of centroids.
According to Boltzmann’s law, at the state of thermodynamic equilibrium, the probability of a particle occupying a specific energy level decreases exponentially with the energy value of that level. Hence, the objective function (34) is equivalent to the expectation of energy for the entire system, and the proposed clustering algorithm seeks centroids that minimize this expectation of energy. Due to this connection, we refer to the proposed algorithm (alternating between (38) and (39)) as equilibrium K-means.
III-G Membership Defined in Equilibrium K-Means
According to the physical interpretation of EKM, the exponential term $e^{-\alpha d_{kn}}$ can be regarded as the unnormalized probability of the $n$-th data point belonging to the $k$-th clusters. Hence, the membership of the $n$-th data point to the $k$-th clusters in EKM can be defined as
$u_{kn}=\frac{e^{-\alpha d_{kn}}}{\sum_{i=1}^{K}e^{-\alpha d_{in}}}.$ (41)
Note that, although the membership formula of EKM is the same as that (10) of MEFC, the membership values obtained from EKM and MEFC are typically different because their centroids are calculated using distinct formulas. This paper does not intend to compare the memberships in different fuzzy clustering algorithms; therefore, we only define membership in EKM without further discussion.
III-H Convergence of Smooth K-Means
HKM, FKM, MEFC, and EKM can be viewed as special cases of SKM, which can be generalized as the following gradient descent algorithm:
$\mathbf{c}_{k}^{(\tau+1)}=\mathbf{c}_{k}^{(\tau)}-\gamma^{(\tau)}_{k}\partial_{\mathbf{c}_{k}}J(\mathbf{c}_{1}^{(\tau)},\cdots,\mathbf{c}_{K}^{(\tau)}),$ (42)
where $J(\mathbf{c}_{1}^{(\tau)},\cdots,\mathbf{c}_{K}^{(\tau)})=\sum_{n=1}^{N}h(d_{1n}^{(\tau)},\cdots,d_{Kn}^{(\tau)})$, $h$ is a smooth minimum function, $d_{kn}^{(\tau)}=\frac{1}{2}\|\mathbf{x}_{n}-\mathbf{c}_{k}^{(\tau)}\|_{2}^{2}$, and the learning rate $\gamma^{(\tau)}_{k}$ is given by
$\gamma^{(\tau)}_{k}=1/\sum_{n=1}^{N}\bigg{(}\frac{\partial h}{\partial d_{kn}}(d_{1n}^{(\tau)},\cdots,d_{Kn}^{(\tau)})\bigg{)}.$ (43)
The convergence condition of this gradient descent algorithm is given as follows:
Theorem 1 (Convergence Condition):
The centroid sequence obtained by (42) converges to a (local) minimizer or saddle point of the objective function $J$ if the following conditions can be satisfied:
  1. 1.
    (Concavity) The function $h$ is a concave function at its domain $[0,\,+\infty)^{K}$.
  2. 2.
    (Boundness) The function $h$ has a lower bound, i.e., $h>-\infty$, and the learning rate set $\{\gamma^{(\tau)}_{k}\}_{\tau,k}$ has a positive lower bound, i.e., $\exists\epsilon>0$, such that $\gamma^{(\tau)}_{k}\geq\epsilon$ for all $\tau$ and $k$.
Proof.
See the Appendix for the proof, which generalizes the proof of convergence of fuzzy K-means in [30]. ∎
It can be easily verified that when the smooth minimum function $h$ is LogSumExp (19) or p-Norm (20), the above convergence condition can be satisfied with any initial centroids. Although the Boltzmann operator (22) is not concave, it is a smooth approximation of a concave function (the minimum function). Hence, EKM also demonstrates good convergence behavior, which has been empirically verified in our experiments.
IV Comparison of Different Smoothed Objectives
IV-A Case Study
This section explores and compares the behavior of different KM algorithms by examining their new objective functions in some examples. We select these examples because their data structures are well-designed, making it easier to illustrate the clustering characteristics.
We generate one-dimensional datasets comprising two classes of data points by sampling from two normal distributions, drawing $N_{1}$ samples from a distribution with mean $\mu_{1}$ and unit variance, and $N_{2}$ samples from another with mean $\mu_{2}$ and unit variance. Using different parameter combinations, we create four datasets: 1. A balanced, non-overlapping dataset ($N_{1}=N_{2}=50$, $\mu_{1}=-5$, and $\mu_{2}=+5$; Fig. 2a); 2. A balanced, overlapping dataset ($N_{1}=N_{2}=50$, $\mu_{1}=-0.5$, and $\mu_{2}=+0.5$; Fig. 2b); 3. An imbalanced, non-overlapping dataset ($N_{1}=5000$, $N_{2}=50$, $\mu_{1}=-5$, and $\mu_{2}=+5$; Fig. 2c); 4. An imbalanced, compact dataset ($N_{1}=2000$, $N_{2}=50$, $\mu_{1}=-2$, and $\mu_{2}=+2$; Fig. 2d).
We fix the first centroid at $\mu_{1}$, and plot the objectives of HKM (WCSS (12)), FKM ($J_{\text{PN}}$ (29)), MEFC ($J_{\text{LSE}}$ (23)), and EKM ($J_{\text{B}}$ (34)) as a function of the position of the second centroid in Fig. 2. Fig. 2a demonstrates the similarity in objective functions for the four clustering algorithms, indicating comparable performance on a balanced dataset. Fig. 2b reveals that FKM and MEFC are better suited for overlapped data, as they are less prone to a local minimum. In Fig. 2c and Fig. 2d, data imbalance cases show that the local and global minimum points of the objective functions of HKM, FKM, and MEFC are biased towards the center of the large class, i.e., $\mu_{1}$. In contrast, EKM does not have an obvious local minimum and its global minimum point aligns with the correct cluster center, i.e., $\mu_{2}$, highlighting its superiority in handling imbalanced data.
Figure 2: Objectives of four KM algorithms as a function of the first centroid fixed at $\mu_{1}$ and the second centroid moving along the $x$-axis. The yellow, purple, green, and blue curves are the objective functions of HKM, FKM, MEFC, and EKM, respectively. (a), (b), (c), and (d) present objective functions for different data distributions.
IV-B The Analysis of EKM’s Effectiveness on Imbalanced Data
We demonstrate the effectiveness of EKM by analyzing the gradient of its objective. The gradient of the smoothed WCSS with respect to the distance between data points and centroids can be interpreted as the force exerted by a spring. Positive gradient values represent attractive forces, while negative values represent repulsive forces. The gradient is also associated with membership, where the membership (7) of FKM is equivalent to a power exponent ($1/m$) of $\partial J_{\text{PN}}/\partial d_{kn}$, and the membership (10) of MEFC is equivalent to $\partial J_{\text{LSE}}/\partial d_{kn}$.
A simple example is given to plot the gradient values of different smoothed objectives. We fix two centroids at $-1$ and $+1$, respectively, and move a data point along the $x$-axis. In Fig. 3a we display the gradients of WCSS, $J_{\text{LSE}}$, $J_{\text{PN}}$, and $J_{\text{B}}$ with respect to the distance between the data point and the second centroid. As evident from the figure, data points on the side of the first centroid do not impact the second centroid of HKM, but they do attract the second centroids of FKM and MEFC. This finding supports the previous result in [31] that FKM has a stronger uniform effect than HKM. However, for EKM, data points near the first centroid have repulsive forces on the second centroid, which compensates for the attraction of points on the side of the first centroid to the second centroid, effectively reducing the uniform effect. Furthermore, it is worth noting that for EKM, data points near the second centroid have the strongest attraction forces. Thus, the second centroid is anchored by these data points, making its position less susceptible to noise or outliers. This is the advantage of EKM over imbalanced data clustering algorithms that assign weights to data points based on cluster size.
Figure 3: Gradients of four KM algorithms as a function of a data point moving along the $x$-axis. (a) Gradients of HKM, FKM, MEFC, and EKM with fixed smoothness parameters. (b) Gradients of FKM. (c) Gradients of MEFC. (d) Gradients of EKM.
IV-C The Choice of $\alpha$
The smoothing parameter $\alpha$ has an impact on the clustering result in EKM, but the optimal choice remains unknown, which is not a difficulty unique to EKM. The FKM algorithm also struggles with selecting the optimal fuzzifier value, $m$. Despite numerous studies discussing the selection of $m$, a widely accepted solution has yet to be found [32].
Determining optimal parameters for a clustering algorithm often involves trial and error, as the real data structure is unknown. As a rule of thumb, when the data space dimension is less than or equal to three, setting $\alpha=1$ appears effective after normalizing the data to have a zero mean and unit variance for each dimension. As the data space dimension increases, the data-centroid distance $d_{kn}$ tends to grow, necessitating a decrease in $\alpha$ to ensure that the exponential term $e^{-\alpha d_{kn}}$ remains within a normal range. In this case, we suggest setting $\alpha$ to ten times the data variance and gradually reducing it. A proper $\alpha$ value indicates the presence of repulsive forces, resulting in a sudden increase in centroid distances. Ultimately, techniques for visualizing clustering results can be employed to fine-tune the $\alpha$ value. This strategy is proven effective by experimental examples of deep clustering on the MNIST dataset in Section VI. We further discuss and validate the choice of $\alpha$ in the supplemental material.
V Numerical Experiments
Numerical experiments are conducted to compare the performance of our proposed EKM algorithm with six other algorithms: (1) HKM, (2) FKM, (3) MEFC, (4) csiFKM [19], (5) siibFKM [20], and (6) the Gaussian mixture model. Multiprototype clustering algorithms are not used as benchmarks, as the multiprototype mechanisms can be combined with EKM for fair comparison. This is beyond the scope of this paper because the paper does not intend to find the best KM algorithm to pair with the multiprototype mechanism.
To avoid local optima convergence, each clustering algorithm is executed ten times, presenting the lowest objective value result. Centroids are initialized independently for each replication using the K-means++ algorithm [33]. Convergence is deemed achieved when the relative moving distance of centroids between successive iterations is less than $1\mathrm{e}{-3}$, i.e.,
$\frac{\bigg{(}\sum_{k=1}^{K}\|\mathbf{c}_{k}^{(\tau)}-\mathbf{c}_{k}^{(\tau-1)}\|_{2}^{2}\bigg{)}^{1/2}}{\bigg{(}\sum_{k=1}^{K}\|\mathbf{c}_{k}^{(\tau)}\|_{2}^{2}\bigg{)}^{1/2}}\leq 1\mathrm{e}{-3}.$ (44)
The maximum number of iterations is set to 100. FKM, csiFKM, and siibFKM employ a typical fuzzifier value of $m=2$ [34]. The fuzzifier value of MEFC is set to $\lambda=1$ and the smoothness value of EKM is set to $\alpha=1$. We test the algorithms on 12 datasets, including two synthetic and ten real datasets. Because of space limitations, only partial experiment results from six datasets are shown in the paper. Full experiment results with quantitative metrics can be found in the supplemental material. The six datasets are introduced below.
V-A Datasets
V-A1 Synthetic 2-D Ball
We create a synthetic dataset with data independently generated from three distinct normal distributions, each having different means: $(-2,2)$, $(2,-2)$, and $(4,4)$. Their covariance matrices are all identity matrices. The dataset consists of 2,000 samples from the normal distribution with means $(-2,2)$, and 50 samples from the other two normal distributions, resulting in a total of 2,100 samples. The clustering results are shown in Fig. 4.
V-A2 Synthetic 2-D Noisy Ball
We add 100 uniformly distributed noise data samples (around 5% of the number of clean samples) to the synthetic 2-D ball dataset to test the algorithms’ robustness against noise. The clustering outcomes are displayed in Fig. 5.
V-A3 Fisher’s Iris Data
The Fisher’s Iris dataset [35, 36] is a well-known dataset for evaluating the performance of clustering algorithms. It comprises 50 samples from each of three iris species (Iris setosa, Iris virginica, and Iris versicolor). Two features, the width and the length of the sepals, are utilized as clustering features and normalized to have zero means and unit variance. Fig. 6 depicts the clustering results.
V-A4 Imbalanced Fisher’s Iris Data
Because the instance distribution in the Iris dataset is balanced. To examine the algorithms’ capability for imbalanced clustering, we generate an imbalanced Fisher’s Iris data by removing the first 30 instances of Iris setosa, and merging the other two species (Iris virginica and Iris versicolor) into a single class. The clustering results for this dataset are presented in Fig. 7.
V-A5 Wisconsin Diagnostic Breast Cancer
Wisconsin Diagnostic Breast Cancer (WDBC) [37] comprises 30 features of cell nuclei from diagnostic benign and malignant breast tumors, obtained from digitized images of fine-needle aspiration of breast masses. This database contains 569 instances, with a fairly even distribution of 357 benign and 212 malignant cases. We utilize the first three features (mean radius, mean texture, and mean perimeter) for clustering, each normalized with zero mean and unit variance. Fig. 9 presents the clustering results where the mean radius and the mean texture are used as coordinates for visualization.
V-A6 Imbalanced Wisconsin Diagnostic Breast Cancer
The data distribution in WDBC is quite balanced. To evaluate the imbalanced data clustering performance, we remove the first 200 malignant instances in WDBC, resulting in an imbalanced WDBC database with a class distribution of 357 benign and 12 malignant instances. We still use the first three features for clustering, each normalized with zero mean and unit variance. The clustering results are shown in Fig. 9.
V-A7 ColoredChips
This is not a dataset but an image named coloredChips.png from Matlab’s image gallery (shown in Fig. 10a). The purpose of this experiment is to test the algorithms’ performance on the task of color-based image segmentation. The number of clusters is set to five to expect the segmentation of the four colored chips and the background (i.e., the desk surface). Clustering is conducted in the RGB space and the segmentation results can be viewed in Fig. 10b - 10e.
V-B Discussion of Results
In general, these results demonstrate that the proposed EKM algorithm performs competitively on balanced data and outperforms the other clustering algorithms on imbalanced data. In particular, HKM, FKM, and MEFC cannot find the correct clusters in imbalanced datasets. The csiFKM and siibFKM algorithms are sensitive to noise data. Fig. 5c and Fig. 5d show that their centroids are highly deviated from the correct positions because of noise. In comparison, EKM is more robust, as seen in Fig. 5e.
Fig. 9 displays the clustering results of HKM, FKM, GMM, and EKM for the imbalanced WDBC dataset. The strong performance of EKM indicates its potential application in anomaly detection, particularly in identifying clustered anomalies. Although HKM and FKM are also used for anomaly detection, they typically detect isolated anomalies, not clustered ones [38, 39, 40, 41].
The color-based image segmentation results are shown in Fig. 10b - 10e. It can be seen that HKM, FKM, and MEFC divide the background (the desk surface) into two clusters. This happens because the background comprises most pixels, and splitting it into two clusters better balances the number of pixels within the clusters. On the other hand, EKM produces the expected segmentation result.
Figure 4: (a) Synthetic 2-D ball data. (b) Clustering by HKM. (c) Clustering by FKM. (d) Clustering by MEFC. (e) Clustering by EKM.
Figure 5: (a) Synthetic 2-D noisy ball data. (b) Clustering by FKM. (c) Clustering by csiFKM. (d) Clustering by siibFKM. (e) Clustering by EKM.
Figure 6: (a) Fisher’s Iris data. (b) Clustering by HKM. (c) Clustering by FKM. (d) Clustering by MEFC. (e) Clustering by EKM.
Figure 7: (a) Imbalanced Fisher’s Iris data. (b) Clustering by HKM. (c) Clustering by FKM. (d) Clustering by MEFC. (e) Clustering by EKM.
Figure 8: (a) Wisconsin Diagnostic Breast Cancer dataset. Three features are used for clustering and two of them are used for visualization. (b) Clustering by HKM. (c) Clustering by FKM. (d) Clustering by GMM. (e) Clustering by EKM.
Figure 9: (a) Imbalanced Wisconsin Diagnostic Breast Cancer dataset. (b) Clustering by HKM. (c) Clustering by FKM. (d) Clustering by GMM. (e) Clustering by EKM.
Figure 10: (a) coloredChips.png (b) Segmentation by HKM. (c) Segmentation by FKM. (d) Segmentation by MEFC. (e) Segmentation by EKM.
VI Deep Clustering
Clustering algorithms using Euclidean distance for defining similarity often face challenges when clustering high-dimensional data, such as images, due to the small distance differences between various point pairs in high-dimensional space [42]. Deep clustering is a technique that addresses this issue by employing DNNs to map high-dimensional data non-linearly to low-dimensional representation, making the subsequent clustering easier. Most existing deep clustering approaches combine DNNs with HKM, e.g., [12, 13, 14]. However, this design is less favorable to imbalanced datasets [13]. In this section, we provide an example of replacing HKM with EKM in a popular deep clustering framework, showing its performance improvement on an imbalanced dataset derived from the MNIST.
VI-A Optimization Procedure
Figure 11: Illustration of the DCN framework [12]. The parameters of the encoder, decoder, and clustering model are optimized jointly to minimize the reconstruction error and clustering error.
Deep clustering network (DCN) [12] is a popular deep clustering framework, as illustrated in Fig. 11. Essentially, DCN maps high-dimensional data to a low-dimensional representation through an autoencoder network. An autodecoder follows the autoencoder, mapping the representation back to the original high-dimension space (i.e., reconstruction). To ensure that the low-dimensional representation maintains the primary information of the original data, the autoencoder and autodecoder are jointly trained to minimize the reconstruction error. Additionally, to create a clustering-friendly structure in the low-dimensional representation, a clustering error is minimized along with the reconstruction error. DCN uses an alternating optimization algorithm to minimize the total error, and the optimization process is described below.
First, the autoencoder and the autodecoder are jointly trained to reduce the following loss for the incoming data $\mathbf{x}_{n}$:
$\min_{\bm{\theta}_{e},\bm{\theta}_{d}}L^{n}=l(\bm{g}(\bm{f}(\mathbf{x}_{n})),\mathbf{x}_{n})+\beta\|\bm{f}(\mathbf{x}_{n})-\mathbf{C}\mathbf{s}_{n}\|_{2}^{2},$ (45)
where $\bm{f}(\cdot)$ and $\bm{g}(\cdot)$ are simplified symbols for autoencoder $\bm{f}(\cdot;\,\bm{\theta}_{e})$ and autodecoder $\bm{g}(\cdot;\,\bm{\theta}_{d})$, respectively. The function $l(\cdot)$ is the least-squares loss $l(\hat{\mathbf{x}},\mathbf{x})=\|\hat{\mathbf{x}}-\mathbf{x}\|_{2}^{2}$ to measure the reconstruction error. The assignment vector $\mathbf{s}_{n}\in\mathbb{R}^{K\times 1}$ has only one non-zero element and $\mathbf{1}^{\mathsf{T}}\mathbf{s}_{n}=1$, indicating which cluster the $n$-th data belongs to, and the $k$-th column of $\mathbf{C}=[\mathbf{c}_{1},\cdots,\mathbf{c}_{K}]$ is the centroid of the $k$-th cluster. The parameter $\beta$ balances the reconstruction error versus the clustering error. Then, the network parameters $\{\bm{\theta}_{e},\bm{\theta}_{d}\}$ are fixed, and the parameters $\{\mathbf{s}_{n}\}$ are updated as follows:
$s_{j,n}\leftarrow\begin{cases}1,&\text{if}\,j=\operatorname*{arg\,min}_{k=\{1,\cdots,K\}}\|\bm{f}(\mathbf{x}_{n})-\mathbf{c}_{k}\|_{2},\\ 0,&\text{otherwise,}\end{cases}$ (46)
where $s_{j,n}$ is the $j$-th element of $\mathbf{s}_{n}$. Finally, $\mathbf{C}$ is updated by the batch-learning version of the HKM algorithm:
$\mathbf{c}_{k}\leftarrow\mathbf{c}_{k}+(1/m_{k}^{n})(\bm{f}(\mathbf{x}_{n})-\mathbf{c}_{k})s_{k,n},$ (47)
where $m_{k}^{n}$ is the number of samples assigned to the $k$-th cluster before the incoming data $\mathbf{x}_{n}$, controlling the learning rate of the $k$-th centroid. Overall, the optimization procedure of DCN with HKM alternates between updating networks parameters $\{\bm{\theta}_{e},\bm{\theta}_{d}\}$ by solving (45) and updating HKM parameters $\{\mathbf{C},\{\mathbf{s}_{n}\}\}$ by (46) and (47).
However, the centroid updating rule (47) is problematic for imbalanced data due to the uniform effect. To address this issue, we propose to replace (47) with the batch-learning version of EKM:
$\mathbf{c}_{k}$ $\leftarrow\mathbf{c}_{k}-(1/m_{k}^{n})\partial_{\mathbf{c}_{k}}J_{B}(\mathbf{c}_{1},\cdots,\mathbf{c}_{K})$ (48)
$=\mathbf{c}_{k}+(1/m_{k}^{n})\frac{e^{-\alpha d_{kn}}}{\sum_{i=1}^{K}e^{-\alpha d_{in}}}\big{[}1-\alpha(d_{kn}$
$-\frac{\sum_{i=1}^{K}d_{in}e^{-\alpha d_{in}}}{\sum_{i=1}^{K}e^{-\alpha d_{in}}})\big{]}(\bm{f}(\mathbf{x}_{n})-\mathbf{c}_{k}),$
where $d_{kn}=\frac{1}{2}\|\bm{f}(\mathbf{x}_{n})-\mathbf{c}_{k}\|_{2}^{2}$. There are other details and tricks to implement DCN, such as the initialization of the networks. We only introduce the part related to our contribution, and for the rest, the readers can refer to the original paper of DCN [12].
VI-B Clustering Performance on MNIST
To implement DCN, we refer to the code provided by one of its authors, available at https://github.com/boyangumn/DCN-New. We use the default neural network structure and hyperparameters. In particular, the deep representation dimension is set to ten, and the parameter $\beta$ is set to one. The smoothness parameter $\alpha$ for EKM is determined by the strategy introduced in Section IV-C. We evaluate clustering performance using three widely accepted metrics: including normalized mutual information (NMI), adjusted Rand index (ARI), and clustering accuracy (ACC). NMI and ACC range from 0 to 1, with zero representing the worst and one representing the best. ARI ranges from -1 to 1, with minus one representing the worst and one representing the best.
We first validate the algorithms on the full MNIST dataset [43], containing 70,000 gray images of handwritten digits from 0 to 9. We set the smoothness parameter of EKM to $\alpha=5\mathrm{e}-3$. The clustering results are presented in TABLE I, where we compare the proposed DCN+EKM with DCN+HKM and stacked autoencoder (SAE). SAE is a specific version of DCN that only minimizes the reconstruction error. We omit the results of DCN+FKM and DCN+MEFC due to space limitations and their poorer performance compared to DCN+HKM. The results show that DCN outperforms SAE, highlighting the importance of a clustering-friendly representation. The table also reveals that DCN+EKM performs similarly to DCN+HKM for this balanced dataset.
To evaluate the performance of the algorithms for imbalanced data clustering, we remove the MNIST dataset’s training set samples with numbers 1 to 9 (leaving only the test set samples). This imbalanced dataset has 15,923 images, with around 6,900 being the number 0 and the remaining numbers (1 to 9) being around 1,000. We set the smoothness parameter of EKM to $\alpha=3.8\mathrm{e}-3$. The clustering results are summarised in TABLE II, showing that DCN+EKM significantly outperforms DCN+HKM. To analyze the reason, we use t-SNE [44] to map the ten-dimensional representation obtained by DCN to a two-dimensional space for visualization, with the results displayed in Fig. 12. We observe that the deep representation obtained by DCN+EKM is more discriminative than DCN+HKM, and DCN+EKM successfully separates the large class (number 0) from other small classes, while DCN+HKM erroneously groups the large class into four clusters.
TABLE I: Evaluation on full MNIST
Methods SAE+HKM DCN+HKM SAE+EKM DCN+EKM NMI 0.725 0.798 0.711 0.813 ARI 0.667 0.744 0.642 0.731 ACC 0.795 0.837 0.782 0.808
TABLE II: Evaluation on imbalanced MNIST
Methods SAE+HKM DCN+HKM SAE+EKM DCN+EKM NMI 0.551 0.584 0.583 0.701 ARI 0.317 0.325 0.396 0.826 ACC 0.413 0.434 0.497 0.784
Figure 12: Visualization of low-dimensional DCN representations by t-SNE. (a) Representations obtained by DCN+HKM, colored by true labels. (b) Representations obtained by DCN+HKM, colored by clusters of HKM. (c) Representations obtained by DCN+EKM, colored by true labels. (d) Representations obtained by DCN+EKM, colored by clusters of EKM.
VII Conclusion
This paper presents equilibrium K-means (EKM), a novel clustering algorithm that is effective for imbalanced data. EKM is simple, interpretable, scalable to large datasets, and capable of finding unequal-sized clusters, making it a better alternative for hard K-means (HKM) and fuzzy K-means (FKM) for imbalanced data. Experimental results on real datasets from various domains demonstrate that EKM outperforms HKM and FKM on datasets with imbalanced data and performs comparably on datasets with balanced data. Furthermore, we provide a general framework that unifies HKM, FKM, and EKM. We encourage readers to explore and develop new K-means algorithms based on this more flexible framework to address other specific K-clustering problems besides imbalanced data distribution.
Appendix A Proof of Theorem 1
For concise and tidy, we denote the partial derivative $\frac{\partial h}{\partial d_{k}}(d_{1n}^{(\tau)},\cdots,d_{Kn}^{(\tau)})$ as $\frac{\partial h}{\partial d_{k}}|_{n,{\tau}}$. Since $h$ is a concave function at $[0,\,+\infty)^{K}$, we have
$h(d_{1n},\cdots,d_{Kn})$ (49)
$\leq h(d_{1n}^{(\tau)},\cdots,d_{Kn}^{(\tau)})+\sum_{k=1}^{K}\frac{\partial h}{\partial d_{k}}\bigg{|}_{n,{\tau}}\cdot(d_{kn}-d_{kn}^{(\tau)}),$
which holds for any $n\in\{1,\cdots,N\}$ and $d_{1n},\cdots,d_{Kn}\in[0,\,+\infty)$. Summing over $n$, it follows that
$J(\mathbf{c}_{1},\cdots,\mathbf{c}_{K})\\ \leq\sum_{n=1}^{N}\bigg{[}h(d_{1n}^{(\tau)},\cdots,d_{Kn}^{(\tau)})+\sum_{k=1}^{K}\frac{\partial h}{\partial d_{k}}\bigg{|}_{n,{\tau}}\cdot(d_{kn}-d_{kn}^{(\tau)})\bigg{]}\\ =J(\mathbf{c}_{1}^{(\tau)},\cdots,\mathbf{c}_{K}^{(\tau)})+\frac{1}{2}\sum_{k=1}^{K}\sum_{n=1}^{N}\bigg{(}\frac{\partial h}{\partial d_{k}}\bigg{|}_{n,{\tau}}\cdot\big{(}\|\mathbf{x}_{n}-\mathbf{c}_{k}\|_{2}^{2}\\ -\|\mathbf{x}_{n}-\mathbf{c}_{k}^{(\tau)}\|_{2}^{2}\big{)}\bigg{)}.$ (50)
Denote the function on the right side of the inequality as $M(\mathbf{c}_{1},\cdots,\mathbf{c}_{K})$. With the boundness condition, we have $\sum_{n=1}^{N}\frac{\partial h}{\partial d_{k}}\big{|}_{n,\tau}>0$ for any $k$ and $\tau$, thus, $M$ is a quadratic function and strictly convex, with the unique global minimizer at $(\mathbf{c}_{1}^{(\tau+1)},\cdots,\mathbf{c}_{K}^{(\tau+1)})$ defined by (42). Denote $J(\mathbf{c}_{1}^{(\tau)},\cdots,\mathbf{c}_{K}^{(\tau)})$ as $J^{(\tau)}$, $M(\mathbf{c}_{1}^{(\tau)},\cdots,\mathbf{c}_{K}^{(\tau)})$ as $M^{(\tau)}$, and $\partial_{\mathbf{c}_{k}}J(\mathbf{c}_{1}^{(\tau)},\cdots,\mathbf{c}_{K}^{(\tau)})$ as $\partial_{\mathbf{c}_{k}}J^{(\tau)}$. Each iteration of the centroid will reduce the objective function by
$J^{(\tau)}-J^{(\tau+1)}$ (51)
$\geq J^{(\tau)}-M^{(\tau+1)}$
$=\frac{1}{2}\sum_{k=1}^{K}\sum_{n=1}^{N}\frac{\partial h}{\partial d_{k}}\bigg{|}_{n,\tau}\cdot\big{(}\|\mathbf{x}_{n}-\mathbf{c}_{k}^{(\tau)}\|_{2}^{2}-\|\mathbf{x}_{n}-\mathbf{c}_{k}^{(\tau+1)}\|_{2}^{2}\big{)}$
substituting (42)
$=\frac{1}{2}\sum_{k=1}^{K}\sum_{n=1}^{N}\frac{\partial h}{\partial d_{k}}\bigg{|}_{n,\tau}\cdot\big{(}\|\mathbf{x}_{n}-\mathbf{c}_{k}^{(\tau)}\|_{2}^{2}-\|\mathbf{x}_{n}-\mathbf{c}_{k}^{(\tau)}$
$+\gamma_{k}^{(\tau)}\partial_{\mathbf{c}_{k}}J^{(\tau)}\|_{2}^{2}\big{)}$
$=\frac{1}{2}\sum_{k=1}^{K}\sum_{n=1}^{N}\frac{\partial h}{\partial d_{k}}\bigg{|}_{n,\tau}\cdot\big{(}-(\gamma_{k}^{(\tau)})^{2}\|\partial_{\mathbf{c}_{k}}J^{(\tau)}\|_{2}^{2}$
$-2\gamma_{k}^{(\tau)}(\mathbf{x}_{n}-\mathbf{c}_{k}^{(\tau)})^{\mathsf{T}}\partial_{\mathbf{c}_{k}}J^{(\tau)}\big{)}$
using $\gamma_{k}^{(\tau)}=\frac{1}{\sum_{n=1}^{N}\frac{\partial h}{\partial d_{k}}|_{n,\tau}}$
and $\partial_{\mathbf{c}_{k}}J^{(\tau)}=-\sum_{n=1}^{N}\frac{\partial h}{\partial d_{k}}\cdot(\mathbf{x}_{n}-\mathbf{c}_{k}^{(\tau)})$
$=\sum_{k=1}^{K}-\frac{1}{2}\gamma^{(\tau)}_{k}\|\partial_{\mathbf{c}_{k}}J^{(\tau)}\|_{2}^{2}+\sum_{k=1}^{K}\gamma_{k}^{(\tau)}\|\partial_{\mathbf{c}_{k}}J^{(\tau)}\|_{2}^{2}$
$=\sum_{k=1}^{K}\frac{1}{2}\gamma_{k}^{(\tau)}\|\partial_{\mathbf{c}_{k}}J^{(\tau)}\|_{2}^{2}$
with the boundness condition of $\gamma_{k}^{(\tau)}$
$\geq\frac{1}{2}\epsilon\sum_{k=1}^{K}\|\partial_{\mathbf{c}_{k}}J^{(\tau)}\|_{2}^{2},$
where $\epsilon$ is a positive number. Hence, the sequence $(J^{(1)},J^{(2)},\cdots)$ is non-increasing, and with the boundness condition that $h>-\infty$, we have $\lim_{\tau\to+\infty}(J^{(\tau)}-J^{(\tau+1)})\to 0$. If the left side of the inequality (51) converges to zero, the right side of the inequality also converges to zero since it is non-negative. Consequently, we have $\lim_{\tau\to+\infty}\partial_{\mathbf{c}_{k}}J^{(\tau)}\to 0$ for all $k$. Therefore, the sequence $(\mathbf{c}_{k}^{(\tau)})$ converges to a stationary point of the objective function $J$. Because $(J^{(1)},J^{(2)},\cdots)$ is non-increasing, only (local) minimizers or saddle points appear as limit points.
Supplemental Materials: Imbalanced Data Clustering using Equilibrium K-Means
supplement
We evaluate the clustering performance of the proposed Equilibrium K-means (EKM) on ten real datasets. Four datasets (Iris, Imbalanced Iris, Wisconsin Diagnostic Breast Cancer (WDBC), and Imbalanced WDBC) are discussed in detail in the main text. This supplemental material presents the results for the remaining six datasets from the UCI repository (available at https://archive.ics.uci.edu/datasets): Wine, Ecoli, Image Segmentation (IS), Htru2, Rice, and Dry Bean. A detailed description of all ten datasets can be found in TABLE SI, ordered according to the imbalance level measured by CV0, which we will define later. Datasets with a lower order have higher CV0 values, indicating a higher imbalance level. We provide some necessary explanations for this table below.
In the “Feature” column, the format “x (y)” indicates that the dataset has y features, with x features used for clustering. For instance, the Iris dataset has four features, with two used for clustering. Only the four datasets in the main text use partial features for clustering, which serves the purpose of visualizing clustering results. For the additional six datasets, all features are used since visualization is not required. In the “CV0” column, we calculate the coefficient of variation (CV) to measure data imbalance. CV is defined as the ratio of the standard deviation to the mean. CV is defined as the ratio of the standard deviation to the mean. Mathematically, given the number of instances in each class as $N_{1},\cdots,N_{K}$, we have
$\text{CV}=\text{std}(\mathbf{N})/\text{mean}(\mathbf{N}),$
where
$\text{mean}(\mathbf{N})=\frac{\sum_{k=1}^{K}N_{k}}{K},$
and
$\text{std}(\mathbf{N})=\sqrt{\frac{\sum_{k=1}^{K}(N_{k}-\text{mean}(\mathbf{N}))^{2}}{K-1}}.$
In general, the larger the CV value, the greater the imbalance of the data.
Each dataset is normalized so that each feature has zero mean and unit variance. The number of clusters is assumed to be equal to the number of classes for all datasets. For the supplementary six datasets, the smoothness parameter $\alpha$ is chosen as follows:
$\alpha=2/\bar{d}_{0n},$
where $\bar{d}_{0n}=\frac{1}{2}\sum_{n=1}^{N}\|\mathbf{x}_{n}\|_{2}^{2}/N$, $\mathbf{x}_{n}$ is the $n$-th data point, and $N$ is the total number of data points. The value $\bar{d}_{0n}$ can be interpreted as a data variance after the normalization of zero data mean. The normalized mutual information (NMI), adjusted Rand index (ARI), and clustering accuracy (ACC) metrics of each algorithm on the ten datasets are presented in TABLE SII. Bold values indicate the best results. We also include three more metrics: “CV1”, ”DCV” and ”Running time” for readers’ interest. “CV1” is the CV value of the clustering result. “DCV” is the difference between CV0 and CV1, i.e., $\text{DCV}=\text{CV}0-\text{CV}1$. A positive DCV value indicates that the data distribution in the true classes is more imbalanced than in the clusters, while a negative DCV value means the data distribution in the true classes is more uniform than in the clusters. “Running time” is the sum of the running time of ten replicates.
We observe that the proposed EKM generally outperforms the other six algorithms on the last four datasets with CV values greater than 0.5. EKM outperforms other algorithms on two of the top six datasets with CV values less than 0.5. On the rest of the datasets, EKM still has a competitive performance. The running time of EKM is also comparable to FKM. Overall, these results demonstrate that EKM is an efficient algorithm for clustering balanced and imbalanced data.
TABLE SI: Detailed description of REAL-WORLD Datasets
Dataset Instance Feature Class CV0
Iris 150 2 (4) 3 0
IS 2310 19 (19) 7 0
Wine 178 13 (13) 3 0.1939
Rice 3810 7 (7) 2 0.2042
WDBC 569 3 (30) 3 0.3604
Dry Bean 13611 16 (16) 7 0.4951
Imbalanced Iris 120 2 (4) 2 0.9428
Ecoli 336 7 (7) 8 1.1604
Imbalanced WDBC 369 3 (30) 2 1.3222
Htru2 17898 8 (8) 2 1.4142
TABLE SII: Experimental results of different algorithms on real-world datasets
Dataset Measurement HKM FKM MEFC csiFKM siibFKM GMM EKM
Iris NMI 0.5945 0.5671 0.5194 0.5671 0.5675 0.6775 0.5457
ARI 0.5462 0.5328 0.5291 0.5328 0.5281 0.5354 0.5134
ACC 0.7867 0.7800 0.7933 0.7800 0.7800 0.7000 0.7733
CV1 0.0346 0.0200 0.1217 0.0200 0.1442 0.8902 0.0346
DCV -0.0346 -$0$0.1217 -$0$0.1442 -0.8902 -0.0346
Running time 0.0113 0.0111 0.0305 0.0142 0.1844 0.0715 0.0328
IS NMI 0.5864 0.5953 0.6222 0.5971 0.5997 0.5657 0.6463
ARI 0.4605 0.4973 0.4995 0.4962 0.3214 0.4456 0.5161
ACC 0.5455 0.6455 0.5944 0.6433 0.5861 0.5641 0.5944
CV1 0.8246 0.2450 0.6433 0.4617 1.0377 0.5037 0.6582
DCV -0.8246 -0.2450 -0.6433 -0.4617 -1.0377 -0.5037 -0.6582
Running time 0.0321 0.4714 0.2616 0.7913 49.5369 1.0463 0.7205
Wine NMI 0.8759 0.8759 0.8759 0.8610 0.3361 0.8466 0.8920
ARI 0.8975 0.8975 0.8975 0.8804 0.1355 0.8649 0.9134
ACC 0.9663 0.9663 0.9663 0.9607 0.5562 0.9551 0.9719
CV1 0.1242 0.1242 0.1242 0.1219 1.1513 0.1070 0.1403
DCV 0.0696 0.0696 0.0696 0.0720 -0.9574 0.0868 0.0535
Running time 0.0075 0.0153 0.0088 0.0126 1.6469 0.0515 0.0094
Rice NMI 0.5685 0.5688 0.5699 0.5554 0.5415 0.4891 0.5653
ARI 0.6815 0.6824 0.6833 0.6642 0.6394 0.5649 0.6772
ACC 0.9129 0.9131 0.9134 0.9076 0.9000 0.8759 0.9115
CV1 0.2442 0.2183 0.2309 0.2977 0.3623 0.0163 0.2643
DCV -0.0401 -0.0141 -0.0267 -0.0935 -0.1581 0.1878 -0.0601
Running time 0.0200 0.0438 0.0357 0.1032 2.1275 0.2303 0.0397
WBDC NMI 0.4920 0.4870 0.5005 0.4423 0.2951 0.4618 0.4906
ARI 0.5844 0.5963 0.6257 0.4402 0.2797 0.5722 0.5340
ACC 0.8840 0.8875 0.8963 0.8366 0.7750 0.8787 0.8682
CV1 0.5990 0.5393 0.4051 0.8227 0.9768 0.2461 0.7133
DCV -0.2386 -0.1790 -0.0447 -0.4623 -0.6164 0.1143 -0.3529
Running time 0.0156 0.0226 0.0171 0.0385 1.7632 0.0894 0.0160
Dry Bean NMI 0.7138 0.7055 0.7050 0.6990 0.6673 0.7429 0.6859
ARI 0.6687 0.6676 0.6565 0.6607 0.6170 0.6914 0.5792
ACC 0.7865 0.8373 0.7834 0.8001 0.8032 0.8160 0.7478
CV1 0.5616 0.4697 0.5441 0.5154 0.5394 0.5480 0.6385
DCV -0.0666 0.0254 -0.0491 -0.0203 -0.0444 -0.0529 -0.1434
Running time 0.2186 1.3271 0.6290 1.8517 139.5876 3.4712 0.8499
Imbalanced Iris NMI 0.0001 0.0247 0.0061 0.6414 0.6870 0.9101 0.9101
ARI -0.0053 0.0049 -0.0023 0.7421 0.7865 0.9582 0.9582
ACC 0.5250 0.5500 0.5250 0.9500 0.9583 0.9917 0.9917
CV1 0.1179 0.0471 0.0236 1.0842 1.0607 0.9664 0.9664
DCV 0.8250 0.8957 0.9192 -0.1414 -0.1179 -0.0236 -0.0236
Running time 0.0057 0.0121 0.0142 0.0257 0.1474 0.0146 0.0115
Ecoli NMI 0.6370 0.5837 0.6087 0.5999 0.6026 0.6015 0.6530
ARI 0.4953 0.4113 0.4704 0.4176 0.4468 0.6149 0.5202
ACC 0.6310 0.5863 0.6190 0.6012 0.6012 0.6964 0.6458
CV1 0.6871 0.4671 0.7517 0.4482 0.7770 1.0893 0.7001
DCV 0.4733 0.6933 0.4087 0.7123 0.3835 0.0711 0.4604
Running time 0.0255 0.2672 0.0760 0.1770 5.3315 0.1939 0.0886
Imbalanced WBDC NMI 0.0966 0.0848 0.0828 0.6846 0.3893 0.3038 0.6907
ARI 0.0337 0.0184 0.0158 0.7855 0.5156 0.3323 0.8308
ACC 0.6260 0.5827 0.5745 0.9892 0.9539 0.9051 0.9892
CV1 0.2644 0.1418 0.1188 1.3529 1.2073 1.0540 1.3069
DCV 1.0578 1.1804 1.2034 -0.0307 0.1150 0.2683 0.0153
Runing time 0.0141 0.0234 0.0094 0.0959 1.2488 0.0870 0.0535
Htru2 NMI 0.4068 0.4101 0.4076 0.0547 0.3854 0.2611 0.5869
ARI 0.6071 0.5798 0.6076 0.0217 0.5967 0.3504 0.7331
ACC 0.9366 0.9253 0.9366 0.9097 0.9428 0.8473 0.9660
CV1 1.0891 1.0125 1.0882 1.4107 1.1813 0.7805 1.2342
DCV 0.3251 0.4017 0.3260 0.0035 0.2329 0.6337 0.1800
Running time 0.0825 0.2913 0.1094 2.0558 68.3254 0.4956 0.1119

### v2

@@
 Abstract
-Imbalanced data, characterized by an unequal distribution of data points across different clusters, poses a challenge for traditional hard and fuzzy clustering algorithms, such as hard K-means (HKM, or Lloyd’s algorithm) and fuzzy K-means (FKM, or Bezdek’s algorithm).
-This paper introduces equilibrium K-means (EKM), a novel and simple K-means-type algorithm that alternates between just two steps, yielding significantly improved clustering results for imbalanced data by reducing the tendency of centroids to crowd together in the center of large clusters.
-We also present a unifying perspective for HKM, FKM, and EKM, showing they are essentially gradient descent algorithms with an explicit relationship to Newton’s method.
-EKM has the same time and space complexity as FKM but offers a clearer physical meaning for its membership definition.
-We illustrate the performance of EKM on two synthetic and ten real datasets, comparing it to various clustering algorithms, including HKM, FKM, maximum-entropy fuzzy clustering, two FKM variations designed for imbalanced data, and the Gaussian mixture model.
-The results demonstrate that EKM performs competitively on balanced data while significantly outperforming other techniques on imbalanced data.
-For high-dimensional data clustering, we demonstrate that a more discriminative representation can be obtained by mapping high-dimensional data via deep neural networks into a low-dimensional, EKM-friendly space.
-Deep clustering with EKM improves clustering accuracy by 35% on an imbalanced dataset derived from MNIST compared to deep clustering based on HKM.
+Traditional centroid-based clustering algorithms, such as hard K-means (HKM, or Lloyd’s algorithm) and fuzzy K-means (FKM, or Bezdek’s algorithm), display degraded performance when true underlying groups of data have varying sizes (i.e., imbalanced data).
+This paper introduces equilibrium K-means (EKM), a novel fuzzy clustering algorithm that has the robustness to imbalanced data by preventing centroids from crowding together in the center of large clusters.
+EKM is simple, alternating between two steps; fast, with the same time and space complexity as FKM; and scalable to large datasets.
+We evaluate the performance of EKM on two synthetic and ten real datasets, comparing it to other centroid-based algorithms, including HKM, FKM, maximum-entropy fuzzy clustering (MEFC), two FKM variations designed for imbalanced data, and the Gaussian mixture model.
+The results show that EKM performs competitively on balanced data and significantly outperforms other algorithms on imbalanced data.
+Deep clustering experiments on the MNIST dataset demonstrate the significance of making representation have an EKM-friendly structure when dealing with imbalanced data; In comparison to deep clustering with HKM, deep clustering with EKM obtains a more discriminative representation and a 35% improvement in clustering accuracy.
+Additionally, we reformulate HKM, FKM, MEFC, and EKM in a general form of gradient descent, where fuzziness is introduced differently and more simply than in Bezdek’s work, and demonstrate how the general form facilitates a uniform study of KM algorithms.
 Index Terms:
-K-means, fuzzy K-means, equilibrium K-means, imbalanced clustering, deep clustering
+K-means, fuzzy clustering, imbalance learning, deep clustering
 I Introduction
-Imbalanced data refers to data that is unevenly distributed across different clusters, which is prevalent in medical diagnosis, fraud detection, and anomaly detection.
-Learning from imbalanced data has been considered a challenge and a valuable direction [1].
-While there is a considerable amount of research on supervised learning from imbalanced data, e.g., [2, 3, 4], unsupervised learning has not been as thoroughly explored, due to the unknown imbalance ratio between clusters making the task more challenging [5].
+Imbalanced data refers to the true underlying groups of data having different sizes, which is common in datasets of medical diagnosis, fraud detection, and anomaly detection.
+Imbalanced data poses a challenge for learning algorithms because these algorithms tend to be biased towards the majority group [1].
+While there is a considerable amount of research on supervised learning (e.g., classification) from imbalanced data [2, 3, 4], unsupervised learning has not been as thoroughly explored, because the unknown cluster sizes make the task more difficult [5].
+Methods like resampling and boosting frequently used in supervised learning cannot be applied in unsupervised learning.
 Clustering is an important unsupervised learning task involving grouping data into clusters based on similarity.
 K-means (KM) is the most popular clustering technique, valued for its simplicity, scalability, and effectiveness with real datasets.
-It is also used as an initialization method for more advanced clustering techniques, such as the Gaussian mixture model [6, 7].
+It can also be used as an initialization method for more advanced clustering techniques, such as the Gaussian mixture model (GMM) [6, 7].
 KM starts with an initial set of centroids (cluster centers) and iteratively refines them to increase cluster compactness.
 The hard KM (HKM, or Lloyd’s algorithm) [8, 9] and fuzzy KM (FKM, or Bezdek’s algorithm) [10] are the two most representative KM algorithms.
-HKM assigns a data point to only one cluster, while FKM allows a data point to belong to multiple clusters with varying degrees of membership.
-In recent, they are also widely employed in deep learning, jointly trained with deep neural networks (DNNs) for clustering high-dimensional data, such as images [11, 12, 13, 14].
-Many FKM variations have been developed to enhance its performance in different aspects.
+HKM assigns a data point to only one cluster, while FKM assigns a data point to multiple clusters with varying degrees of membership.
+Research on HKM and FKM has been ongoing.
 For example, the possibilistic FKM (PFKM) [11] was proposed to address the sensitivity of FKM to noise and outliers.
-FKM-$\sigma$ [16] was proposed to improve FKM’s performance on data points with uneven densities or non-spherical shapes in individual clusters.
-Fuzzy local information K-means (FLICM) [17] was designed to promote FKM’s performance in image segmentation.
-For more information on other FKM variants, please refer to the review paper [18].
+FKM-$\sigma$ [12] was proposed to improve FKM’s performance on data points with uneven variations or non-spherical shapes in individual clusters.
+Fuzzy local information K-means (FLIKM) [13] was designed to promote FKM’s performance in image segmentation.
+Deep clustering, a technique that combines deep neural networks (DNNs) with HKM and FKM, was proposed to cluster high-dimensional data [14, 15, 16, 17].
+Although many variations based on HKM and FKM have been developed to deal with different situations, most of them display degraded effectiveness in cases of imbalanced data.
+This is due to the so-called “uniform effect” that causes the clusters generated by these algorithms to have similar sizes even when input data has highly varying group sizes [18].
+An illustration of the uniform effect of HKM and FKM is given in Fig.
+1 where we can observe that the centroids of HKM and FKM crowd together in the large true cluster.
 Figure 1: Clustering results of a highly imbalanced dataset.
@@
 (d) Clustering by the proposed equilibrium K-means.
-They are all two-step alternating algorithms that iteratively update centroids and clusters.
-Clustering imbalanced data also poses a challenge for both HKM and FKM.
-This is due to an undesirable property known as the “uniform effect”, which causes centroids to crowd together in large clusters, resulting in clusters of equal size even when input data has varied cluster sizes.
-An example illustrating the uniform effect of HKM and FKM is given in Fig.
-1.
+They are all two-step alternating algorithms that iteratively compute centroids.
 I-A Existing Efforts to Overcome Uniform Effect
-To improve the poor performance of FKM on imbalanced data, Noordam et al.
-[19] developed a modified FKM called cluster-size insensitive FKM (csiFKM).
-This method assigns weight to adjust the contribution of data points according to the size of the clusters they belong to.
-Lin et al.
-[20] proposed a size-insensitive integrity-based FKM (siibFKM) to reduce the sensitivity of csiFKM to the distance between adjacent clusters by increasing the cluster’s integrity.
+To the best of our knowledge, there are two popular methods to overcome the uniform effect in the field of fuzzy systems.
+The first method is to introduce more weight on the data points in small clusters at each updating iteration, biasing learning towards them, like a modified FKM called cluster-size insensitive FKM (csiFKM) developed by Noordam et al.
+[19].
+Later, Lin et al.
+[20] proposed a size-insensitive integrity-based FKM (siibFKM) based on csiFKM to reduce the sensitivity of csiFKM to the distance between adjacent clusters.
 However, weighting based on cluster size inadvertently increases the influence of outliers, making these algorithms sensitive to noise [21].
-Another approach to tackle the challenge of imbalanced data clustering is to use a multiprototype mechanism.
-This method first groups data into multiple subclusters, ensuring that the data points in each subcluster are relatively balanced.
-The final clusters are obtained by merging adjacent subclusters.
+The second method is called multiprototype clustering.
+This method first groups data into multiple subclusters with similar sizes and the final clusters are obtained by merging adjacent subclusters.
 Liang et al.
-[21] introduced a multiprototype clustering algorithm that employs FKM to generate subclusters.
+[22] proposed a multiprototype clustering algorithm that employs FKM to generate subclusters.
 Later, Lu et al.
-[5] proposed a self-adaptive multiprototype clustering algorithm that can automatically adjust the number of subclusters.
-However, multiprototype clustering algorithms involve a complex process and high complexity, with a time complexity of $O(N^{2})$, where $N$ is the number of data points in the dataset, making them computationally expensive for large datasets.
+[5] proposed a self-adaptive multiprototype clustering algorithm that automatically adjusts the number of subclusters.
+However, multiprototype clustering algorithms have a complex process and high time complexity of $O(N^{2})$, where $N$ is the number of data points in the dataset.
+Thus, they are computationally expensive for large datasets.
 We should additionally mention that Zeng et al.
 [23] recently proposed a soft multiprototype clustering algorithm with time complexity linear to $N$.
-However, the clustering process of this algorithm remains complex and is aimed at clustering high-dimensional and complex-structured data rather than imbalanced data.
+However, their clustering process remains complex and is aimed at clustering high-dimensional and complex-structured data rather than imbalanced data.
 I-B Our Contributions
-In this paper, we propose equilibrium K-means (EKM) for imbalanced data clustering, which avoids the limitations of the aforementioned techniques.
+In this paper, we propose equilibrium K-means (EKM) to address the issue of imbalanced data clustering and the aforementioned algorithms’ limitations.
 Our contributions can be summarized in three main aspects.
-First, we offer a new framework that unifies HKM, FKM, and EKM with three key advantages.
-1.
-This framework can directly re-derive HKM and FKM from unconstrained optimization problems with differentiable objectives without applying Lagrangian multipliers, simplifying the previous derivation from constrained optimization problems and non-differentiable objectives.
-2.
-The new objective functions of HKM and FKM can be compared in the same plot, which used to be unfeasible due to the old objective function of FKM having more optimization variables than that of HKM.
-Such comparison can facilitate a more comprehensive and fundamental view than existing comparisons (such as [23, 24]) based on clustering outcomes.
-3.
-This new framework allows for the uniform study of KM-type algorithms, such as the convergence conditions provided in this paper.
-We also reveal an explicit relationship between these KM algorithms and Newton’s method.
-Second, we present EKM, derived from the new framework.
-Similar to FKM, EKM is a two-step alternating algorithm between updating the weighted centroid and calculating the weight, but the membership defined in EKM has a clearer physical meaning than that defined in FKM.
-Besides, EKM prevents centroids crowd together by introducing appropriate repulsive forces between centroids, making it more effective than HKM and FKM on imbalanced data.
-In addition to having the same time and space complexity as FKM, EKM also supports batch learning, making it scalable for large datasets.
-We demonstrate EKM’s performance through numerical experiments on two synthetic and ten real datasets from various domains.
-Because of space limitations, quantified metrics of performance are presented in the supplemental material, but the scatter graphs of partial clustering outcomes that we present in the paper are sufficient to demonstrate the superiority of EKM.
-Finally, we investigate the cooperation between EKM and DNNs for high-dimensional data clustering.
-We find that mapping high-dimensional data via DNNs to a low-dimensional EKM-friendly space results in a more discriminative representation than mapping to an HKM-friendly space.
-When applied to an imbalanced dataset derived from MNIST, joint learning of DNNs and EKM improves clustering accuracy by 35% compared to joint learning of DNNs and HKM.
+First, we reformulate HKM, FKM, the maximum-entropy fuzzy clustering (MEFC) [24, 25], and EKM in a general form of gradient descent.
+We show that these algorithms aim to optimize different approximations of the same objective.
+This general form facilitates the uniform study of KM algorithms.
+Second, we develop EKM based on the first contribution.
+EKM belongs to the family of fuzzy clustering and membership defined in EKM has a clear physical meaning.
+Repulsive forces appear among centroids of EKM, successfully reducing the uniform effect by preventing centroids from crowding together in a large cluster (see Fig.
+1d for an example).
+Therefore, EKM is effective in dealing with imbalanced data and is not sensitive to noise.
+Similar to HKM and FKM, EKM is a two-step alternating algorithm that iteratively computes centroids.
+In addition to having the same time ($O(N)$) and space complexity as FKM, EKM has a batch-learning version that can be applied to large datasets.
+We demonstrate the effectiveness of EKM on balanced and imbalanced data by conducting numerical experiments on two synthetic and ten real datasets from different domains.
+Due to space limitations, quantitative performance metrics are provided in the supplementary material.
+The scatter plots of partial clustering results we provide in the paper are sufficient to demonstrate the superiority of EKM on imbalanced data.
+Finally, we investigate the joint learning of DNNs and EKM.
+We find that mapping high-dimensional data via DNNs to an EKM-friendly space can result in more discriminative low-dimensional representation than mapping to an HKM-friendly space.
+When tested on an imbalanced dataset derived from MNIST, joint learning of DNNs and EKM improves clustering accuracy by 35% compared to joint learning of DNNs and HKM.
 I-C Organization
-We introduce the HKM, FKM, and a variation of FKM, the maximum-entropy fuzzy clustering (MEFC) in Section II.
-In Section III and Section IV, we introduce the proposed clustering framework and the proposed EKM.
-We demonstrate the performance of EKM on classic clustering tasks in Section V and deep clustering in Section VI.
+We introduce HKM, FKM, and MEFC in Section II.
+In Section III, we derive their general form and the proposed EKM.
+The properties of EKM are studied in Section IV.
+We evaluate the performance of EKM on classic clustering tasks in Section V and on deep clustering in Section VI.
 Finally, we conclude in Section VII.
@@
 II-A The Hard K-Means Algorithm (Lloyd’s Algorithm)
-KM clustering aims to partition $N$ data points into $K$ clusters, minimizing the sum of the variances within each cluster. Mathematically, the objective can be expressed as:
+KM aims to partition $N$ data points into $K$ clusters, minimizing the sum of the variances within each cluster. Mathematically, the objective can be expressed as:
 $\underset{\mathbb{S}_{1},\cdots,\mathbb{S}_{K}}{\arg\min}\sum_{k=1}^{K}\sum_{\mathbf{x}\in\mathbb{S}_{k}}\|\mathbf{x}-\bm{\mu}_{k}\|_{2}^{2},$ (1)
-where $\mathbb{S}_{k}$ represents the set of data points in the $k$-th cluster, $\mathbf{x}\in\mathbb{S}_{k}$ denotes a data point belonging to $\mathbb{S}_{k}$, $\bm{\mu}_{k}$ is the centroid of points in $\mathbb{S}_{k}$, expressed by
+where $\mathbb{S}_{k}$ represents the set of data points in the $k$-th cluster, $\mathbf{x}\in\mathbb{S}_{k}$ denotes a data point belonging to $\mathbb{S}_{k}$, $\bm{\mu}_{k}$ is the centroid of $\mathbb{S}_{k}$, expressed by
 $\bm{\mu}_{k}=\frac{1}{|\mathbb{S}_{k}|}\sum_{\mathbf{x}\in\mathbb{S}_{k}}\mathbf{x},$ (2)
 $|\mathbb{S}_{k}|$ signifies the size of $\mathbb{S}_{k}$, and $\|\cdot\|_{2}$ is the $l_{2}$ norm.
-The optimization problem (1) is NP-hard [25], and Lloyd‘s algorithm is a popular heuristic algorithm for solving this problem.
+This optimization problem is NP-hard [26] and is commonly solved heuristically by Lloyd‘s algorithm [9].
 Given $K$ points ($\mathbf{c}_{1}^{(1)},\cdots,\mathbf{c}_{K}^{(1)})$ as initial centroids, Lloyd’s algorithm alternates between two steps:
@@
 II-B The Uniform Effect of K-Means
-The uniform effect refers to the propensity to generate clusters of equal size, which is implicitly implied in the objective (1) of KM.
+The uniform effect refers to the propensity to generate clusters of similar sizes, which is implicitly implied in the objective (1) of KM.
 For simplicity, kindly consider a case with two clusters (i.e., $K=2$).
@@
 FKM attempts to minimize the sum of weighted distances between data points and centroids, with the following objective and constraints [10]:
-$\min_{\mathbf{c}_{1},\cdots,\mathbf{c}_{K},\{u_{kn}\}}\quad$ $\sum_{n=1}^{N}\sum_{k=1}^{K}(u_{kn})^{m}\|\mathbf{x}_{n}-\mathbf{c}_{k}\|_{2}^{2}$ (6)
+$\min_{\mathbf{c}_{1},\cdots,\mathbf{c}_{K},\{u_{kn}\}}$ $\sum_{n=1}^{N}\sum_{k=1}^{K}(u_{kn})^{m}\|\mathbf{x}_{n}-\mathbf{c}_{k}\|_{2}^{2}$ (6)
 […]
-$\sum_{k}u_{kn}=1\,\forall n,$
-$0<\sum_{n}u_{kn}<N\,\forall k,$
+$\sum_{k=1}^{K}u_{kn}=1\,\forall n,$
+$0<\sum_{n=1}^{N}u_{kn}<N\,\forall k,$
 where $\mathbf{x}_{n}$ represents the $n$-th data point, $\mathbf{c}_{k}$ denotes the $k$-th centroid, $u_{kn}$ is a coefficient called membership that indicates the degree of $\mathbf{x}_{n}$ belonging to the $k$-th cluster, and $m\in(1,+\infty)$ is a hyperparameter controlling the degree of fuzziness level.
@@
   1. 1.
-    Calculate the membership values of data points of being in the cluster:
+    Calculate the membership value of the $n$-th data point belonging to the $k$-th cluster:
     $u_{kn}^{(\tau)}=\frac{1}{\sum_{i=1}^{K}\bigg{(}\frac{\|\mathbf{x}_{n}-\mathbf{c}_{k}^{(\tau)}\|_{2}}{\|\mathbf{x}_{n}-\mathbf{c}_{i}^{(\tau)}\|_{2}}\bigg{)}^{\frac{2}{m-1}}}.$ (7)
   2. 2.
-    Recalculate the centroid of each cluster by taking the mean of all data points, weighted by their membership to the cluster:
+    Recalculate the weighted centroid of the $k$-th cluster by:
     $\mathbf{c}_{k}^{(\tau+1)}=\frac{\sum_{n}(u_{kn}^{(\tau)})^{m}\mathbf{x}_{n}}{\sum_{n}(u_{kn}^{(\tau)})^{m}}.$ (8)
-The time complexity of one iteration of the above two steps is $O(NK^{2})$.
-The higher time complexity of FKM than that of HKM is due to the extra calculation of membership.
-A practical time consumption comparison can be found in [27].
+The time complexity of one iteration of the above two steps is $O(NK^{2})$. The higher time complexity of FKM than HKM is due to the extra membership calculation.
 II-D Maximum-Entropy Fuzzy Clustering
 Karayiannis [24] added an entropy term to the objective function of FKM, resulting in MEFC. The new objective function is given as follows:
-$\min_{\mathbf{c}_{1},\cdots,\mathbf{c}_{K},\{u_{kn}\}}\quad$ $\eta\sum_{n=1}^{N}\sum_{k=1}^{K}u_{kn}\ln{u_{kn}}$ (9)
+$\min_{\mathbf{c}_{1},\cdots,\mathbf{c}_{K},\{u_{kn}\}}$ $\eta\sum_{n=1}^{N}\sum_{k=1}^{K}u_{kn}\ln{u_{kn}}$ (9)
 $+(1-\eta)\frac{1}{N}\sum_{n=1}^{N}\sum_{k=1}^{K}u_{kn}\|\mathbf{x}_{n}-\mathbf{c}_{k}\|_{2}^{2}$
 […]
-$\sum_{k}u_{kn}=1\,\forall n,$
+$\sum_{k=1}^{K}u_{kn}=1\,\forall n,$
 where $\eta\in(0,1)$ is a hyperparameter, controlling the transition from maximization of the entropy to the minimization of centroid-data distances.
-MEFC is similar to FKM but with a different membership:
-$u^{(\tau)}_{kn}=\frac{\exp{(-\lambda\|\mathbf{x}_{n}-\mathbf{c}_{k}\|_{2}^{2})}}{\sum_{i=1}^{K}\exp{(-\lambda\|\mathbf{x}_{n}-\mathbf{c}_{i}\|_{2}^{2})}},$ (10)
+MEFC is similar to FKM but with a different definition of membership:
+$u^{(\tau)}_{kn}=\frac{\exp{(-\lambda\|\mathbf{x}_{n}-\mathbf{c}_{k}^{(\tau)}\|_{2}^{2})}}{\sum_{i=1}^{K}\exp{(-\lambda\|\mathbf{x}_{n}-\mathbf{c}_{i}^{(\tau)}\|_{2}^{2})}},$ (10)
 where $\lambda=\frac{1}{N}\frac{1-\eta}{\eta}$. The time complexity of MEFC is the same as FKM, which is $O(NK^{2})$ for one iteration.
@@
 III-A Objective of Smooth K-Means
-Here, we propose a novel framework called smooth K-means (SKM) and demonstrate that the three KM algorithms introduced in the previous section are special cases of SKM.
-Based on this framework, we further develop a new clustering algorithm, aiming to improve the degraded performance of KM algorithms on imbalanced data.
+We propose a novel framework called smooth K-means (SKM) and demonstrate that the three KM algorithms introduced in the previous section are special cases of SKM.
 Denote the squared Euclidean distance [^1] Other distance metrics, such as the absolute difference and the angle between points, can also be used to define $d_{kn}$.
-Considering the derivation process remains the same, we use the Euclidean distance in the paper for simplicity.
+Considering the derivation process is the same, we use the Euclidean distance in the paper for simplicity.
 between the $k$-th centroid and the $n$-th data point as
 $d_{kn}=\frac{1}{2}\|\mathbf{x}_{n}-\mathbf{c}_{k}\|_{2}^{2},$ (11)
-and define the within-cluster sum of squares (WCSS) as the sum of squared Euclidean distances between data points and their nearest centroid, i.e.,
+and define the within-cluster sum of squares (WCSS) as the sum of squared Euclidean distances between data points and their nearest centroids, i.e.,
 $\text{WCSS}:=\sum_{n=1}^{N}\min(d_{1n},\cdots,d_{Kn}).$ (12)
@@
 where $h(d_{1n},\cdots,d_{Kn})$ is a smooth approximation to $\min(d_{1n},\cdots,d_{Kn})$ and is referred to as the smooth minimum function.
-Below we explain the significance of replacing the minimum function with a smooth minimum function from a statistical point of view.
-III-B Statistical Interpretation of Smooth K-Means
-Statistically, the centroids obtained by minimizing WCSS are maximum likelihood estimators (MLEs) of parameters in a specific type of Gaussian mixture model (GMM) referred to as “hard” GMM.
-Assuming that the dataset $(\mathbf{x}_{1},\cdots,\mathbf{x}_{N})$ is generated by $K$ independent multivariate normal distributions, we denote the mean vector and covariance matrix of the $k$-th normal distribution as $\bm{\mu}_{k}$ and $\mathbf{\Sigma}_{k}$, respectively.
-The standard GMM assumes that data points are generated from multiple normal distributions based on a certain probability distribution, with the following likelihood
+Below we reveal the relationship between the SKM and the GMM.
+III-B The Relationship Between SKM and GMM
+The centroids obtained by minimizing WCSS are maximum likelihood estimators (MLEs) of parameters of “hard” GMM.
+The derivation is as follows.
+Assuming that the dataset $(\mathbf{x}_{1},\cdots,\mathbf{x}_{N})$ is sampled from $K$ independent multivariate normal distributions.
+We denote the mean vector and covariance matrix of the $k$-th normal distribution as $\bm{\mu}_{k}$ and $\mathbf{\Sigma}_{k}$, respectively.
+The standard GMM has the following well-known likelihood
 $L(\bm{\mu},\mathbf{\Sigma},p\,|\,\mathbf{x})\propto$ $\prod_{n=1}^{N}\sum_{k=1}^{K}p_{kn}\det(\mathbf{\Sigma}_{k})^{-1/2}$ (14)
@@
 where $p_{kn}$ is the probability that the $n$-th data point is generated from the $k$-th normal distribution.
-The hard GMM, on the other hand, assumes that data points can only be generated from a single normal distribution.
+The standard GMM assumes that data points are generated from multiple normal distributions based on a certain probability distribution.
+On the other hand, the hard GMM assumes that a data point is generated from a single normal distribution.
 Accordingly, for all $n$, $p_{kn}=1$ for one $k\in\{1,\cdots,N\}$ and $p_{in}=0$ if $i\neq k$.
@@
 Now it is clear that the MLEs of $\{\bm{\mu}_{1},\cdots,\bm{\mu}_{K}\}$ are the centroids minimizing WCSS.
-While the hard GMM model is more computationally efficient than the standard GMM, it only considers the impact of a data point on its closest centroid.
-However, as seen in the success of FKM, it can be advantageous to consider the effect of data points on all centroids for certain real datasets.
-This can be accomplished by SKM in a parameter-free manner, which applies a smooth function to WCSS.
-Different smoothing methods result in unique impact patterns, allowing for the capture of various data structures.
-In the following sections, we will explore three popular smooth minimum functions and their impact on clustering.
+The hard GMM model simplifies the standard GMM by only considering the impact of a data point on its closest centroid.
+However, as seen in the success of FKM, it can be more advantageous to consider the impact of a data point on all centroids.
+This can be accomplished in SKM by applying an approximation function to smooth WCSS.
+Hence, SKM can be viewed as a model between hard and standard GMM.
+In the following sections, we will introduce three common smooth minimum functions and explore the resulting distinct clustering algorithms.
 III-C Three Common Smooth Minimum Functions
@@
 where $f^{\prime}$ is the first derivative of $f$. Let $g(x)=1/f(x)$, a smooth minimum function $h_{1}:[0,\,+\infty)^{K}\mapsto[0,+\infty)$ can be constructed by
-$h_{1}(x_{1},\cdots,x_{K})=g^{-1}(g(x_{1})+\cdots+g(x_{K}))$ (18)
+$h_{1}(x_{1},\cdots,x_{K})=g^{-1}(g(x_{1})+\cdots+g(x_{K})).$ (18)
 Let $f(x)$ be defined as $e^{\lambda x}$, then the function $h_{1}$ takes on a specific form known as LogSumExp:
@@
 $=-\frac{1}{\lambda}\ln(e^{-\lambda x_{1}}+\cdots+e^{-\lambda x_{K}}),$
-where $\lambda$ serves as a parameter that controls the degree of approximation, with $\text{LSE}_{\lambda}\to\min$ as $\lambda\to+\infty$.
-Another common smooth minimum function, called p-Norm, can be defined as $f(x)=x^{p}$, which has the following specific form
+where $\lambda$ is a parameter controlling the degree of approximation, with $\text{LSE}_{\lambda}\to\min$ as $\lambda\to+\infty$.
+Define $f(x)=x^{p}$, we can have another common smooth minimum function, called p-Norm, which has the following specific form
 $h_{1}(x_{1},\cdots,x_{K})$ $=\text{PN}_{p}(x_{1},\cdots,x_{K})$ (20)
@@
 and converges to $\min(x_{1},\cdots,x_{K})$ as $p\to+\infty$.
-Smooth minimum functions can also be constructed using
+An additional definition is necessary for the domain of the p-Norm function to legally include zeros: $\text{PN}_{p}(x_{1},\cdots,x_{K})=0$ if some $x_{i}=0$ where $1\leq i\leq K$.
+Smooth minimum functions can also be constructed by
 $h_{2}(x_{1},\cdots,x_{K})=\frac{x_{1}g(x_{1})+\cdots+x_{K}g(x_{K})}{g(x_{1})+\cdots+g(x_{K})}.$ (21)
-A specific example is the Boltzmann operator, which is defined as $g(x)=e^{-\alpha x}$. The Boltzmann operator takes on the form of
+A specific example is the Boltzmann operator, where $g(x)=e^{-\alpha x}$. The Boltzmann operator takes on the form of
 $h_{2}(x_{1},\cdots,x_{K})$ $=\text{boltz}_{\alpha}(x_{1},\cdots,x_{K})$ (22)
@@
 $\mathbf{c}_{k}^{(\tau+1)}=\mathbf{c}_{k}^{(\tau)}-\gamma^{(\tau)}_{k}\partial_{\mathbf{c}_{k}}{J_{\text{LSE}}}(\mathbf{c}_{k}^{(\tau)}),$ (25)
-where $\gamma^{(\tau)}_{k}$ is the learning rate at the $\tau$-th iteration. This updating procedure (25) is equivalent to the MEFC algorithm if one set
+where $\gamma^{(\tau)}_{k}$ is the learning rate at the $\tau$-th iteration.
+This updating procedure (25) is equivalent [^2] It should be noted that the equivalence mentioned in this paper is in the sense of the algorithm level, not in the criterion level.
+to the MEFC algorithm if one set
 $\gamma^{(\tau)}_{k}=1/\sum_{n=1}^{N}\bigg{(}\frac{e^{-\lambda d_{kn}^{(\tau)}}}{\sum_{i=1}^{K}e^{-\lambda d_{in}^{(\tau)}}}\bigg{)}.$ (26)
 Such learning rate value is related to the second-order partial derivative of $J_{\text{LSE}}$, which we will discuss later.
+The membership (10) of MEFC is identical to $\partial J_{\text{LSE}}/\partial d_{kn}$.
 III-D2 Towards Lloyd’s algorithm
-In the limit $\lambda\to\infty$, $\frac{e^{-\lambda d_{kn}}}{\sum_{i}e^{-\lambda d_{in}}}\to 1$ if $d_{kn}\leq d_{in}\,\forall i\in\{1,\cdots,K\}$, and $\frac{e^{-\lambda d_{kn}}}{\sum_{i}e^{-\lambda d_{in}}}\to 0$ otherwise [^2] We assume $\forall n$, if $k\neq i$, then $d_{kn}\neq d_{in}$..
+In the limit of $\lambda\to\infty$, $\frac{e^{-\lambda d_{kn}}}{\sum_{i}e^{-\lambda d_{in}}}\to 1$ if $d_{kn}\leq d_{in}\,\forall i\in\{1,\cdots,K\}$, and $\frac{e^{-\lambda d_{kn}}}{\sum_{i}e^{-\lambda d_{in}}}\to 0$ otherwise [^3] We assume $\forall n$, if $k\neq i$, then $d_{kn}\neq d_{in}$..
 In this case, the learning rate $\gamma_{k}^{(\tau)}\to 1/N_{k}^{(\tau)}$ where $N_{k}^{(\tau)}$ is the number of data points closest to $\mathbf{c}_{k}^{(\tau)}$, and the updating procedure (25) approaches to Lloyd’s algorithm.
-III-D3 Lloyd’s algorithm as a Newton’s method
+III-D3 Lloyd’s algorithm and Newton’s method
 The smooth objective function $J_{\text{LSE}}$ possesses the second-order partial derivative given by
@@
 $\mathbf{c}_{k}^{(\tau+1)}=\mathbf{c}_{k}^{(\tau)}-[\partial_{\mathbf{c}_{k}}^{2}{J_{\text{LSE}}}(\mathbf{c}_{k}^{(\tau)})]^{-1}\partial_{\mathbf{c}_{k}}{J_{\text{LSE}}}(\mathbf{c}_{k}^{(\tau)}).$ (28)
-As $\lambda\to+\infty$, $\partial_{\mathbf{c}_{k}}^{2}{J_{\text{LSE}}(\mathbf{c}_{k}^{(\tau)})}\to N_{k}^{(\tau)}\mathbf{I}$, and the updating procedure (28) asymptotically approaches Lloyd’s algorithm.
+As $\lambda\to+\infty$, $\partial_{\mathbf{c}_{k}}^{2}{J_{\text{LSE}}(\mathbf{c}_{k}^{(\tau)})}\to N_{k}^{(\tau)}\mathbf{I}$, and the updating procedure (28) approaches Lloyd’s algorithm.
 This indicates that Lloyd’s algorithm is essentially Newton’s method, which aligns with the perspective presented in [28].
-There is a close relationship between the learning rate value (26) and the second-order partial derivative of $J_{\text{LSE}}$ (27): the gradient descent with the learning rate of  (26) is equivalent to Newton’s method considering only a portion of the Hessian matrix (ignoring a rank-1 matrix $\mathbf{D}_{kn}$).
-Consequently, MEFC can be viewed as an approximated Newton’s method.
+There is a close relationship between the learning rate value (26) and the second-order partial derivative of $J_{\text{LSE}}$ (27): the gradient descent with the learning rate of  (26) is equivalent to an approximated Newton’s method in the sense that one term (the rank-1 matrix $\mathbf{D}_{kn}$) in (27) is ignored.
 A discussion of the pros and cons of optimizing $J_{\text{LSE}}$ using gradient descent or Newton’s method would be valuable, but it is beyond the scope of this paper.
-In the following section, we will observe a similar connection between Newton’s method and FKM.
+In the following section, we will observe the same connection between Newton’s method and FKM.
 III-D4 The relationship between p-Norm and fuzzy K-means
@@
 $\min_{\mathbf{c}_{1},\cdots,\mathbf{c}_{K}}J_{\text{PN}}(\mathbf{c}_{1},\cdots,\mathbf{c}_{K})=\sum_{n=1}^{N}(d_{1n}^{-p}+\cdots+d_{Kn}^{-p})^{-1/p}.$ (29)
-The objective $J_{\text{PN}}$ has the first partial derivative as
+The objective $J_{\text{PN}}$ has the first partial derivative of
 $\partial_{\mathbf{c}_{k}}{J_{\text{PN}}}=\frac{\partial{J_{\text{PN}}}}{\partial{\mathbf{c}_{k}}}=-\sum_{n=1}^{N}\frac{d_{kn}^{-p-1}}{(\sum_{i=1}^{K}d_{in}^{-p})^{1/p+1}}(\mathbf{x}_{n}-\mathbf{c}_{k}).$ (30)
@@
 $\gamma^{(\tau)}_{k}=1/\sum_{n=1}^{N}\frac{(d_{kn}^{(\tau)})^{-p-1}}{(\sum_{i=1}^{K}(d_{in}^{(\tau)})^{-p})^{1/p+1}},$ (32)
-this gradient descent optimization is identical to the FKM algorithm.
-The connection between FKM and Newton’s method becomes evident through the second-order partial derivative of $J_{\text{PN}}$, which is
+this gradient descent iteration is equivalent to the FKM algorithm. The membership (7) of FKM is equivalent to a power exponent ($1/m$) of $\partial J_{\text{PN}}/\partial d_{kn}$.
+The connection between FKM and Newton’s method becomes evident by taking the second-order partial derivative of $J_{\text{PN}}$, which is
 $\partial^{2}_{\mathbf{c}_{k}}{J_{\text{PN}}}=\frac{\partial^{2}{J_{\text{PN}}}}{\partial{\mathbf{c}_{k}}^{2}}=\sum_{n=1}^{N}\frac{d_{kn}^{-p-1}}{(\sum_{i}d_{in}^{-p})^{1/p+1}}(\mathbf{I}+\zeta_{kn}\mathbf{D}_{kn}),$ (33)
-where $\zeta_{kn}=(p+1)\frac{\sum_{i\neq k}d_{in}^{-p}}{d_{kn}\sum_{i}d_{in}^{-p}}$. Hence, FKM can also be viewed as an approximated Newton’s method that ignores the rank-1 matrix $\mathbf{D}_{kn}$.
+where $\zeta_{kn}=(p+1)\frac{\sum_{i\neq k}d_{in}^{-p}}{d_{kn}\sum_{i}d_{in}^{-p}}$.
+Hence, FKM can also be viewed as an approximated Newton’s method that ignores the term of the rank-1 matrix $\mathbf{D}_{kn}$.
 III-E From Boltzmann Operator to A Novel Clustering Algorithm
-Employing the Boltzmann operator to smooth WCSS results in a new objective function, which is
-$\min_{\mathbf{c}_{1},\cdots,\mathbf{c}_{K}}J_{\text{B}}(\mathbf{c}_{1},\cdots,\mathbf{c}_{K})=\sum_{n=1}^{N}\frac{\sum_{i}d_{in}e^{-\alpha d_{in}}}{\sum_{i}e^{-\alpha d_{in}}}.$ (34)
+Employing the Boltzmann operator to smooth WCSS results in
+$\min_{\mathbf{c}_{1},\cdots,\mathbf{c}_{K}}J_{\text{B}}(\mathbf{c}_{1},\cdots,\mathbf{c}_{K})=\sum_{n=1}^{N}\frac{\sum_{i=1}^{K}d_{in}e^{-\alpha d_{in}}}{\sum_{i=1}^{K}e^{-\alpha d_{in}}}.$ (34)
 The objective $J_{B}$ possesses the first-order partial derivative of
@@
 $\gamma^{(\tau)}_{k}=1/\bigl{(}\sum_{n=1}^{N}\frac{e^{-\alpha d_{kn}^{(\tau)}}}{\sum_{i}e^{-\alpha d_{in}^{(\tau)}}}[1-\alpha(d_{kn}^{(\tau)}-\frac{\sum_{i}d_{in}^{(\tau)}e^{-\alpha d_{in}^{(\tau)}}}{\sum_{i}e^{-\alpha d_{in}^{(\tau)}}})]\bigl{)}.$ (37)
-This updating procedure can be reformulated into a two-step iteration procedure, akin to FKM, which entails finding centroids using
+This updating procedure can be reformulated into a two-step iteration procedure akin to FKM, that is
 $\mathbf{c}_{k}^{(\tau+1)}=\frac{\sum_{n}w_{kn}^{(\tau)}\mathbf{x}_{n}}{\sum_{n}w_{kn}^{(\tau)}},$ (38)
-where weights $w_{kn}^{(\tau)}$ are given by
-$w_{kn}^{(\tau)}=\frac{e^{-\alpha d_{kn}^{(\tau)}}}{\sum_{i}e^{-\alpha d_{in}^{(\tau)}}}[1-\alpha(d_{kn}^{(\tau)}-\frac{\sum_{i}d_{in}^{(\tau)}e^{-\alpha d_{in}^{(\tau)}}}{\sum_{i}e^{-\alpha d_{in}^{(\tau)}}})].$ (39)
-The time complexity of one iteration is $O(NK^{2})$. Although the sum of weights for each instance equals one:
-$\sum_{k=1}^{K}w_{kn}^{(\tau)}=1,$ (40)
-it is worth noting that these weights cannot be interpreted as probabilities nor memberships, since some values are negative. Section III-G will discuss how to define membership in EKM.
+where weights $w_{kn}^{(\tau)}$ are calculated by
+$w_{kn}^{(\tau)}=\frac{e^{-\alpha d_{kn}^{(\tau)}}}{\sum_{i=1}^{K}e^{-\alpha d_{in}^{(\tau)}}}[1-\alpha(d_{kn}^{(\tau)}-\frac{\sum_{i=1}^{K}d_{in}^{(\tau)}e^{-\alpha d_{in}^{(\tau)}}}{\sum_{i=1}^{K}e^{-\alpha d_{in}^{(\tau)}}})].$ (39)
+The time complexity of one iteration is $O(NK^{2})$. Although the sum of weights equals one, i.e.,
+$\sum_{k=1}^{K}w_{kn}^{(\tau)}=1,\,\forall n,$ (40)
+it is worth noting that these weights cannot be interpreted as probabilities or memberships since some weight values are negative. The membership of EKM is defined in Section III-G.
 III-F Physical Interpretation of Equilibrium K-Means
 The second law of thermodynamics asserts that in a closed system with constant external parameters (e.g., volume) and fixed entropy, the internal energy will reach its minimum value at the state of thermal equilibrium.
-The objective (34) of EKM adheres to this minimum energy principle.
-This connection can be established by envisioning data points as particles with discrete/quantized energy levels, where the energy value corresponds to the squared Euclidean distance between a data point and a centroid, and the number of energy levels is equivalent to the number of centroids.
-According to Boltzmann’s law, at the state of thermodynamic equilibrium, the probability of a particle occupying a specific energy level decreases exponentially with the energy value of that level.
-Hence, the objective function (34) is equivalent to the expectation of energy for the entire system, and the proposed clustering algorithm seeks centroids that minimize this expectation of energy.
+The objective (34) of EKM follows this minimum energy principle.
+This connection can be established by envisioning data points as particles with discrete/quantized energy levels, where the number of energy levels is equivalent to the number of centroids, and the energy value corresponds to the squared Euclidean distance between a data point and a centroid.
+Boltzmann’s law tells that at the state of thermodynamic equilibrium, the probability of a particle occupying a specific energy level decreases exponentially with the increase of the energy value of that level.
+Hence, the objective function (34) equals the expectation of the entire system’s energy, and EKM seeks centroids to minimize this energy expectation.
 Due to this connection, we refer to the proposed algorithm (alternating between (38) and (39)) as equilibrium K-means.
 III-G Membership Defined in Equilibrium K-Means
-According to the physical interpretation of EKM, the exponential term $e^{-\alpha d_{kn}}$ can be regarded as the unnormalized probability of the $n$-th data point belonging to the $k$-th clusters.
-Hence, the membership of the $n$-th data point to the $k$-th clusters in EKM can be defined as
+According to the physical interpretation of EKM, the exponential term $e^{-\alpha d_{kn}}$ can be interpreted as the unnormalized probability of the $n$-th data point belonging to the $k$-th clusters.
+Hence, the membership of the $n$-th data point to the $k$-th cluster can be defined as
 $u_{kn}=\frac{e^{-\alpha d_{kn}}}{\sum_{i=1}^{K}e^{-\alpha d_{in}}}.$ (41)
-Note that, although the membership formula of EKM is the same as that (10) of MEFC, the membership values obtained from EKM and MEFC are typically different because their centroids are calculated using distinct formulas.
-This paper does not intend to compare the memberships in different fuzzy clustering algorithms; therefore, we only define membership in EKM without further discussion.
+Note that, although the membership formula of EKM is the same as that (10) of MEFC, the values are different because centroids are calculated using distinct formulas.
+We have no intention in this paper to compare membership defined in different fuzzy clustering algorithms; therefore, we only define membership in EKM and leave the further discussion to subsequent research.
 III-H Convergence of Smooth K-Means
-HKM, FKM, MEFC, and EKM can be viewed as special cases of SKM, which can be generalized as the following gradient descent algorithm:
+HKM, FKM, MEFC, and EKM are special cases of SKM, which can be generalized as the following gradient descent algorithm:
 $\mathbf{c}_{k}^{(\tau+1)}=\mathbf{c}_{k}^{(\tau)}-\gamma^{(\tau)}_{k}\partial_{\mathbf{c}_{k}}J(\mathbf{c}_{1}^{(\tau)},\cdots,\mathbf{c}_{K}^{(\tau)}),$ (42)
@@
 $\gamma^{(\tau)}_{k}=1/\sum_{n=1}^{N}\bigg{(}\frac{\partial h}{\partial d_{kn}}(d_{1n}^{(\tau)},\cdots,d_{Kn}^{(\tau)})\bigg{)}.$ (43)
-The convergence condition of this gradient descent algorithm is given as follows:
+Different KM algorithms can be obtained by taking $h$ the corresponding explicit form (refer to Section III-D to III-E).
+This general form facilitates the uniform study of KM algorithms.
+Below we give a convergence guarantee conditioning on the properties of $h$:
 Theorem 1 (Convergence Condition):
@@
 See the Appendix for the proof, which generalizes the proof of convergence of fuzzy K-means in [29]. ∎
-It can be easily verified that when the smooth minimum function $h$ is LogSumExp (19) or p-Norm (20), the above convergence condition can be satisfied with any initial centroids.
+It can be easily verified that when the smooth minimum function $h$ is LogSumExp (19) and p-Norm (20) in which case SKM (42) is equivalent to MEFC and FKM, respectively, the above convergence condition can be satisfied with any initial centroids.
 Although the Boltzmann operator (22) is not concave, it is a smooth approximation of a concave function (the minimum function).
-Hence, EKM also demonstrates good convergence behavior, which has been empirically verified in our experiments.
+Hence, EKM also exhibits good convergence behavior in numerical experiments.
 IV Comparison of Different Smoothed Objectives
 IV-A Case Study
-This section explores and compares the behavior of different KM algorithms by examining their new objective functions in some examples.
-We select these examples because their data structures are well-designed, making it easier to illustrate the clustering characteristics.
-We generate one-dimensional datasets comprising two classes of data points by sampling from two normal distributions, drawing $N_{1}$ samples from a distribution with mean $\mu_{1}$ and unit variance, and $N_{2}$ samples from another with mean $\mu_{2}$ and unit variance.
-Using different parameter combinations, we create four datasets: 1.
+This section presents an empirical analysis of the behavior of different KM algorithms by examining their reformulated objective functions in some examples with well-designed data structures.
+Datasets comprising two classes of one-dimensional data points are generated by sampling from two normal distributions, drawing $N_{1}$ samples from a distribution with mean $\mu_{1}$ and unit variance, and $N_{2}$ samples from another with mean $\mu_{2}$ and unit variance.
+Using different parameter combinations, we generate four datasets: 1.
 A balanced, non-overlapping dataset ($N_{1}=N_{2}=50$, $\mu_{1}=-5$, and $\mu_{2}=+5$; Fig.
@@
 2c); 4.
-An imbalanced, compact dataset ($N_{1}=2000$, $N_{2}=50$, $\mu_{1}=-2$, and $\mu_{2}=+2$; Fig.
+An imbalanced, overlapping dataset ($N_{1}=2000$, $N_{2}=50$, $\mu_{1}=-2$, and $\mu_{2}=+2$; Fig.
 2d).
-We fix the first centroid at $\mu_{1}$, and plot the objectives of HKM (WCSS (12)), FKM ($J_{\text{PN}}$ (29)), MEFC ($J_{\text{LSE}}$ (23)), and EKM ($J_{\text{B}}$ (34)) as a function of the position of the second centroid in Fig.
+We fix the first centroid at $\mu_{1}$, and plot the reformulated objectives of HKM (WCSS (12)), FKM ($J_{\text{PN}}$ (29)), MEFC ($J_{\text{LSE}}$ (23)), and EKM ($J_{\text{B}}$ (34)) as a function of the position of the second centroid in Fig.
 2.
+The four objective functions behave similarly on the first two balanced datasets, but it is worth noting that the last two imbalanced datasets.
 Fig.
-2a demonstrates the similarity in objective functions for the four clustering algorithms, indicating comparable performance on a balanced dataset.
-Fig.
-2b reveals that FKM and MEFC are better suited for overlapped data, as they are less prone to a local minimum.
-In Fig.
 2c and Fig.
-2d, data imbalance cases show that the local and global minimum points of the objective functions of HKM, FKM, and MEFC are biased towards the center of the large class, i.e., $\mu_{1}$.
-In contrast, EKM does not have an obvious local minimum and its global minimum point aligns with the correct cluster center, i.e., $\mu_{2}$, highlighting its superiority in handling imbalanced data.
-Figure 2: Objectives of four KM algorithms as a function of the first centroid fixed at $\mu_{1}$ and the second centroid moving along the $x$-axis.
-The yellow, purple, green, and blue curves are the objective functions of HKM, FKM, MEFC, and EKM, respectively.
-(a), (b), (c), and (d) present objective functions for different data distributions.
+2d show that the local and global minimum points of the objective functions of HKM, FKM, and MEFC are biased towards the center of the large cluster, i.e., $\mu_{1}$.
+In contrast, EKM does not have an obvious local minimum and its global minimum point aligns with the true cluster center, i.e., $\mu_{2}$, highlighting EKM’s superiority in handling imbalanced data.
+Figure 2: Objectives as a function of the second centroid moving along the $x$-axis (the first centroid is fixed at $\mu_{1}$).
+The yellow, purple, green, and blue curves are reformulated objective functions of HKM, FKM, MEFC, and EKM, respectively.
+(a), (b), (c), and (d) present objective functions under different data distributions.
 IV-B The Analysis of EKM’s Effectiveness on Imbalanced Data
-We demonstrate the effectiveness of EKM by analyzing the gradient of its objective.
-The gradient of the smoothed WCSS with respect to the distance between data points and centroids can be interpreted as the force exerted by a spring.
-Positive gradient values represent attractive forces, while negative values represent repulsive forces.
-The gradient is also associated with membership, where the membership (7) of FKM is equivalent to a power exponent ($1/m$) of $\partial J_{\text{PN}}/\partial d_{kn}$, and the membership (10) of MEFC is equivalent to $\partial J_{\text{LSE}}/\partial d_{kn}$.
-A simple example is given to plot the gradient values of different smoothed objectives.
+We analyze the effectiveness of EKM on imbalanced data based on the gradient of its objective.
+The gradient of the smoothed WCSS with respect to the data-centroid distance (i.e., $\partial J/\partial d_{kn}$) can be interpreted as the force exerted by a spring.
+A positive gradient value represents an attractive force, while a negative value represents a repulsive force.
+A simple example is given to plot the gradient values of different KM’s objectives.
 We fix two centroids at $-1$ and $+1$, respectively, and move a data point along the $x$-axis.
@@
 As evident from the figure, data points on the side of the first centroid do not impact the second centroid of HKM, but they do attract the second centroids of FKM and MEFC.
-This finding supports the previous result in [31] that FKM has a stronger uniform effect than HKM.
-However, for EKM, data points near the first centroid have repulsive forces on the second centroid, which compensates for the attraction of points on the side of the first centroid to the second centroid, effectively reducing the uniform effect.
-Furthermore, it is worth noting that for EKM, data points near the second centroid have the strongest attraction forces.
-Thus, the second centroid is anchored by these data points, making its position less susceptible to noise or outliers.
-This is the advantage of EKM over imbalanced data clustering algorithms that assign weights to data points based on cluster size.
-Figure 3: Gradients of four KM algorithms as a function of a data point moving along the $x$-axis.
-(a) Gradients of HKM, FKM, MEFC, and EKM with fixed smoothness parameters.
-(b) Gradients of FKM.
-(c) Gradients of MEFC.
-(d) Gradients of EKM.
+This finding supports the statement in [30] that FKM has a stronger uniform effect than HKM.
+On the other hand, data points near the first centroid have repulsive forces on the second centroid of EKM, which compensate for the attraction from other data points, effectively reducing the uniform effect.
+It is worth noting that data points near the second centroid of EKM have the strongest attraction forces to this centroid.
+Consequently, EKM’s centroids are anchored by their surrounding data points, making their position less influenced by noise and outliers.
+Figure 3: Gradients as a function of a data point moving along the $x$-axis.
+(a) Gradients of WCSS, $J_{\text{PN}}$, $J_{\text{LSE}}$, and $J_{\text{B}}$ with fixed smoothing parameters.
+(b), (c), and (d) are gradients of $J_{\text{PN}}$, $J_{\text{LSE}}$, and $J_{\text{B}}$ with varied smoothing parameters, respectively.
 IV-C The Choice of $\alpha$
-The smoothing parameter $\alpha$ has an impact on the clustering result in EKM, but the optimal choice remains unknown, which is not a difficulty unique to EKM.
+The smoothing parameter $\alpha$ impacts the performance of EKM, but the optimal choice remains unknown.
+This is not a difficulty unique to EKM.
 The FKM algorithm also struggles with selecting the optimal fuzzifier value, $m$.
 Despite numerous studies discussing the selection of $m$, a widely accepted solution has yet to be found [31].
-Determining optimal parameters for a clustering algorithm often involves trial and error, as the real data structure is unknown.
-As a rule of thumb, when the data space dimension is less than or equal to three, setting $\alpha=1$ appears effective after normalizing the data to have a zero mean and unit variance for each dimension.
-As the data space dimension increases, the data-centroid distance $d_{kn}$ tends to grow, necessitating a decrease in $\alpha$ to ensure that the exponential term $e^{-\alpha d_{kn}}$ remains within a normal range.
-In this case, we suggest setting $\alpha$ to ten times the data variance and gradually reducing it.
-A proper $\alpha$ value indicates the presence of repulsive forces, resulting in a sudden increase in centroid distances.
-Ultimately, techniques for visualizing clustering results can be employed to fine-tune the $\alpha$ value.
-This strategy is proven effective by experimental examples of deep clustering on the MNIST dataset in Section VI.
-We further discuss and validate the choice of $\alpha$ in the supplemental material.
+As a rule of thumb, when the dimension of the data space is less than or equal to three, setting $\alpha=1$ appears effective after normalizing the data to have a zero mean and unit variance for each dimension.
+As the data space dimension increases, the data-centroid distance $d_{kn}$ increases, necessitating a decrease in $\alpha$ to ensure that the exponential term $e^{-\alpha d_{kn}}$ falls within a normal range.
+Hence, in the case of a higher dimensional data clustering, we suggest initially setting $\alpha$ to ten times the data variance and gradually reducing it until a sudden increase in centroid-centroid distance is observed.
+Because the increase in centroid distance implies the emergence of repulsive forces that reduce the uniform effect.
+This strategy is applied in deep clustering on the MNIST dataset in Section VI.
+An adaptive selection formula of $\alpha$ and its validation are given in the supplemental material.
 V Numerical Experiments
-Numerical experiments are conducted to compare the performance of our proposed EKM algorithm with six other algorithms: (1) HKM, (2) FKM, (3) MEFC, (4) csiFKM [19], (5) siibFKM [20], and (6) the Gaussian mixture model.
-Multiprototype clustering algorithms are not used as benchmarks, as the multiprototype mechanisms can be combined with EKM for fair comparison.
-This is beyond the scope of this paper because the paper does not intend to find the best KM algorithm to pair with the multiprototype mechanism.
-To avoid local optima convergence, each clustering algorithm is executed ten times, presenting the lowest objective value result.
-Centroids are initialized independently for each replication using the K-means++ algorithm [33].
-Convergence is deemed achieved when the relative moving distance of centroids between successive iterations is less than $1\mathrm{e}{-3}$, i.e.,
+Numerical experiments are conducted to compare the performance of our proposed EKM algorithm with six other centroid-based algorithms: (1) HKM, (2) FKM, (3) MEFC, (4) csiFKM [19], (5) siibFKM [20], and (6) the GMM.
+Multiprototype-based algorithms such as [22, 5] are not appropriate as baseline models because they are too complex to be benchmarks for gauging the efficiency of the proposed EKM.
+To avoid local optima convergence, each clustering algorithm is executed ten times, and the result with the lowest objective value is presented.
+Centroids are initialized independently for each replication by the K-means++ algorithm [32].
+Convergence is achieved when the relative moving distance of centroids between successive iterations is less than $1\mathrm{e}{-3}$, i.e.,
 $\frac{\bigg{(}\sum_{k=1}^{K}\|\mathbf{c}_{k}^{(\tau)}-\mathbf{c}_{k}^{(\tau-1)}\|_{2}^{2}\bigg{)}^{1/2}}{\bigg{(}\sum_{k=1}^{K}\|\mathbf{c}_{k}^{(\tau)}\|_{2}^{2}\bigg{)}^{1/2}}\leq 1\mathrm{e}{-3}.$ (44)
@@
 FKM, csiFKM, and siibFKM employ a typical fuzzifier value of $m=2$ [33].
-The fuzzifier value of MEFC is set to $\lambda=1$ and the smoothness value of EKM is set to $\alpha=1$.
-We test the algorithms on 12 datasets, including two synthetic and ten real datasets.
-Because of space limitations, only partial experiment results from six datasets are shown in the paper.
-Full experiment results with quantitative metrics can be found in the supplemental material.
-The six datasets are introduced below.
+The fuzzifier value of MEFC is set to $\lambda=1$ and the smoothing parameter of EKM is set to $\alpha=1$.
+We evaluate EKM’s performance on 12 datasets, including two synthetic and ten real datasets.
+Besides, we also apply EKM to the application of image segmentation.
+Because of space limitations, experiment results from six datasets and the image segmentation are shown in the paper.
+Full experiment results can be found in the supplemental material.
+The six datasets and the test image are introduced below.
 V-A Datasets
 V-A1 Synthetic 2-D Ball
-We create a synthetic dataset with data independently generated from three distinct normal distributions, each having different means: $(-2,2)$, $(2,-2)$, and $(4,4)$.
-Their covariance matrices are all identity matrices.
-The dataset consists of 2,000 samples from the normal distribution with means $(-2,2)$, and 50 samples from the other two normal distributions, resulting in a total of 2,100 samples.
+We generate data from three distinct normal distributions, each having different means: $(-2,2)$, $(2,-2)$, and $(4,4)$, and the same identity covariance matrix.
+The dataset consists of 2,100 samples, of which 2000 samples from the normal distribution with means $(-2,2)$, and 100 samples from the other two normal distributions (each distribution for 50 samples).
 The clustering results are shown in Fig.
@@
 V-A2 Synthetic 2-D Noisy Ball
-We add 100 uniformly distributed noise data samples (around 5% of the number of clean samples) to the synthetic 2-D ball dataset to test the algorithms’ robustness against noise.
+We add 100 uniformly distributed noise data (around 5% of the number of clean samples) to the synthetic 2-D ball dataset in order to test the algorithms’ robustness against noise.
 The clustering outcomes are displayed in Fig.
@@
 Because the instance distribution in the Iris dataset is balanced.
-To examine the algorithms’ capability for imbalanced clustering, we generate an imbalanced Fisher’s Iris data by removing the first 30 instances of Iris setosa, and merging the other two species (Iris virginica and Iris versicolor) into a single class.
-The clustering results for this dataset are presented in Fig.
+To examine the capability of imbalanced data clustering, we generate an imbalanced Fisher’s Iris dataset by removing the first 30 instances of Iris setosa, and merging the other two species (Iris virginica and Iris versicolor) into a single class.
+The clustering results of this imbalanced dataset are presented in Fig.
 7.
@@
 Wisconsin Diagnostic Breast Cancer (WDBC) [36] comprises 30 features of cell nuclei from diagnostic benign and malignant breast tumors, obtained from digitized images of fine-needle aspiration of breast masses.
-This database contains 569 instances, with a fairly even distribution of 357 benign and 212 malignant cases.
-We utilize the first three features (mean radius, mean texture, and mean perimeter) for clustering, each normalized with zero mean and unit variance.
+This database contains 569 instances, with an instance distribution of 357 benign and 212 malignant cases.
+We use the first three features (mean radius, mean texture, and mean perimeter) for clustering, each normalized with zero mean and unit variance.
 Fig.
@@
 The data distribution in WDBC is quite balanced.
-To evaluate the imbalanced data clustering performance, we remove the first 200 malignant instances in WDBC, resulting in an imbalanced WDBC database with a class distribution of 357 benign and 12 malignant instances.
+To examine the ability of imbalanced data clustering, we remove the first 200 malignant instances in WDBC, resulting in an imbalanced WDBC database with a distribution of 357 benign and 12 malignant instances.
 We still use the first three features for clustering, each normalized with zero mean and unit variance.
@@
 9.
-V-A7 ColoredChips
-This is not a dataset but an image named coloredChips.png from Matlab’s image gallery (shown in Fig.
+V-A7 Color-based image segmentation
+The test image coloredChips.png is chosen from Matlab’s image gallery (shown in Fig.
 10a).
-The purpose of this experiment is to test the algorithms’ performance on the task of color-based image segmentation.
 The number of clusters is set to five to expect the segmentation of the four colored chips and the background (i.e., the desk surface).
-Clustering is conducted in the RGB space and the segmentation results can be viewed in Fig.
+Clustering is conducted in the RGB color space and the segmentation results can be viewed in Fig.
 10b - 10e.
@@
 In general, these results demonstrate that the proposed EKM algorithm performs competitively on balanced data and outperforms the other clustering algorithms on imbalanced data.
-In particular, HKM, FKM, and MEFC cannot find the correct clusters in imbalanced datasets.
-The csiFKM and siibFKM algorithms are sensitive to noise data.
-Fig.
+In particular, HKM, FKM, and MEFC suffer from the uniform effect when data has an imbalanced distribution.
+The csiFKM and siibFKM algorithms are sensitive to noise data: Fig.
 5c and Fig.
-5d show that their centroids are highly deviated from the correct positions because of noise.
-In comparison, EKM is more robust, as seen in Fig.
+5d show that their centroids are highly deviated from the correct positions in the presence of noise.
+In comparison, EKM is more robust to noise because of the anchoring effect produced by the positive gradients (see Section IV-B), as seen in Fig.
 5e.
-Fig.
-9 displays the clustering results of HKM, FKM, GMM, and EKM for the imbalanced WDBC dataset.
-The strong performance of EKM indicates its potential application in anomaly detection, particularly in identifying clustered anomalies.
-Although HKM and FKM are also used for anomaly detection, they typically detect isolated anomalies, not clustered ones [38, 39, 40, 41].
-The color-based image segmentation results are shown in Fig.
+The strong performance of EKM shown in Fig.
+9 demonstrates its ability to identify clustered anomalies.
+This makes up a missing function of HKM and FKM in anomaly detection: they are typically used to detect isolated anomalies, not clustered anomalies [37, 38, 39, 40].
+The results of color-based image segmentation are shown in Fig.
 10b - 10e.
-It can be seen that HKM, FKM, and MEFC divide the background (the desk surface) into two clusters.
-This happens because the background comprises most pixels, and splitting it into two clusters better balances the number of pixels within the clusters.
+HKM, FKM, and MEFC erroneously separate the background (the desk surface) into two clusters.
+This happens because the background comprises most pixels, and splitting it into two clusters better balances the number of pixels within each clusters.
 On the other hand, EKM produces the expected segmentation result.
@@
 Figure 9: (a) Imbalanced Wisconsin Diagnostic Breast Cancer dataset. (b) Clustering by HKM. (c) Clustering by FKM. (d) Clustering by GMM. (e) Clustering by EKM.
-Figure 10: (a) coloredChips.png (b) Segmentation by HKM. (c) Segmentation by FKM. (d) Segmentation by MEFC. (e) Segmentation by EKM.
+Figure 10: (a) coloredChips.png. (b) Segmentation by HKM. (c) Segmentation by FKM. (d) Segmentation by MEFC. (e) Segmentation by EKM.
 VI Deep Clustering
-Clustering algorithms using Euclidean distance for defining similarity often face challenges when clustering high-dimensional data, such as images, due to the small distance differences between various point pairs in high-dimensional space [42].
-Deep clustering is a technique that addresses this issue by employing DNNs to map high-dimensional data non-linearly to low-dimensional representation, making the subsequent clustering easier.
-Most existing deep clustering approaches combine DNNs with HKM, e.g., [12, 13, 14].
-However, this design is less favorable to imbalanced datasets [13].
-In this section, we provide an example of replacing HKM with EKM in a popular deep clustering framework, showing its performance improvement on an imbalanced dataset derived from the MNIST.
+Clustering algorithms using Euclidean distance for defining similarity face a challenge when clustering high-dimensional data, such as images, due to the small distance differences between various point pairs in high-dimensional space [41].
+Deep clustering is a technique that addresses this issue by employing DNNs to map high-dimensional data to low-dimensional representation.
+Many deep clustering methods combine DNNs with HKM, e.g., [15, 16, 17].
+However, this design is less favorable to imbalanced data [16].
+In this section, we show the promise of EKM in the deep clustering of imbalanced data by replacing HKM with EKM in a popular deep clustering framework.
 VI-A Optimization Procedure
 Figure 11: Illustration of the DCN framework [15]. The parameters of the encoder, decoder, and clustering model are optimized jointly to minimize the reconstruction error and clustering error.
-Deep clustering network (DCN) [12] is a popular deep clustering framework, as illustrated in Fig.
-11.
-Essentially, DCN maps high-dimensional data to a low-dimensional representation through an autoencoder network.
-An autodecoder follows the autoencoder, mapping the representation back to the original high-dimension space (i.e., reconstruction).
-To ensure that the low-dimensional representation maintains the primary information of the original data, the autoencoder and autodecoder are jointly trained to minimize the reconstruction error.
-Additionally, to create a clustering-friendly structure in the low-dimensional representation, a clustering error is minimized along with the reconstruction error.
+Deep clustering network (DCN) [15] is a popular deep clustering framework.
+As illustrated in Fig.
+11, DCN maps high-dimensional data to low-dimensional representation through an autoencoder network.
+An autodecoder follows the autoencoder, mapping the representation back to the original high-dimensional space (i.e., reconstruction).
+To ensure that the low-dimensional representation maintains the primary information of the original data, the autoencoder and the autodecoder are jointly trained to minimize the reconstruction error.
+Additionally, to make the low-dimensional representation have a clustering-friendly structure, a clustering error is minimized along with the reconstruction error.
 DCN uses an alternating optimization algorithm to minimize the total error, and the optimization process is described below.
@@
 where $m_{k}^{n}$ is the number of samples assigned to the $k$-th cluster before the incoming data $\mathbf{x}_{n}$, controlling the learning rate of the $k$-th centroid.
-Overall, the optimization procedure of DCN with HKM alternates between updating networks parameters $\{\bm{\theta}_{e},\bm{\theta}_{d}\}$ by solving (45) and updating HKM parameters $\{\mathbf{C},\{\mathbf{s}_{n}\}\}$ by (46) and (47).
+Overall, the optimization procedure of DCN alternates between updating networks parameters $\{\bm{\theta}_{e},\bm{\theta}_{d}\}$ by solving (45) and updating HKM parameters $\{\mathbf{C},\{\mathbf{s}_{n}\}\}$ by (46) and (47).
 However, the centroid updating rule (47) is problematic for imbalanced data due to the uniform effect. To address this issue, we propose to replace (47) with the batch-learning version of EKM:
@@
 There are other details and tricks to implement DCN, such as the initialization of the networks.
-We only introduce the part related to our contribution, and for the rest, the readers can refer to the original paper of DCN [12].
+We only introduce the part related to our contribution, and kindly refer to [15] for more implementation details.
 VI-B Clustering Performance on MNIST
-To implement DCN, we refer to the code provided by one of its authors, available at https://github.com/boyangumn/DCN-New.
+To implement DCN, we refer to the code one of its authors provided, available at https://github.com/boyangumn/DCN-New.
 We use the default neural network structure and hyperparameters.
-In particular, the deep representation dimension is set to ten, and the parameter $\beta$ is set to one.
-The smoothness parameter $\alpha$ for EKM is determined by the strategy introduced in Section IV-C.
+In particular, the dimension of the low-dimensional representation is set to ten, and the parameter $\beta$ is set to one.
+The smoothing parameter $\alpha$ is tuned on the representation obtained by the initialized DCN networks according to the strategy introduced in Section IV-C.
 We evaluate clustering performance using three widely accepted metrics: including normalized mutual information (NMI), adjusted Rand index (ARI), and clustering accuracy (ACC).
@@
 ARI ranges from -1 to 1, with minus one representing the worst and one representing the best.
-We first validate the algorithms on the full MNIST dataset [43], containing 70,000 gray images of handwritten digits from 0 to 9.
-We set the smoothness parameter of EKM to $\alpha=5\mathrm{e}-3$.
-The clustering results are presented in TABLE I, where we compare the proposed DCN+EKM with DCN+HKM and stacked autoencoder (SAE).
+We first evaluate the algorithm’s performance on the full MNIST dataset [42], which contains 70,000 gray images of handwritten digits from 0 to 9.
+Each image has $28\times 28=784$ pixels.
+We set the smoothing parameter of EKM to $\alpha=5\mathrm{e}-3$.
+The clustering results are presented in TABLE I.
+We compare the proposed DCN+EKM with DCN+HKM and stacked autoencoder (SAE).
 SAE is a specific version of DCN that only minimizes the reconstruction error.
-We omit the results of DCN+FKM and DCN+MEFC due to space limitations and their poorer performance compared to DCN+HKM.
-The results show that DCN outperforms SAE, highlighting the importance of a clustering-friendly representation.
-The table also reveals that DCN+EKM performs similarly to DCN+HKM for this balanced dataset.
-To evaluate the performance of the algorithms for imbalanced data clustering, we remove the MNIST dataset’s training set samples with numbers 1 to 9 (leaving only the test set samples).
-This imbalanced dataset has 15,923 images, with around 6,900 being the number 0 and the remaining numbers (1 to 9) being around 1,000.
-We set the smoothness parameter of EKM to $\alpha=3.8\mathrm{e}-3$.
-The clustering results are summarised in TABLE II, showing that DCN+EKM significantly outperforms DCN+HKM.
-To analyze the reason, we use t-SNE [44] to map the ten-dimensional representation obtained by DCN to a two-dimensional space for visualization, with the results displayed in Fig.
+Thus, the learned representation by SAE does not have a clustering-friendly structure.
+The results show that DCN outperforms SAE, highlighting the importance of a clustering-friendly structure.
+We can also see that DCN+EKM performs similarly to DCN+HKM.
+We omit the results of DCN+FKM and DCN+MEFC due to space limitations.
+Their performance is not better than DCN+HKM.
+MNIST contains 60,000 training images and 10,000 testing images.
+To evaluate the performance of imbalanced data clustering, we generate an imbalanced MNIST dataset by removing the training images of digits 1 to 9.
+This imbalanced dataset contains 15,923 images, of which approximately 6,900 are digit 0, while the remaining digits (1 to 9) each have about 1,000 images.
+We set the smoothing parameter of EKM to $\alpha=3.8\mathrm{e}-3$.
+The results are summarized in TABLE II.
+In conclusion, we find that an EKM-friendly representation structure is crucial for deep clustering of imbalanced data.
+We map the ten-dimensional representation obtained by DCN to a two-dimensional space by t-SNE [43] for visualization.
+The results are displayed in Fig.
 12.
-We observe that the deep representation obtained by DCN+EKM is more discriminative than DCN+HKM, and DCN+EKM successfully separates the large class (number 0) from other small classes, while DCN+HKM erroneously groups the large class into four clusters.
+The visualization result implies that the deep representation obtained by DCN+EKM is more discriminative than that obtained by DCN+HKM.
+We observe that EKM successfully identifies the large class (number 0) from other small classes without referring to the true labels, while HKM incorrectly divides the large class into four clusters.
 TABLE I: Evaluation on full MNIST
@@
 Methods SAE+HKM DCN+HKM SAE+EKM DCN+EKM NMI 0.551 0.584 0.583 0.701 ARI 0.317 0.325 0.396 0.826 ACC 0.413 0.434 0.497 0.784
-Figure 12: Visualization of low-dimensional DCN representations by t-SNE.
+Figure 12: Visualization of DCN representations by t-SNE.
 (a) Representations obtained by DCN+HKM, colored by true labels.
-(b) Representations obtained by DCN+HKM, colored by clusters of HKM.
+(b) Representations obtained by DCN+HKM, colored by labels generated by HKM.
 (c) Representations obtained by DCN+EKM, colored by true labels.
-(d) Representations obtained by DCN+EKM, colored by clusters of EKM.
+(d) Representations obtained by DCN+EKM, colored by labels generated by EKM.
 VII Conclusion
-This paper presents equilibrium K-means (EKM), a novel clustering algorithm that is effective for imbalanced data.
-EKM is simple, interpretable, scalable to large datasets, and capable of finding unequal-sized clusters, making it a better alternative for hard K-means (HKM) and fuzzy K-means (FKM) for imbalanced data.
-Experimental results on real datasets from various domains demonstrate that EKM outperforms HKM and FKM on datasets with imbalanced data and performs comparably on datasets with balanced data.
-Furthermore, we provide a general framework that unifies HKM, FKM, and EKM.
-We encourage readers to explore and develop new K-means algorithms based on this more flexible framework to address other specific K-clustering problems besides imbalanced data distribution.
+This paper presents equilibrium K-means (EKM), a novel clustering algorithm effective for both balanced and imbalanced data.
+EKM is simple, interpretable, scalable to large datasets, and capable of creating unequal-sized clusters.
+Experimental results on real datasets from various domains show that EKM outperforms HKM, FKM, and other popular centroid-based algorithms on datasets with imbalanced data and performs comparably on datasets with balanced data.
+In the same deep clustering framework, compared with HKM and FKM, the clustering accuracy of imbalanced data using EKM has been noticeably improved.
+Furthermore, we reformulate HKM, FKM, and EKM in a general form of gradient descent.
+We encourage readers to study more properties of this general form to develop new K-means algorithms with non-heuristic memberships to address different issues.
 Appendix A Proof of Theorem 1
@@
 Hence, the sequence $(J^{(1)},J^{(2)},\cdots)$ is non-increasing, and with the boundness condition that $h>-\infty$, we have $\lim_{\tau\to+\infty}(J^{(\tau)}-J^{(\tau+1)})\to 0$.
-If the left side of the inequality (51) converges to zero, the right side of the inequality also converges to zero since it is non-negative.
+If the left side of the inequality (LABEL:theorem1_proof_inequality) converges to zero, the right side of the inequality also converges to zero since it is non-negative.
 Consequently, we have $\lim_{\tau\to+\infty}\partial_{\mathbf{c}_{k}}J^{(\tau)}\to 0$ for all $k$.
@@
 We evaluate the clustering performance of the proposed equilibrium K-means (EKM) on ten real datasets.
-Four datasets (Iris, Imbalanced Iris, Wisconsin Diagnostic Breast Cancer (WDBC), and Imbalanced WDBC) are discussed in detail in the main text.
-This supplemental material presents the results for the remaining six datasets from the UCI repository (available at https://archive.ics.uci.edu/datasets): Wine, Ecoli, Image Segmentation (IS), Htru2, Rice, and Dry Bean.
-A detailed description of all ten datasets can be found in TABLE SI, ordered according to the imbalance level measured by CV0, which we will define later.
-Datasets with a lower order have higher CV0 values, indicating a higher imbalance level.
-We provide some necessary explanations for this table below.
-In the “Feature” column, the format “x (y)” indicates that the dataset has y features, with x features used for clustering.
-For instance, the Iris dataset has four features, with two used for clustering.
-Only the four datasets in the main text use partial features for clustering, which serves the purpose of visualizing clustering results.
-For the additional six datasets, all features are used since visualization is not required.
-In the “CV0” column, we calculate the coefficient of variation (CV) to measure data imbalance.
-CV is defined as the ratio of the standard deviation to the mean.
+Four datasets (Iris, Imbalanced Iris, Wisconsin Diagnostic Breast Cancer (WDBC), and Imbalanced WDBC) are introduced in detail in the main text.
+Herein we supplement the results for the remaining six datasets.
+They are from the UCI repository (available at https://archive.ics.uci.edu/datasets), which are: Wine, Ecoli, Image Segmentation (IS), Htru2, Rice, and Dry Bean.
+A detailed description of all ten datasets are summarized in TABLE SI, ordered according to the imbalance level measured by CV0, which we will define later.
+The dataset with a lower order has a higher CV0 value, indicating a higher imbalance level.
+In the “Feature” column of TABLE SI, the format “x (y)” indicates that the dataset has y features, of which x features are used for clustering.
+For instance, the Iris dataset has four features, of which two are used for clustering.
+Partial features are used for the sake of visualizing clustering results.
+In the “CV0” column, we calculate the coefficient of variation (CV) of the dataset to measure the level of data imbalance.
 CV is defined as the ratio of the standard deviation to the mean.
@@
 The number of clusters is assumed to be equal to the number of classes for all datasets.
-For the supplementary six datasets, the smoothness parameter $\alpha$ is chosen as follows:
+For the supplementary six datasets, the smoothness parameter $\alpha$ is chosen adaptively as follows:
 $\alpha=2/\bar{d}_{0n},$
@@
 The value $\bar{d}_{0n}$ can be interpreted as a data variance after the normalization of zero data mean.
-The normalized mutual information (NMI), adjusted Rand index (ARI), and clustering accuracy (ACC) metrics of each algorithm on the ten datasets are presented in TABLE SII.
+The normalized mutual information (NMI), adjusted Rand index (ARI), and clustering accuracy (ACC) of each algorithm on the ten datasets are presented in TABLE SII.
 Bold values indicate the best results.
-We also include three more metrics: “CV1”, ”DCV” and ”Running time” for readers’ interest.
+We also include three metrics: “CV1”, ”DCV” and ”Running time” for readers’ interest.
 “CV1” is the CV value of the clustering result.
 “DCV” is the difference between CV0 and CV1, i.e., $\text{DCV}=\text{CV}0-\text{CV}1$.
-A positive DCV value indicates that the data distribution in the true classes is more imbalanced than in the clusters, while a negative DCV value means the data distribution in the true classes is more uniform than in the clusters.
+A positive DCV value indicates that the data distribution of the true classes is more imbalanced than that of the generated clusters, while a negative DCV value means the data distribution of the true classes is more uniform than that of the generated clusters.
 “Running time” is the sum of the running time of ten replicates.
@@
 On the rest of the datasets, EKM still has a competitive performance.
-The running time of EKM is also comparable to FKM.
-Overall, these results demonstrate that EKM is an efficient algorithm for clustering balanced and imbalanced data.
+Overall, these results demonstrate that EKM is an efficient clustering algorithm for both balanced and imbalanced data.
 TABLE SI: Detailed description of REAL-WORLD Datasets

### v3

@@
 Abstract
-Traditional centroid-based clustering algorithms, such as hard K-means (HKM, or Lloyd’s algorithm) and fuzzy K-means (FKM, or Bezdek’s algorithm), display degraded performance when true underlying groups of data have varying sizes (i.e., imbalanced data).
-This paper introduces equilibrium K-means (EKM), a novel fuzzy clustering algorithm that has the robustness to imbalanced data by preventing centroids from crowding together in the center of large clusters.
-EKM is simple, alternating between two steps; fast, with the same time and space complexity as FKM; and scalable to large datasets.
-We evaluate the performance of EKM on two synthetic and ten real datasets, comparing it to other centroid-based algorithms, including HKM, FKM, maximum-entropy fuzzy clustering (MEFC), two FKM variations designed for imbalanced data, and the Gaussian mixture model.
-The results show that EKM performs competitively on balanced data and significantly outperforms other algorithms on imbalanced data.
-Deep clustering experiments on the MNIST dataset demonstrate the significance of making representation have an EKM-friendly structure when dealing with imbalanced data; In comparison to deep clustering with HKM, deep clustering with EKM obtains a more discriminative representation and a 35% improvement in clustering accuracy.
-Additionally, we reformulate HKM, FKM, MEFC, and EKM in a general form of gradient descent, where fuzziness is introduced differently and more simply than in Bezdek’s work, and demonstrate how the general form facilitates a uniform study of KM algorithms.
+Centroid-based clustering algorithms, such as hard K-means (HKM) and fuzzy K-means (FKM), have suffered from learning bias towards large clusters.
+Their centroids tend to be crowded in large clusters, compromising performance when the true underlying data groups vary in size (i.e., imbalanced data).
+To address this, we propose a new clustering objective function based on the Boltzmann operator, which introduces a novel centroid repulsion mechanism, where data points surrounding the centroids repel other centroids.
+Larger clusters repel more, effectively mitigating the issue of large cluster learning bias.
+The proposed new algorithm, called equilibrium K-means (EKM), is simple, alternating between two steps; resource-saving, with the same time and space complexity as FKM; and scalable to large datasets via batch learning.
+We substantially evaluate the performance of EKM on synthetic and real-world datasets.
+The results show that EKM performs competitively on balanced data and significantly outperforms benchmark algorithms on imbalanced data.
+Deep clustering experiments demonstrate that EKM is a better alternative to HKM and FKM on imbalanced data as more discriminative representation can be obtained.
+Additionally, we reformulate HKM, FKM, and EKM in a general form of gradient descent and demonstrate how this general form facilitates a uniform study of K-means algorithms.
 Index Terms:
-K-means, fuzzy clustering, imbalance learning, deep clustering
+K-means, fuzzy clustering, imbalance learning, uniform effect, deep clustering
 I Introduction
@@
 While there is a considerable amount of research on supervised learning (e.g., classification) from imbalanced data [2, 3, 4], unsupervised learning has not been as thoroughly explored, because the unknown cluster sizes make the task more difficult [5].
-Methods like resampling and boosting frequently used in supervised learning cannot be applied in unsupervised learning.
+Methods like resampling and boosting frequently used in supervised learning cannot be applied in unsupervised learning due to the lack of labels.
 Clustering is an important unsupervised learning task involving grouping data into clusters based on similarity.
@@
 HKM assigns a data point to only one cluster, while FKM assigns a data point to multiple clusters with varying degrees of membership.
-Research on HKM and FKM has been ongoing.
-For example, the possibilistic FKM (PFKM) [11] was proposed to address the sensitivity of FKM to noise and outliers.
+HKM and FKM stimulate many subsequent research.
+For example, the possibilistic K-means (PKM) [11] was proposed with reformulated membership more meaningful in terms of typicality.
+Later, the possibilistic fuzzy K-means (PFKM) [12] was proposed to address the noise sensitivity defect of FKM and overcome the coincident clusters problem of PKM.
 FKM-$\sigma$ [13] was proposed to improve FKM’s performance on data points with uneven variations or non-spherical shapes in individual clusters.
-Fuzzy local information K-means (FLIKM) [13] was designed to promote FKM’s performance in image segmentation.
-Deep clustering, a technique that combines deep neural networks (DNNs) with HKM and FKM, was proposed to cluster high-dimensional data [14, 15, 16, 17].
-Although many variations based on HKM and FKM have been developed to deal with different situations, most of them display degraded effectiveness in cases of imbalanced data.
+Fuzzy local information K-means [14] was designed to promote FKM’s performance in image segmentation.
+Recently, feature-weighted PKM (FWPKM) [15, 16] was proposed to give non-uniform importance to features.
+Research of combining K-means and kernel mechanisms [17, 18, 19] has garnered interest, which non-linearly mapped data from low-dimensional space to high-dimensional space through kernel functions to enhance data separability.
+Conversely, deep clustering, a modern technique that integrates deep neural networks (DNNs) and K-means (or other clustering algorithms), has been proposed to cluster high-dimensional data by mapping them to a low-dimensional space to overcome the curse of dimensionality [20, 21, 22, 23].
+Although many variations derived from HKM and FKM have been developed to deal with different situations, most display degraded effectiveness in cases of imbalanced data.
 This is due to the so-called “uniform effect” that causes the clusters generated by these algorithms to have similar sizes even when input data has highly varying group sizes [24].
+It is reported that FKM has a stronger uniform effect than HKM [25].
 An illustration of the uniform effect of HKM and FKM is given in Fig.
-1 where we can observe that the centroids of HKM and FKM crowd together in the large true cluster.
+1 where we can observe that the centroids of HKM and FKM crowd together in the large cluster.
 Figure 1: Clustering results of a highly imbalanced dataset.
 (a) Ground truth.
-The colors represent the labels of the data points.
+The colors represent the reference labels of the data points.
 (b) Clustering by hard K-means.
@@
 (d) Clustering by the proposed equilibrium K-means.
-They are all two-step alternating algorithms that iteratively compute centroids.
+They are all two-step alternating algorithms that iteratively update centroids.
 I-A Existing Efforts to Overcome Uniform Effect
-To the best of our knowledge, there are two popular methods to overcome the uniform effect in the field of fuzzy systems.
+There are two popular methods to overcome the uniform effect.
 The first method is to introduce more weight on the data points in small clusters at each updating iteration, biasing learning towards them, like a modified FKM called cluster-size insensitive FKM (csiFKM) developed by Noordam et al.
@@
 I-B Our Contributions
-In this paper, we propose equilibrium K-means (EKM) to address the issue of imbalanced data clustering and the aforementioned algorithms’ limitations.
+In this paper, we propose equilibrium K-means (EKM), which addresses the imbalanced data clustering issue by a new centroid repulsion mechanism without the above limitations.
 Our contributions can be summarized in three main aspects.
@@
 This general form facilitates the uniform study of KM algorithms.
-Second, we develop EKM based on the first contribution.
-EKM belongs to the family of fuzzy clustering and membership defined in EKM has a clear physical meaning.
-Repulsive forces appear among centroids of EKM, successfully reducing the uniform effect by preventing centroids from crowding together in a large cluster (see Fig.
+Second, we develop EKM based on the first contribution where the Boltzmann operator is utilized as an approximation method.
+Such approximation brings a brand-new centroid repulsion mechanism.
+Data surrounding centroids repel other centroids, and large clusters repel more as they contain more data, successfully reducing the uniform effect by preventing centroids from crowding together in a large cluster (see Fig.
 1d for an example).
-Therefore, EKM is effective in dealing with imbalanced data and is not sensitive to noise.
+Hence, EKM is robust to imbalanced data and not sensitive to noise.
 Similar to HKM and FKM, EKM is a two-step alternating algorithm that iteratively computes centroids.
 In addition to having the same time ($O(N)$) and space complexity as FKM, EKM has a batch-learning version that can be applied to large datasets.
-We demonstrate the effectiveness of EKM on balanced and imbalanced data by conducting numerical experiments on two synthetic and ten real datasets from different domains.
-Due to space limitations, quantitative performance metrics are provided in the supplementary material.
-The scatter plots of partial clustering results we provide in the paper are sufficient to demonstrate the superiority of EKM on imbalanced data.
-Finally, we investigate the joint learning of DNNs and EKM.
+A comprehensive study is conducted using four synthetic and 16 real datasets (hence 20 datasets in total) to demonstrate the effectiveness of EKM on balanced and imbalanced data.
+Finally, we investigate the combination of DNNs and EKM.
 We find that mapping high-dimensional data via DNNs to an EKM-friendly space can result in more discriminative low-dimensional representation than mapping to an HKM-friendly space.
-When tested on an imbalanced dataset derived from MNIST, joint learning of DNNs and EKM improves clustering accuracy by 35% compared to joint learning of DNNs and HKM.
+Compared to the combination of DNNs and HKM, using EKM improves clustering accuracy by 35% on an imbalanced dataset derived from MNIST.
 I-C Organization
@@
 In Section III, we derive their general form and the proposed EKM.
-The properties of EKM are studied in Section IV.
+The properties of EKM and its centroid repulsion mechanism are studied in Section IV.
 We evaluate the performance of EKM on classic clustering tasks in Section V and on deep clustering in Section VI.
@@
 II-B The Uniform Effect of K-Means
-The uniform effect refers to the propensity to generate clusters of similar sizes, which is implicitly implied in the objective (1) of KM.
+The uniform effect refers to the propensity to generate clusters of similar sizes.
+It is essentially a learning bias towards large clusters and is implicitly implied in the objective (1) of KM.
 For simplicity, kindly consider a case with two clusters (i.e., $K=2$).
@@
 $\gamma^{(\tau)}_{k}=1/\bigl{(}\sum_{n=1}^{N}\frac{e^{-\alpha d_{kn}^{(\tau)}}}{\sum_{i}e^{-\alpha d_{in}^{(\tau)}}}[1-\alpha(d_{kn}^{(\tau)}-\frac{\sum_{i}d_{in}^{(\tau)}e^{-\alpha d_{in}^{(\tau)}}}{\sum_{i}e^{-\alpha d_{in}^{(\tau)}}})]\bigl{)}.$ (37)
-This updating procedure can be reformulated into a two-step iteration procedure akin to FKM, that is
-$\mathbf{c}_{k}^{(\tau+1)}=\frac{\sum_{n}w_{kn}^{(\tau)}\mathbf{x}_{n}}{\sum_{n}w_{kn}^{(\tau)}},$ (38)
-where weights $w_{kn}^{(\tau)}$ are calculated by
+This updating procedure which we call EKM can be reformulated into a two-step iteration procedure akin to FKM:
+  1. 1.
+    Calculate the weight value of the $n$-th data point to the $k$-th cluster by
     $w_{kn}^{(\tau)}=\frac{e^{-\alpha d_{kn}^{(\tau)}}}{\sum_{i=1}^{K}e^{-\alpha d_{in}^{(\tau)}}}[1-\alpha(d_{kn}^{(\tau)}-\frac{\sum_{i=1}^{K}d_{in}^{(\tau)}e^{-\alpha d_{in}^{(\tau)}}}{\sum_{i=1}^{K}e^{-\alpha d_{in}^{(\tau)}}})].$ (38)
-The time complexity of one iteration is $O(NK^{2})$. Although the sum of weights equals one, i.e.,
-$\sum_{k=1}^{K}w_{kn}^{(\tau)}=1,\,\forall n,$ (40)
-it is worth noting that these weights cannot be interpreted as probabilities or memberships since some weight values are negative. The membership of EKM is defined in Section III-G.
+  2. 2.
+    Recalculate the weighted centroid of the $k$-th cluster by
+    $\mathbf{c}_{k}^{(\tau+1)}=\frac{\sum_{n}w_{kn}^{(\tau)}\mathbf{x}_{n}}{\sum_{n}w_{kn}^{(\tau)}}.$ (39)
+EKM converges when centroids cease to change or the maximum number of iterations is reached.
+The time complexity of one iteration of the above two steps is $O(NK^{2})$.
+For convenience, we summarise the complete procedure of EKM in Algorithm 1.
+Input: A dataset $X=\{\mathbf{x}_{n}\}_{n=1}^{N}$, cluster number $K$, initial centroids $\{\mathbf{c}_{k}^{(0)}\}_{k=1}^{K}$, smoothing parameter $\alpha$
+Output: Centroids $\{\mathbf{c}_{1},\cdots,\mathbf{c}_{K}\}$
+$\tau=0$;
+repeat
+       Compute weight $w_{kn}^{(\tau)}$ by (38) for all $k,\,n$;
+       Update centroid $\mathbf{c}_{k}^{(\tau+1)}$ by (39) for all $k$;
+       $\tau=\tau+1$;
+until convergence;
+return $\{\mathbf{c}_{k}^{(\tau)}\}_{k=1}^{K}$
+Algorithm 1 Equilibrium K-Means Algorithm
 III-F Physical Interpretation of Equilibrium K-Means
@@
 Hence, the objective function (34) equals the expectation of the entire system’s energy, and EKM seeks centroids to minimize this energy expectation.
-Due to this connection, we refer to the proposed algorithm (alternating between (38) and (39)) as equilibrium K-means.
+Due to this connection, we refer to the proposed Algorithm 1 as equilibrium K-means.
 III-G Membership Defined in Equilibrium K-Means
-According to the physical interpretation of EKM, the exponential term $e^{-\alpha d_{kn}}$ can be interpreted as the unnormalized probability of the $n$-th data point belonging to the $k$-th clusters.
+Membership can be defined in EKM. Although the sum of weights (38) of EKM equals one, i.e.,
+$\sum_{k=1}^{K}w_{kn}^{(\tau)}=1,\,\forall n,$ (40)
+it is worth noting that these weights cannot be interpreted as probabilities or memberships since some weight values are negative.
+We define membership of EKM from its physical perspective.
+From the physical interpretation of EKM, the exponential term $e^{-\alpha d_{kn}}$ can be interpreted as the unnormalized probability of the $n$-th data point belonging to the $k$-th clusters.
 Hence, the membership of the $n$-th data point to the $k$-th cluster can be defined as
 $u_{kn}=\frac{e^{-\alpha d_{kn}}}{\sum_{i=1}^{K}e^{-\alpha d_{in}}}.$ (41)
-Note that, although the membership formula of EKM is the same as that (10) of MEFC, the values are different because centroids are calculated using distinct formulas.
-We have no intention in this paper to compare membership defined in different fuzzy clustering algorithms; therefore, we only define membership in EKM and leave the further discussion to subsequent research.
+Since we have no intention in this paper to discuss the membership (it is just an intermediate product), we only define the membership in EKM and leave further discussion to subsequent research.
 III-H Convergence of Smooth K-Means
@@
 It can be easily verified that when the smooth minimum function $h$ is LogSumExp (19) and p-Norm (20) in which case SKM (42) is equivalent to MEFC and FKM, respectively, the above convergence condition can be satisfied with any initial centroids.
-Although the Boltzmann operator (22) is not concave, it is a smooth approximation of a concave function (the minimum function).
-Hence, EKM also exhibits good convergence behavior in numerical experiments.
+Since the Boltzmann operator (22) is not globally concave, discussing the convergence of EKM is not simple.
+If centroids are updated within the concave region of the Boltzmann operator, the convergence of EKM can be guaranteed by Theorem 1.
+However, the practical situation may be more complex.
+A rigorous convergence condition of EKM may involve a tedious discussion of a specific value of $\alpha$, positions of initial centroids, and data structures.
+Instead, we demonstrate the empirical convergence of EKM in Section V-F.
 IV Comparison of Different Smoothed Objectives
@@
 2d).
-We fix the first centroid at $\mu_{1}$, and plot the reformulated objectives of HKM (WCSS (12)), FKM ($J_{\text{PN}}$ (29)), MEFC ($J_{\text{LSE}}$ (23)), and EKM ($J_{\text{B}}$ (34)) as a function of the position of the second centroid in Fig.
-2.
+We plot $J(\mathbf{c}_{1}=\mu_{1},\mathbf{c}_{2})$ as a function of $\mathbf{c}_{2}$ in Fig.
+2, with $\mathbf{c}_{2}$ being the position of the second centroid on the $x$-axix, and $J$ being the reformulated objectives of HKM (WCSS (12)), FKM ($J_{\text{PN}}$ (29)), MEFC ($J_{\text{LSE}}$ (23)), and EKM ($J_{\text{B}}$ (34)).
 The four objective functions behave similarly on the first two balanced datasets, but it is worth noting that the last two imbalanced datasets.
@@
 In contrast, EKM does not have an obvious local minimum and its global minimum point aligns with the true cluster center, i.e., $\mu_{2}$, highlighting EKM’s superiority in handling imbalanced data.
-Figure 2: Objectives as a function of the second centroid moving along the $x$-axis (the first centroid is fixed at $\mu_{1}$).
+Figure 2: Objectives $J({\mathbf{c}_{1}=\mu_{1},\mathbf{c}_{2}})$ as a function of $\mathbf{c}_{2}$ (the position of the second centroid on the $x$-axis).
 The yellow, purple, green, and blue curves are reformulated objective functions of HKM, FKM, MEFC, and EKM, respectively.
 (a), (b), (c), and (d) present objective functions under different data distributions.
-IV-B The Analysis of EKM’s Effectiveness on Imbalanced Data
+IV-B The Analysis of EKM’s Effectiveness on Imbalanced Data: A Centroid Repulsion Mechanism
 We analyze the effectiveness of EKM on imbalanced data based on the gradient of its objective.
@@
 As evident from the figure, data points on the side of the first centroid do not impact the second centroid of HKM, but they do attract the second centroids of FKM and MEFC.
-This finding supports the statement in [30] that FKM has a stronger uniform effect than HKM.
-On the other hand, data points near the first centroid have repulsive forces on the second centroid of EKM, which compensate for the attraction from other data points, effectively reducing the uniform effect.
-It is worth noting that data points near the second centroid of EKM have the strongest attraction forces to this centroid.
-Consequently, EKM’s centroids are anchored by their surrounding data points, making their position less influenced by noise and outliers.
+This finding supports the claim in [25] that FKM has a stronger uniform effect than HKM.
+On the other hand, data points near the first centroid have repulsive forces on the second centroid of EKM, which compensate for the attraction from other data points.
+This new mechanism effectively reduces the algorithm’s learning bias towards other clusters, particularly large ones, because the centers of large clusters often contain more data.
+The existence of this mechanism suggests that the uniform effect of EKM is weak, a hypothesis that our experiments will test.
+Notably, data points near the second centroid of EKM exert the strongest attraction forces on this centroid.
+As a result, EKM’s centroids are stabilized by their surrounding data points, reducing their susceptibility to noise and outliers.
 Figure 3: Gradients as a function of a data point moving along the $x$-axis.
@@
 (b), (c), and (d) are gradients of $J_{\text{PN}}$, $J_{\text{LSE}}$, and $J_{\text{B}}$ with varied smoothing parameters, respectively.
-IV-C The Choice of $\alpha$
+IV-C The Choice of EKM’s parameter
 The smoothing parameter $\alpha$ impacts the performance of EKM, but the optimal choice remains unknown.
@@
 As the data space dimension increases, the data-centroid distance $d_{kn}$ increases, necessitating a decrease in $\alpha$ to ensure that the exponential term $e^{-\alpha d_{kn}}$ falls within a normal range.
-Hence, in the case of a higher dimensional data clustering, we suggest initially setting $\alpha$ to ten times the data variance and gradually reducing it until a sudden increase in centroid-centroid distance is observed.
+Hence, one can set $\alpha$ as a reciprocal of the data variance or manually tune it as follows.
+Initially setting $\alpha$ as ten times the reciprocal of data variance and gradually reducing it until a sudden increase in centroid-centroid distance is observed.
 Because the increase in centroid distance implies the emergence of repulsive forces that reduce the uniform effect.
-This strategy is applied in deep clustering on the MNIST dataset in Section VI.
-An adaptive selection formula of $\alpha$ and its validation are given in the supplemental material.
 V Numerical Experiments
-Numerical experiments are conducted to compare the performance of our proposed EKM algorithm with six other centroid-based algorithms: (1) HKM, (2) FKM, (3) MEFC, (4) csiFKM [19], (5) siibFKM [20], and (6) the GMM.
-Multiprototype-based algorithms such as [22, 5] are not appropriate as baseline models because they are too complex to be benchmarks for gauging the efficiency of the proposed EKM.
-To avoid local optima convergence, each clustering algorithm is executed ten times, and the result with the lowest objective value is presented.
-Centroids are initialized independently for each replication by the K-means++ algorithm [32].
-Convergence is achieved when the relative moving distance of centroids between successive iterations is less than $1\mathrm{e}{-3}$, i.e.,
+V-A Experimental Setup
+Numerical experiments are conducted to compare the performance of our proposed EKM algorithm with seven related centroid-based algorithms including (1) HKM, (2) FKM, (3) MEFC, (4) PFKM [12], (5) csiFKM [26], (6) siibFKM [27], and (7) FWPKM [15].
+Multiprototype clustering algorithms (e.g., [29, 5]) are not appropriate as baseline algorithms because they are too complex to be benchmarks for gauging the efficiency of EKM.
+We also avoid using hybrid methods, such as algorithms combining K-means with kernels proposed in [17, 18, 19], as benchmarks to ensure fairness.
+Implementation is conducted in Matlab R2022a and the operation system is Ubuntu 18.04.1 LTS with Intel Core i9-9900K CPU @ 3.60GHz X 16 and 62.7GiB memory.
+All datasets used and codes are available at https://github.com/ydcnanhe/Imbalanced-Data-Clustering-using-Equilibrium-K-Means.git.
+The experimental datasets contain four artificial datasets generated by us (including Data-A, Data-B, Data-C, and Data-D), 13 UCI [38] datasets (consisting of Image Segmentation (IS), Seeds, Wine, Rice, Wisconsin Diagnostic Breast Cancer (WDBC), Ecoli, Htru2, Zoo, Glass, Shill Bidding, Anuran Calls, Occupancy Detection, Machine Failure), and three Kaggle datasets (incorporating Heart Disease, Pulsar Cleaned, and Bert-Embedded Spam).
+So there are 20 datasets and Table I provides their information, including name, instance number, feature number, reference class number, and coefficient of variation (CV).
+CV is used in previous literature [34] to measure the level of data dispersion.
+It is calculated as the ratio of the standard deviation of class sizes to the mean.
+Given the number of instances in each class as $N_{1},\cdots,N_{K}$, we have
+$\text{CV}=s/\bar{N},$ (44)
+where
+$\bar{N}=\frac{\sum_{k=1}^{K}N_{k}}{K},\,s=\sqrt{\frac{\sum_{k=1}^{K}(N_{k}-\bar{N})^{2}}{K-1}}.$
+In [34], a CV value exceeding 1 indicates highly varying class sizes, while a value below 0.3 signifies uniform class sizes.
+However, there isn’t a widely accepted critical CV value implying that data is imbalanced.
+For rigorous statements, we establish that if a dataset’s CV is less than 0.4, it is considered balanced.
+Conversely, if the CV exceeds 0.7, the dataset is deemed imbalanced.
+Consequently, we have six balanced datasets and 14 imbalanced datasets.
+It should be noted that none of the datasets used has a CV value between 0.4 and 0.7, hence this range remains undefined.
+TABLE I: Detailed description of 20 Datasets
+ID Name Instances Feature Classes CV
+D1 Data-A 2250 2 3 1.4468
+D2 Data-B 2250 2 3 1.4468
+D3 Data-C 5200 2 2 1.3054
+D4 Data-D 5400 2 9 2.7500
+D5 IS 2310 19 7 0
+D6 Seeds 210 7 3 0
+D7 Heart Disease 1125 13 2 0.0373
+D8 Wine 178 13 3 0.1939
+D9 Rice 3810 7 2 0.2042
+D10 WDBC 569 30 3 0.3604
+D11 Zoo 101 16 7 0.8937
+D12 Glass 214 9 6 1.0767
+D13 Ecoli 336 7 8 1.1604
+D14 Htru2 17898 8 2 1.1552
+D15 Shill Bidding 6321 9 2 1.1122
+D16 Anuran Calls 7195 22 10 1.6016
+D17 Occupancy Detection 20560 5 2 0.7608
+D18 Machine Failure 9815 7 2 1.3318
+D19 Pulsar Cleaned 14987 7 2 1.3561
+D20 Bert-Embedded Spam 5572 768 2 1.0350
+All datasets undergo normalization, ensuring that each feature has zero mean and unit variance.
+All features are used for clustering purposes.
+The number of clusters is set to the reference class number.
+Convergence is achieved when the moving distance of centroids between successive iterations is sufficiently small relative to the magnitude of centroids, i.e.,
 $\frac{\bigg{(}\sum_{k=1}^{K}\|\mathbf{c}_{k}^{(\tau)}-\mathbf{c}_{k}^{(\tau-1)}\|_{2}^{2}\bigg{)}^{1/2}}{\bigg{(}\sum_{k=1}^{K}\|\mathbf{c}_{k}^{(\tau)}\|_{2}^{2}\bigg{)}^{1/2}}\leq 1\mathrm{e}{-3}.$ (45)
-The maximum number of iterations is set to 100.
-FKM, csiFKM, and siibFKM employ a typical fuzzifier value of $m=2$ [33].
-The fuzzifier value of MEFC is set to $\lambda=1$ and the smoothing parameter of EKM is set to $\alpha=1$.
-We evaluate EKM’s performance on 12 datasets, including two synthetic and ten real datasets.
-Besides, we also apply EKM to the application of image segmentation.
-Because of space limitations, experiment results from six datasets and the image segmentation are shown in the paper.
-Full experiment results can be found in the supplemental material.
-The six datasets and the test image are introduced below.
-V-A Datasets
-V-A1 Synthetic 2-D Ball
-We generate data from three distinct normal distributions, each having different means: $(-2,2)$, $(2,-2)$, and $(4,4)$, and the same identity covariance matrix.
-The dataset consists of 2,100 samples, of which 2000 samples from the normal distribution with means $(-2,2)$, and 100 samples from the other two normal distributions (each distribution for 50 samples).
-The clustering results are shown in Fig.
-4.
-V-A2 Synthetic 2-D Noisy Ball
-We add 100 uniformly distributed noise data (around 5% of the number of clean samples) to the synthetic 2-D ball dataset in order to test the algorithms’ robustness against noise.
-The clustering outcomes are displayed in Fig.
-5.
-V-A3 Fisher’s Iris Data
-The Fisher’s Iris dataset [34, 35] is a well-known dataset for evaluating the performance of clustering algorithms.
-It comprises 50 samples from each of three iris species (Iris setosa, Iris virginica, and Iris versicolor).
-Two features, the width and the length of the sepals, are utilized as clustering features and normalized to have zero means and unit variance.
+We set the maximum number of iterations to 500 to prevent algorithms from infinitely iterating due to failure to converge.
+Except for sporadic cases, all algorithms can reach the specified convergence within 500 iterations.
+Centroids are initialized using the K-means++ algorithm [39].
+Given that each run of K-means++ yields different outputs and the objectives of the tested clustering algorithms are known to be non-convex (i.e., multiple local optima exist), convergence may happen at one of the local optimal points.
+A common way of finding the global optimum is to carry out a number of replications followed by a selection of the best (lowest) objective value.
+Hence, each trial includes 100 repetitions, and we select the repetition with the lowest objective value as the final result for that trial.
+We average the performance of 50 trials to ensure the rationality of the experiment.
+This means we conduct $50\times 100=5000$ runs for each algorithm.
+100 repetitions in each trial is still within the practical executable range and each trial can give almost consistent results as indicated by a small standard deviation.
+We utilize widely accepted measures of clustering performance for our evaluation indexes, including the normalized mutual information (NMI) [40], the adjusted rand index (ARI) [41], and the clustering accuracy index (ACC) [42].
+Both NMI and ACC range from 0 to 1, with 0 indicating the worst, and 1 representing the best.
+ARI has a range from -1 to 1, in which -1 means two data clusterings are completely dissimilar, 0 indicates that data clusterings are essentially random, and 1 means two data clusterings are perfectly aligned.
+FKM, PFKM, csiFKM, siibFKM, and FWPKM employ a typical fuzzifier value of $m=2$ [43].
+The fuzzifier value of MEFC is set to $\lambda=1$.
+For EKM, we set $\alpha=1$ for the four artificial datasets.
+For the 16 real-world datasets with varying feature numbers, the parameter $\alpha$ is set proportional to the data variance, as follows
+$\alpha=2/\bar{d}_{0n},$ (46)
+where $\bar{d}_{0n}=\frac{1}{2}\sum_{n=1}^{N}\|\mathbf{x}_{n}\|_{2}^{2}/N$, $\mathbf{x}_{n}$ is the $n$-th data point, and $N$ is the total number of data points.
+The reason for such selection is explained in Section IV-C.
+V-B Artificial Datasets
 Fig.
-6 depicts the clustering results.
-V-A4 Imbalanced Fisher’s Iris Data
-Because the instance distribution in the Iris dataset is balanced.
-To examine the capability of imbalanced data clustering, we generate an imbalanced Fisher’s Iris dataset by removing the first 30 instances of Iris setosa, and merging the other two species (Iris virginica and Iris versicolor) into a single class.
-The clustering results of this imbalanced dataset are presented in Fig.
-7.
-V-A5 Wisconsin Diagnostic Breast Cancer
-Wisconsin Diagnostic Breast Cancer (WDBC) [36] comprises 30 features of cell nuclei from diagnostic benign and malignant breast tumors, obtained from digitized images of fine-needle aspiration of breast masses.
-This database contains 569 instances, with an instance distribution of 357 benign and 212 malignant cases.
-We use the first three features (mean radius, mean texture, and mean perimeter) for clustering, each normalized with zero mean and unit variance.
+4 displays scatter plots of the four artificial datasets along with selected data clusterings.
+The average and the standard deviation of evaluation indexes over 50 trials are provided in Table II, where the best and the second-best results are highlighted in bold for easy comparison.
+On Data-A, the data points are sampled from Gaussian distributions.
+While on Data-B the data points are sampled from uniform distributions.
+Data-C and Data-D are a mixture of Gaussian and uniform distributions.
+The difference is that the majority of data points on Data-C are sampled from a Gaussian distribution while the majority of data points on Data-D are sampled from a uniform distribution.
+Additionally, Data-D has more classes than Data-C.
+All four artificial datasets are highly imbalanced with CV values greater than one.
+We can see from Table II that the performance of benchmark algorithms is poor on these artificial datasets, especially on Data-C and Data-D.
+Nevertheless, the proposed EKM is effective, noticeably outperforming the best benchmark algorithm.
+The csiFKM has a certain effectiveness on Data-A and Data-B but does not exceed EKM.
+As seen in Fig.
+4, the benchmark algorithms erroneously divide the majority class into multiple clusters to balance the data size (i.e., the uniform effect) while EKM does not.
+This experiment also proves that EKM is versatile in handling different data distributions.
+Figure 4: Scatter diagrams of artificial datasets.
+(a)-(d) Primitive scatter diagrams with reference class labels (indicated by colors).
+(e)-(f) Clustering results of some benchmark algorithms.
+(i)-(l) Clustering results of the proposed EKM.
+The black crosses represent centroids obtained by each algorithm.
+TABLE II: Experimental results on four artificial datasets with $CV>1$ (the best and second-best performances are in bold)
+Dataset Measurement HKM FKM MEFC PFKM csiFKM siibFKM FWPKM EKM
+Data-A NMI 0.5150 $\pm$0.0001 0.4982 $\pm$0.0001 0.2270 $\pm$0.0002 0.4984 $\pm$0.0001 0.7579 $\pm$0.0471 0.4924 $\pm$0.0503 0.1427 $\pm$0.0488 0.9126 $\pm$0.0000
+ARI 0.2906 $\pm$0.0002 0.2485 $\pm$0.0002 0.0756 $\pm$0.0003 0.2489 $\pm$0.0003 0.8545 $\pm$0.0329 0.2974 $\pm$0.0490 0.0116 $\pm$0.0357 0.9616 $\pm$0.0000
+ACC 0.6893 $\pm$0.0003 0.6075 $\pm$0.0006 0.4818 $\pm$0.0003 0.6089 $\pm$0.0009 0.9429 $\pm$0.0164 0.7443 $\pm$0.0192 0.4562 $\pm$0.0393 0.9933 $\pm$0.0000
+Data-B NMI 0.5193 $\pm$0.0000 0.5160 $\pm$0.0000 0.2121 $\pm$0.0001 0.5181 $\pm$0.0001 0.7989 $\pm$0.0677 0.5090 $\pm$0.0419 0.1591 $\pm$0.0577 0.9981 $\pm$0.0057
+ARI 0.2529 $\pm$0.0000 0.2452 $\pm$0.0000 0.0547 $\pm$0.0000 0.2498 $\pm$0.0003 0.8794 $\pm$0.0406 0.2461 $\pm$0.0486 0.0332 $\pm$0.0497 0.9992 $\pm$0.0023
+ACC 0.6116 $\pm$0.0000 0.5852 $\pm$0.0002 0.4607 $\pm$0.0016 0.6022 $\pm$0.0009 0.9424 $\pm$0.0194 0.6304 $\pm$0.0278 0.4767 $\pm$0.0671 0.$9$$\pm$0.0004
+Data-C NMI 0.0909 $\pm$0.0000 0.0797 $\pm$0.0000 0.0892 $\pm$0.0000 0.0782 $\pm$0.0001 0.1069 $\pm$0.0928 0.0769 $\pm$0.1281 0.0783 $\pm$0.0151 0.9514 $\pm$0.0000
+ARI 0.0181 $\pm$0.0000 0.0034 $\pm$0.0001 0.0159 $\pm$0.0001 0.0015 $\pm$0.0001 0.0728 $\pm$0.1242 0.0199 $\pm$0.1713 0.0053 $\pm$0.0090 0.9807 $\pm$0.0000
+ACC 0.5754 $\pm$0.0000 0.5302 $\pm$0.0002 0.5689 $\pm$0.0002 0.5237 $\pm$0.0002 0.6981 $\pm$0.2108 0.8207 $\pm$0.0436 0.5371 $\pm$0.0272 0.9987 $\pm$0.0000
+Data-D NMI 0.4799 $\pm$0.0316 0.3185 $\pm$0.0010 0.3963 $\pm$0.0670 0.2353 $\pm$0.0554 0.3984 $\pm$0.0440 0.3022 $\pm$0.0729 0.1733 $\pm$0.0268 0.9463 $\pm$0.0104
+ARI 0.1096 $\pm$0.0195 0.0467 $\pm$0.0007 0.1506 $\pm$0.0360 0.0389 $\pm$0.0183 0.0813 $\pm$0.0266 0.0339 $\pm$0.0260 0.0083 $\pm$0.0068 0.9954 $\pm$0.0049
+ACC 0.3593 $\pm$0.0555 0.2559 $\pm$0.0080 0.4893 $\pm$0.0133 0.3699 $\pm$0.0380 0.3357 $\pm$0.0564 0.2330 $\pm$0.0398 0.2206 $\pm$0.0325 0.9763 $\pm$0.0045
+V-C Study of Parameter Impact
+We investigate the influence of the parameter $\alpha$ on EKM in this experiment.
+We perform EKM with $\alpha$ values of $0.1,0.2,0.5,0.8,1,2,5,8,10$ on the four artificial datasets.
 Fig.
-8 presents the clustering results where the mean radius and the mean texture are used as coordinates for visualization.
-V-A6 Imbalanced Wisconsin Diagnostic Breast Cancer
-The data distribution in WDBC is quite balanced.
-To examine the ability of imbalanced data clustering, we remove the first 200 malignant instances in WDBC, resulting in an imbalanced WDBC database with a distribution of 357 benign and 12 malignant instances.
-We still use the first three features for clustering, each normalized with zero mean and unit variance.
-The clustering results are shown in Fig.
-9.
-V-A7 Color-based image segmentation
-The test image coloredChips.png is chosen from Matlab’s image gallery (shown in Fig.
-10a).
-The number of clusters is set to five to expect the segmentation of the four colored chips and the background (i.e., the desk surface).
-Clustering is conducted in the RGB color space and the segmentation results can be viewed in Fig.
-10b - 10e.
-V-B Discussion of Results
-In general, these results demonstrate that the proposed EKM algorithm performs competitively on balanced data and outperforms the other clustering algorithms on imbalanced data.
-In particular, HKM, FKM, and MEFC suffer from the uniform effect when data has an imbalanced distribution.
-The csiFKM and siibFKM algorithms are sensitive to noise data: Fig.
-5c and Fig.
-5d show that their centroids are highly deviated from the correct positions in the presence of noise.
-In comparison, EKM is more robust to noise because of the anchoring effect produced by the positive gradients (see Section IV-B), as seen in Fig.
-5e.
-The strong performance of EKM shown in Fig.
-9 demonstrates its ability to identify clustered anomalies.
-This makes up a missing function of HKM and FKM in anomaly detection: they are typically used to detect isolated anomalies, not clustered anomalies [37, 38, 39, 40].
-The results of color-based image segmentation are shown in Fig.
-10b - 10e.
-HKM, FKM, and MEFC erroneously separate the background (the desk surface) into two clusters.
-This happens because the background comprises most pixels, and splitting it into two clusters better balances the number of pixels within each clusters.
-On the other hand, EKM produces the expected segmentation result.
-Figure 4: (a) Synthetic 2-D ball data. (b) Clustering by HKM. (c) Clustering by FKM. (d) Clustering by MEFC. (e) Clustering by EKM.
-Figure 5: (a) Synthetic 2-D noisy ball data. (b) Clustering by FKM. (c) Clustering by csiFKM. (d) Clustering by siibFKM. (e) Clustering by EKM.
-Figure 6: (a) Fisher’s Iris data. (b) Clustering by HKM. (c) Clustering by FKM. (d) Clustering by MEFC. (e) Clustering by EKM.
-Figure 7: (a) Imbalanced Fisher’s Iris data. (b) Clustering by HKM. (c) Clustering by FKM. (d) Clustering by MEFC. (e) Clustering by EKM.
-Figure 8: (a) Wisconsin Diagnostic Breast Cancer dataset.
-Three features are used for clustering and two of them are used for visualization.
-(b) Clustering by HKM.
-(c) Clustering by FKM.
-(d) Clustering by GMM.
-(e) Clustering by EKM.
-Figure 9: (a) Imbalanced Wisconsin Diagnostic Breast Cancer dataset. (b) Clustering by HKM. (c) Clustering by FKM. (d) Clustering by GMM. (e) Clustering by EKM.
-Figure 10: (a) coloredChips.png. (b) Segmentation by HKM. (c) Segmentation by FKM. (d) Segmentation by MEFC. (e) Segmentation by EKM.
+5 presents the corresponding NMI values for these $\alpha$ values.
+The NMI value of HKM is considered as a reference.
+It is evident that when $\alpha>5$, the NMI of EKM closely aligns with that of HKM.
+This similarity arises because a sufficiently large $\alpha$ makes the objective of EKM nearly identical to the objective of HKM.
+As the value of $\alpha$ decreases, the NMI of EKM increases abruptly, which implies the emergence of repulsive force between centroids overcoming the uniform effect.
+When $\alpha<0.2$, EKM becomes inferior to HKM.
+This is because the centroids of EKM tend to overlap when $\alpha$ is particularly small (repulsion becomes attraction, see Fig.
+3d when $\alpha=0.5$).
+Fortunately, the range of $\alpha$ that causes the centroids to overlap is narrow.
+Figure 5: NMI of EKM with different $\alpha$ values on the four artificial datasets. The blue curve is the NMI of EKM, and the red line is the NMI of HKM as a reference.
+V-D Real Datasets with Balanced Data
+To evaluate EKM’s effectiveness on balanced data, we perform it on six real datasets: IS, Seeds, Heart Disease, Wine, Rice, and WDBC.
+These datasets are routinely used in the field of machine learning and are characterized by uniform class sizes, CV values less than 0.4, and are categorized as balanced datasets as per our protocol.
+Table III shows the clustering results.
+EKM delivers superior performance on the Wine dataset and is comparable to the baseline performance on the remaining five datasets.
+Although EKM’s mechanism makes it more advantageous for imbalanced data, this experiment proves that EKM is competitive for balanced data.
+TABLE III: Experimental results on selected six real-world datasets with $CV<0.4$ (The best performance is in bold)
+Dataset Measurement HKM FKM MEFC PFKM csiFKM siibFKM FWPKM EKM
+IS NMI 0.5873 $\pm$0.0008 0.5950 $\pm$0.0057 0.6206 $\pm$0.0009 0.4962 $\pm$0.0241 0.5195 $\pm$0.0317 0.5925 $\pm$0.0371 0.5452 $\pm$0.0664 0.6618 $\pm$0.0138
+ARI 0.4608 $\pm$0.0005 0.4960 $\pm$0.0059 0.4979 $\pm$0.0005 0.3650 $\pm$0.0149 0.3472 $\pm$0.0587 0.3776 $\pm$0.0794 0.4022 $\pm$0.0884 0.4810 $\pm$0.0089
+ACC 0.5456 $\pm$0.0003 0.6578 $\pm$0.0146 0.5943 $\pm$0.0002 0.5276 $\pm$0.0139 0.4977 $\pm$0.0670 0.5840 $\pm$0.0511 0.5576 $\pm$0.0775 0.5609 $\pm$0.0118
+Seeds NMI 0.7279 $\pm$0.0000 0.7318 $\pm$0.0056 0.7496 $\pm$0.0016 0.7170 $\pm$0.0000 0.7200 $\pm$0.0077 0.6893 $\pm$0.0015 0.5409 $\pm$0.0336 0.7315 $\pm$0.0000
+ARI 0.7733 $\pm$0.0000 0.7768 $\pm$0.0058 0.7968 $\pm$0.0017 0.7607 $\pm$0.0000 0.7640 $\pm$0.0082 0.7143 $\pm$0.0034 0.4724 $\pm$0.0513 0.7715 $\pm$0.0000
+ACC 0.9190 $\pm$0.0000 0.9209 $\pm$0.0023 0.9285 $\pm$0.0007 0.9143 $\pm$0.0000 0.9156 $\pm$0.0033 0.8952 $\pm$0.0015 0.7069 $\pm$0.0649 0.9190 $\pm$0.0000
+Heart Disease NMI 0.3162 $\pm$0.0000 0.0142 $\pm$0.0132 0.2500 $\pm$0.0000 0.2326 $\pm$0.0975 0.1548 $\pm$0.0398 0.0364 $\pm$0.0476 0.0868 $\pm$0.0820 0.2571 $\pm$ 0.0000
+ARI 0.3641 $\pm$0.0000 0.0033 $\pm$0.0042 0.3140 $\pm$0.0000 0.2995 $\pm$0.1230 0.0509 $\pm$0.0172 0.0163 $\pm$0.0523 0.1030 $\pm$0.1050 0.3162 $\pm$ 0.0000
+ACC 0.8020 $\pm$0.0000 0.5249 $\pm$0.0202 0.7805 $\pm$0.0000 0.7653 $\pm$0.0689 0.6101 $\pm$0.0284 0.5354 $\pm$0.0549 0.6344 $\pm$0.0898 0.7815 $\pm$0.0000
+Wine NMI 0.8759 $\pm$0.0000 0.8759 $\pm$0.0000 0.8759 $\pm$0.0000 0.8097 $\pm$0.0171 0.6966 $\pm$0.1663 0.5250 $\pm$0.0803 0.3976 $\pm$0.0794 0.8920 $\pm$0.0000
+ARI 0.8975 $\pm$0.0000 0.8975 $\pm$0.0000 0.8975 $\pm$0.0000 0.8249 $\pm$0.0142 0.6283 $\pm$0.2549 0.3954 $\pm$0.1252 0.3518 $\pm$0.0987 0.9134 $\pm$0.0000
+ACC 0.9663 $\pm$0.0000 0.9663 $\pm$0.0000 0.9663 $\pm$0.0000 0.9396 $\pm$0.0051 0.7797 $\pm$0.1834 0.7129 $\pm$0.0900 0.6476 $\pm$0.0742 0.9719 $\pm$0.0000
+Rice NMI 0.5685 $\pm$0.0000 0.5688 $\pm$0.0000 0.5682 $\pm$0.0004 0.5742 $\pm$0.0008 0.5544 $\pm$0.0001 0.5437 $\pm$0.0032 0.0536 $\pm$0.0268 0.5659 $\pm$0.0005
+ARI 0.6815 $\pm$0.0000 0.6824 $\pm$0.0000 0.6817 $\pm$0.0004 0.6876 $\pm$0.0008 0.6634 $\pm$0.0001 0.6471 $\pm$0.0109 0.0338 $\pm$0.0134 0.6777 $\pm$0.0004
+ACC 0.9129 $\pm$0.0000 0.9131 $\pm$0.0000 0.9129 $\pm$0.0001 0.9147 $\pm$0.0002 0.9074 $\pm$0.0000 0.9023 $\pm$0.0033 0.5941 $\pm$0.0155 0.9117 $\pm$0.0001
+WBDC NMI 0.5547 $\pm$0.0000 0.5612 $\pm$0.0000 0.5547 $\pm$0.0000 0.5700 $\pm$0.0000 0.3516 $\pm$0.2432 0.4202 $\pm$0.1246 0.1138 $\pm$0.0626 0.5513 $\pm$0.0025
+ARI 0.6707 $\pm$0.0000 0.6829 $\pm$0.0000 0.6707 $\pm$0.0000 0.6895 $\pm$0.0000 0.4093 $\pm$0.3070 0.5250 $\pm$0.1734 0.0281 $\pm$0.0409 0.6444 $\pm$0.0028
+ACC 0.9104 $\pm$0.0000 0.9139 $\pm$0.0000 0.9104 $\pm$0.0000 0.9156 $\pm$0.0000 0.7989 $\pm$0.1380 0.8557 $\pm$0.0810 0.5910 $\pm$0.0441 0.9027 $\pm$0.0009
+V-E Real Datasets with Imbalanced Data
+The considered datasets here are Ecoli, Htru2, Zoo, Glass, Shill Bidding, Anuran Calls, Occupancy Detection, Machine Failure, Pulsar Cleaned, and Bert-Embedded Spam.
+Among them, Htru2 and Occupancy Detection have a large amount of data, Glass and Ecoli have 6 and 8 classes respectively, and Zoo has high dimensionality.
+They have varying class sizes with CV values greater than 0.7.
+Except for Occupancy Detection and Zoo, the other eight datasets have CV values greater than 1.
+Bert-Embedded Spam has over 700 features.
+Since the Euclidean distance is ineffective for clustering algorithms on high-dimensional data, we execute principal component analysis on the Bert-Embedded Spam data.
+The first five principal components are used for clustering as adding more does not improve and even compromise performance.
+Table IV displays the results, where the best and the second-best results are in bold.
+As can be seen from Table IV, in general, EKM obtains the best performances on eight datasets, and achieves the second-best clustering performances on the remaining two datasets (Zoo and Anuran Calls), as confirmed by all three measured clustering indexes.
+Specifically, although csiFKM outperforms EKM on the Zoo dataset, it performs substantially worse on Htru2, Occupancy Detection, Shill Bidding, etc.
+HKM performs well on Anuran Calls but underperforms on Shill Bidding, Machine Failure, and Bert-Embedded Spam.
+Due to the structural complexity of real datasets, it is reasonable that some algorithms perform well on certain datasets.
+In comparison, EKM performs consistently well on all datasets, especially on the Htru2, Shill Bidding, Machine Failure, and Bert-Embedded Spam datasets, where it significantly outperforms the benchmark algorithms.
+Fig.
+6 illustrates the scatter diagrams of Htru2, Shill Bidding, Machine Failure, and Bert-Embedded Spam datasets, along with some clustering results.
+Comparing Fig.
+LABEL:sub@fig:htru2_fwpkm - LABEL:sub@fig:spam_mefc with Fig.
+LABEL:sub@fig:htru2_ekm - LABEL:sub@fig:spam_ekm, again, we can see the uniform effect exists in the benchmark algorithms.
+Conversely, the centroids obtained by EKM are farther away from each other, implying the existence of the repulsive force overcoming the uniform effect.
+Note that, we can only see one centroid in some figures because centroids overlap on the selected two coordinates.
+Overall, the experiment results provide strong evidence to support the superiority and consistent performance of EKM on imbalanced data.
+Note that, we simply set $\alpha$ proportional to the data variance.
+EKM’s performance can be further enhanced by fine-tuning the parameter $\alpha$.
+TABLE IV: Experimental results on selected ten real-world datasets with $CV>0.7$ (the best and second-best performances are in bold)
+Dataset Measurement HKM FKM MEFC PFKM csiFKM siibFKM FWPKM EKM
+Zoo NMI 0.8381 $\pm$0.0268 0.7764 $\pm$0.0019 0.8179 $\pm$0.0000 0.6611 $\pm$0.0806 0.8657 $\pm$0.0054 0.7625 $\pm$0.0415 0.6652 $\pm$0.0804 0.7912 $\pm$0.0188
+ARI 0.7546 $\pm$0.0487 0.6235 $\pm$0.0010 0.6444 $\pm$0.0000 0.4868 $\pm$0.0829 0.8715 $\pm$0.0034 0.6423 $\pm$0.1030 0.4743 $\pm$0.0948 0.8181 $\pm$0.0783
+ACC 0.8248 $\pm$0.0310 0.6832 $\pm$0.0000 0.7228 $\pm$0.0000 0.6267 $\pm$0.0536 0.8703 $\pm$0.0078 0.7305 $\pm$0.0831 0.5958 $\pm$0.0604 0.8507 $\pm$0.0566
+Glass NMI 0.3140 $\pm$0.0046 0.3073 $\pm$0.0003 0.3120 $\pm$0.0012 0.0721 $\pm$0.0069 0.3390 $\pm$0.0549 0.3424 $\pm$0.0413 0.1654 $\pm$0.0823 0.3764 $\pm$0.0370
+ARI 0.1702 $\pm$0.0026 0.1529 $\pm$0.0005 0.1651 $\pm$0.0009 0.0070 $\pm$0.0059 0.1930 $\pm$0.0435 0.1648 $\pm$0.0366 0.0734 $\pm$0.0659 0.1978 $\pm$0.0193
+ACC 0.4586 $\pm$0.0076 0.4021 $\pm$0.0011 0.4487 $\pm$0.0007 0.3368 $\pm$0.0054 0.4684 $\pm$0.0476 0.4202 $\pm$0.0270 0.4096 $\pm$0.0584 0.4796 $\pm$0.0086
+Ecoli NMI 0.6379 $\pm$0.0040 0.5695 $\pm$0.0115 0.6096 $\pm$0.0076 0.5012 $\pm$0.0649 0.5985 $\pm$0.0241 0.5541 $\pm$0.0464 0.4746 $\pm$0.0424 0.6426 $\pm$0.0024
+ARI 0.5032 $\pm$0.0070 0.4142 $\pm$0.0204 0.4815 $\pm$0.0270 0.3357 $\pm$0.0752 0.4807 $\pm$0.0898 0.3697 $\pm$0.0862 0.2832 $\pm$0.0500 0.5157 $\pm$0.0013
+ACC 0.6461 $\pm$0.0103 0.5769 $\pm$0.0160 0.6238 $\pm$0.0218 0.5302 $\pm$0.0435 0.6276 $\pm$0.0711 0.5624 $\pm$0.0638 0.4598 $\pm$0.0410 0.6482 $\pm$0.0043
+Htru2 NMI 0.4068 $\pm$0.0000 0.4100 $\pm$0.0004 0.4075 $\pm$0.0002 0.1289 $\pm$0.0002 0.0204 $\pm$0.0506 0.3671 $\pm$0.0290 0.0660 $\pm$0.0268 0.5872 $\pm$0.0003
+ARI 0.6071 $\pm$0.0000 0.5796 $\pm$0.0005 0.6075 $\pm$0.0002 0.0462 $\pm$0.0002 0.0077 $\pm$0.0649 0.5704 $\pm$0.0379 -0.0117 $\pm$0.0074 0.7333 $\pm$0.0002
+ACC 0.9366 $\pm$0.0000 0.9252 $\pm$0.0001 0.9366 $\pm$0.0000 0.6091 $\pm$0.0002 0.$8$$\pm$0.0637 0.9337 $\pm$0.0126 0.5170 $\pm$0.0151 0.9661 $\pm$0.0000
+Shill Bidding NMI 0.0021 $\pm$0.0001 0.0023 $\pm$0.0000 0.0042 $\pm$0.0002 0.0012 $\pm$0.0000 0.0295 $\pm$0.0359 0.0351 $\pm$0.0047 0.0105 $\pm$0.0330 0.6216 $\pm$0.0000
+ARI -0.0006 $\pm$0.$0$0.0002 $\pm$0.$0$0.0002 $\pm$0.$0$0.0002 $\pm$0.$0$0.0485 $\pm$0.0482 -0.0854 $\pm$0.0118 0.0079 $\pm$0.0252 0.7914 $\pm$0.0000
+ACC 0.5020 $\pm$0.0002 0.5059 $\pm$0.0000 0.5090 $\pm$0.0003 0.5035 $\pm$0.0002 0.7457 $\pm$0.0961 0.7869 $\pm$0.0182 0.5937 $\pm$0.0750 0.9674 $\pm$0.0000
+Anuran Calls NMI 0.6775 $\pm$0.0195 0.5699 $\pm$0.0002 0.6155 $\pm$0.0022 0.4270 $\pm$0.0153 0.5575 $\pm$0.0516 0.5178 $\pm$0.0339 0.5159 $\pm$0.0247 0.6229 $\pm$0.0134
+ARI 0.5826 $\pm$0.0174 0.3414 $\pm$0.0009 0.4017 $\pm$0.0027 0.2254 $\pm$0.0694 0.4329 $\pm$0.1363 0.2368 $\pm$0.0694 0.4268 $\pm$0.0500 0.5145 $\pm$0.0329
+ACC 0.6441 $\pm$0.0134 0.4363 $\pm$0.0009 0.5103 $\pm$0.0073 0.3705 $\pm$0.0230 0.5682 $\pm$0.0799 0.4101 $\pm$0.0348 0.5123 $\pm$0.0271 0.5821 $\pm$0.0420
+Occupancy Detection NMI 0.4890 $\pm$0.0001 0.4720 $\pm$0.0003 0.4710 $\pm$0.0002 0.0539 $\pm$0.0007 0.0934 $\pm$0.1560 0.2982 $\pm$0.1379 0.0591 $\pm$0.0464 0.5389 $\pm$0.0002
+ARI 0.5959 $\pm$0.0001 0.5699 $\pm$0.0004 0.5670 $\pm$0.0002 0.0209 $\pm$0.0002 0.0521 $\pm$0.2248 0.2175 $\pm$0.2226 0.0363 $\pm$0.0682 0.6750 $\pm$0.0002
+ACC 0.8912 $\pm$0.0000 0.8825 $\pm$0.0001 0.8815 $\pm$0.0001 0.5772 $\pm$0.0004 0.7597 $\pm$0.0596 0.7026 $\pm$0.1361 0.6117 $\pm$0.0595 0.9163 $\pm$0.0001
+Machine Failure NMI 0.0248 $\pm$0.0000 0.0281 $\pm$0.0000 0.0310 $\pm$0.0000 0.0284 $\pm$0.0003 0.2285 $\pm$0.0569 0.0540 $\pm$0.0292 0.0137 $\pm$0.0161 0.5825 $\pm$0.1869
+ARI -0.0116 $\pm$0.$0$0.0083 $\pm$0.$0$0.0052 $\pm$0.0000 0.0007 $\pm$0.0001 0.2217 $\pm$0.0510 -0.0037 $\pm$0.0382 0.0031 $\pm$0.0083 0.6506 $\pm$0.1987
+ACC 0.5673 $\pm$0.0000 0.5418 $\pm$0.0002 0.5202 $\pm$0.0001 0.5146 $\pm$0.0007 0.9653 $\pm$0.0543 0.6125 $\pm$0.0412 0.5481 $\pm$0.0358 0.9863 $\pm$0.0068
+Pulsar Cleaned NMI 0.0311 $\pm$0.0002 0.0167 $\pm$0.0000 0.0201 $\pm$0.0000 0.0236 $\pm$0.0011 0.0523 $\pm$0.0813 0.0044 $\pm$0.0081 0.0090 $\pm$0.0067 0.0544 $\pm$0.0001
+ARI 0.0370 $\pm$0.0003 0.0069 $\pm$0.0000 0.0131 $\pm$0.0001 -0.0011 $\pm$0.0002 0.0731 $\pm$0.1193 -0.0067 $\pm$0.0160 -0.0039 $\pm$0.0080 0.1082 $\pm$0.0002
+ACC 0.7329 $\pm$0.0008 0.5714 $\pm$0.0001 0.6171 $\pm$0.0004 0.5019 $\pm$0.0012 0.9488 $\pm$0.0973 0.9043 $\pm$0.1231 0.6036 $\pm$0.0659 0.8834 $\pm$0.0002
+Bert-Embedded Spam NMI 0.1517 $\pm$0.0001 0.1249 $\pm$0.0000 0.0472 $\pm$0.0374 0.1122 $\pm$0.0023 0.0330 $\pm$0.0791 0.0431 $\pm$0.0946 0.0238 $\pm$0.0004 0.6704 $\pm$0.0000
+ARI 0.0830 $\pm$0.0003 0.0550 $\pm$0.0000 0.0289 $\pm$0.0265 -0.0026 $\pm$0.0007 0.0080 $\pm$0.1084 0.0296 $\pm$0.1287 -0.0669 $\pm$0.0004 0.8216 $\pm$0.0000
+ACC 0.6453 $\pm$0.0002 0.6174 $\pm$0.0000 0.5837 $\pm$0.0390 0.5482 $\pm$0.0012 0.8210 $\pm$0.0816 0.8281 $\pm$0.0832 0.8038 $\pm$0.0002 0.9664 $\pm$0.0000
+Figure 6: Scatter diagrams of some selected real imbalanced datasets (from left to right: Htru2, Shill Bidding, Machine Failure, and Bert-Embedded Spam).
+The first two features are used as coordinates for visualization.
+(a)-(d) Primitive scatter diagrams with reference class labels.
+(e)-(f) Clustering results of some benchmark algorithms.
+(i)-(l) Clustering results of the proposed EKM.
+The black crosses represent centroids obtained by each algorithm.
+V-F Empirical Convergence and Algorithm Efficiency
+Table V shows the average number of iterations and the average time per run (each algorithm is run 5000 times).
+The numbers in brackets in the table are the ranking numbers of the corresponding algorithms.
+AVK and MDK give the average and median number of iterations and computational time of each algorithm on all test datasets, respectively.
+As evident from Table V, AVK and MDK are consistent except that EKM is faster than FKM in AVK and slower in MDK.
+This is because FKM is slow on the Bert-Embedded Spam dataset.
+HKM is the fastest algorithm as it does not require membership calculation.
+MEFC comes second, followed by FKM and EKM.
+EKM is 30% faster than FKM in AVK and slower than FKM by 9% in MDK.
+In terms of absolute speed, MEFC, FKM, and EKM operate within the same order of magnitude.
+This aligns with their identical time complexity of $O(NK^{2}P)$ per iteration, where $N$ is the instance number, $K$ is the cluster number, and $P$ is the data dimension or feature number.
+The table also demonstrates the empirical convergence of EKM, averaging 44 iterations or a median of 20 iterations to reach the specified convergence (45).
+In summary, EKM has a low computational cost.
+TABLE V: Average number of iterations and average calculation time of each algorithm
+Dataset Measurement HKM FKM MEFC PFKM csiFKM siibFKM FWPKM EKM
+Data-A Iter 17.3 13.7 29.0 33.5 30.7 28.0 44.6 12.1
+Time 0.005(1) 0.008(3) 0.013(4) 0.033(7) 0.021(6) 0.277(8) 0.016(5) 0.007(2)
+Data-B Iter 9.2 13.2 30.1 28.8 15.1 17.4 24.1 17.5
+Time 0.003(1) 0.008(2) 0.013(6) 0.031(7) 0.011(5) 0.168(8) 0.010(4) 0.009(3)
+Data-C Iter 16.3 17.6 36.0 23.9 35.0 42.1 73.7 23.5
+Time 0.008(1) 0.014(2) 0.023(4) 0.048(7) 0.040(6) 0.624(8) 0.035(5) 0.017(3)
+Data-D Iter 14.8 23.6 25.4 39.8 24.5 25.2 7.3 15.6
+Time 0.033(2) 0.091(5) 0.082(4) 0.364(7) 0.142(6) 1.867(8) 0.030(1) 0.058(3)
+IS Iter 15.9 40.1 20.9 53.0 49.8 32.1 29.6 24.1
+Time 0.021(1) 0.110(5) 0.058(2) 0.218(7) 0.104(4) 0.824(8) 0.152(6) 0.074(3)
+Seeds Iter 8.5 10.4 11.1 13.3 12.6 26.0 32.5 12.0
+Time 0.001(1) 0.001(3) 0.001(2) 0.003(6) 0.002(4) 0.028(8) 0.005(7) 0.002(5)
+Heart Disease Iter 9.9 269.3 12.5 22.7 500.0 498.8 3.3 18.7
+Time 0.002(1) 0.074(5) 0.003(3) 0.081(6) 2.095(8) 1.646(7) 0.002(2) 0.005(4)
+Wine Iter 7.0 13.5 10.6 25.3 15.7 40.4 19.4 11.7
+Time 0.000(1) 0.002(4) 0.001(2) 0.004(6) 0.002(5) 0.037(8) 0.005(7) 0.001(3)
+Rice Iter 8.4 8.7 8.9 12.7 14.0 17.8 72.8 9.1
+Time 0.004(1) 0.009(4) 0.006(2) 0.021(6) 0.011(5) 0.214(8) 0.072(7) 0.009(3)
+WDBC Iter 7.4 10.8 10.6 32.4 500.0 500.0 16.2 9.9
+Time 0.001(1) 0.003(4) 0.002(2) 0.009(5) 2.307(8) 0.933(7) 0.013(6) 0.002(3)
+Zoo Iter 4.4 31.0 23.8 35.0 34.4 37.8 2.3 498.9
+Time 0.001(1) 0.005(4) 0.003(3) 0.009(6) 0.006(5) 0.048(7) 0.001(2) 0.069(8)
+Glass Iter 8.5 27.2 29.4 18.8 31.4 45.0 25.7 25.3
+Time 0.001(1) 0.006(5) 0.006(2) 0.009(7) 0.006(4) 0.096(8) 0.007(6) 0.006(3)
+Ecoli Iter 14.1 42.9 23.6 32.2 40.5 45.2 3.0 22.9
+Time 0.002(2) 0.019(6) 0.010(3) 0.029(7) 0.015(5) 0.197(8) 0.002(1) 0.011(4)
+Htru2 Iter 12.7 33.8 15.2 21.5 500.0 500.0 52.5 10.8
+Time 0.046(1) 0.175(4) 0.067(3) 0.357(5) 17.884(7) 27.836(8) 0.450(6) 0.059(2)
+Shill Bidding Iter 12.0 15.0 24.3 22.7 36.8 40.4 25.6 22.3
+Time 0.011(1) 0.026(2) 0.029(3) 0.064(7) 0.046(5) 0.794(8) 0.056(6) 0.033(4)
+Anuran Calls Iter 22.5 35.7 24.1 39.3 42.6 45.4 24.0 44.3
+Time 0.339(1) 0.527(3) 0.377(2) 1.502(6) 0.614(4) 5.119(8) 2.089(7) 0.719(5)
+Occupancy Detection Iter 11.5 25.6 28.6 24.6 28.0 194.9 19.8 14.8
+Time 0.032(1) 0.132(5) 0.097(3) 0.291(7) 0.109(4) 12.090(8) 0.163(6) 0.067(2)
+Machine Failure Iter 13.4 22.0 11.9 127.4 422.1 34.5 122.6 22.5
+Time 0.015(1) 0.055(4) 0.022(2) 0.442(6) 0.724(7) 1.058(8) 0.331(5) 0.049(3)
+Pulsar Cleaned Iter 25.3 17.1 18.4 60.7 500.0 500.0 50.2 21.4
+Time 0.045(1) 0.068(3) 0.052(2) 0.435(6) 12.560(7) 22.953(8) 0.268(5) 0.075(4)
+Bert-Embedded Spam Iter 31.6 500.0 498.7 21.2 500.0 500.0 28.5 41.6
+Time 0.015(1) 0.618(5) 0.388(4) 0.642(6) 4.866 (7) 8.520(8) 0.030(2) 0.045(3)
+AVK Iter 13.5 58.5 44.7 34.4 166.4 158.3 33.9 44.0
+Time 0.029(1) 0.098(4) 0.063(2) 0.230(6) 2.078(7) 4.266(8) 0.187(5) 0.066(3)
+MDK Iter 12.4 22.8 23.7 27.1 35.9 41.3 25.6 20.0
+Time 0.007(1) 0.023(3) 0.018(2) 0.056(6) 0.075(7) 0.809(8) 0.030(5) 0.025(4)
 VI Deep Clustering
-Clustering algorithms using Euclidean distance for defining similarity face a challenge when clustering high-dimensional data, such as images, due to the small distance differences between various point pairs in high-dimensional space [41].
+Clustering algorithms using Euclidean distance for defining similarity face a challenge when clustering high-dimensional data, such as images, due to the distance differences between point pairs vanishes in high-dimensional space [44].
 Deep clustering is a technique that addresses this issue by employing DNNs to map high-dimensional data to low-dimensional representation.
@@
 However, this design is less favorable to imbalanced data [22].
-In this section, we show the promise of EKM in the deep clustering of imbalanced data by replacing HKM with EKM in a popular deep clustering framework.
+In this section, we show that EKM is a better alternative for deep clustering of imbalanced data.
 VI-A Optimization Procedure
@@
 As illustrated in Fig.
-11, DCN maps high-dimensional data to low-dimensional representation through an autoencoder network.
+7, DCN maps high-dimensional data to low-dimensional representation through an autoencoder network.
 An autodecoder follows the autoencoder, mapping the representation back to the original high-dimensional space (i.e., reconstruction).
@@
 where $m_{k}^{n}$ is the number of samples assigned to the $k$-th cluster before the incoming data $\mathbf{x}_{n}$, controlling the learning rate of the $k$-th centroid.
-Overall, the optimization procedure of DCN alternates between updating networks parameters $\{\bm{\theta}_{e},\bm{\theta}_{d}\}$ by solving (45) and updating HKM parameters $\{\mathbf{C},\{\mathbf{s}_{n}\}\}$ by (46) and (47).
+Overall, the optimization procedure of DCN alternates between updating network parameters $\{\bm{\theta}_{e},\bm{\theta}_{d}\}$ by solving (47) and updating HKM parameters $\{\mathbf{C},\{\mathbf{s}_{n}\}\}$ by (48) and (49).
 However, the centroid updating rule (49) is problematic for imbalanced data due to the uniform effect. To address this issue, we propose to replace (49) with the batch-learning version of EKM:
@@
 In particular, the dimension of the low-dimensional representation is set to ten, and the parameter $\beta$ is set to one.
-The smoothing parameter $\alpha$ is tuned on the representation obtained by the initialized DCN networks according to the strategy introduced in Section IV-C.
-We evaluate clustering performance using three widely accepted metrics: including normalized mutual information (NMI), adjusted Rand index (ARI), and clustering accuracy (ACC).
-NMI and ACC range from 0 to 1, with zero representing the worst and one representing the best.
-ARI ranges from -1 to 1, with minus one representing the worst and one representing the best.
-We first evaluate the algorithm’s performance on the full MNIST dataset [42], which contains 70,000 gray images of handwritten digits from 0 to 9.
-Each image has $28\times 28=784$ pixels.
+The smoothing parameter $\alpha$ is tuned on the representation obtained by the initialized DCN network according to the strategy introduced in Section IV-C.
+We first evaluate the algorithm’s performance on a balanced dataset.
+The considered dataset is the full MNIST [45], which contains 70,000 gray images of handwritten digits from 0 to 9.
+Each digit has approximately 7,000 images and each image has $28\times 28=784$ pixels.
 We set the smoothing parameter of EKM to $\alpha=5\mathrm{e}-3$.
-The clustering results are presented in TABLE I.
+The clustering results are presented in TABLE VI.
 We compare the proposed DCN+EKM with DCN+HKM and stacked autoencoder (SAE).
@@
 We can also see that DCN+EKM performs similarly to DCN+HKM.
-We omit the results of DCN+FKM and DCN+MEFC due to space limitations.
+We omit the results of DCN+FKM and DCN+MEFC due to save space.
 Their performance is not better than DCN+HKM.
-MNIST contains 60,000 training images and 10,000 testing images.
-To evaluate the performance of imbalanced data clustering, we generate an imbalanced MNIST dataset by removing the training images of digits 1 to 9.
-This imbalanced dataset contains 15,923 images, of which approximately 6,900 are digit 0, while the remaining digits (1 to 9) each have about 1,000 images.
+Then we evaluate the algorithm performance on an imbalanced dataset.
+The considered dataset is derived from the full MNIST by removing the training images of digits 1 to 9.
+This imbalanced dataset (we call imbalanced MNIST) contains 15,923 images, of which approximately 7,000 are digit 0, while the remaining digits (1 to 9) each have about 1,000 images.
 We set the smoothing parameter of EKM to $\alpha=3.8\mathrm{e}-3$.
-The results are summarized in TABLE II.
-In conclusion, we find that an EKM-friendly representation structure is crucial for deep clustering of imbalanced data.
+The results are summarized in TABLE VII.
+We can see that DCN+EKM greatly outperforms other algorithms.
+This finding implies that the EKM-friendly structure is crucial for deep clustering of imbalanced data.
 We map the ten-dimensional representation obtained by DCN to a two-dimensional space by t-SNE [46] for visualization.
-The results are displayed in Fig.
-12.
-The visualization result implies that the deep representation obtained by DCN+EKM is more discriminative than that obtained by DCN+HKM.
-We observe that EKM successfully identifies the large class (number 0) from other small classes without referring to the true labels, while HKM incorrectly divides the large class into four clusters.
-TABLE I: Evaluation on full MNIST
+The representation learned on the full MNIST dataset is displayed in Fig.
+8 and the representation learned on the imbalanced MNIST dataset is displayed in Fig.
+9.
+The visualization results suggest that the deep representation obtained by DCN+EKM is more discriminative than that obtained by DCN+HKM as the former has clearer inter-cluster margins, especially for those obtained from the imbalanced MNIST dataset.
+This should be attributed to the centroid repulsion mechanism of EKM, which makes data farther away in representation space.
+We also observe from Fig.
+9 that EKM successfully identifies the large class (digit 0) from other small classes without referring to the true labels, while HKM incorrectly divides the large class into four clusters to balance the cluster sizes.
+TABLE VI: Evaluation on full MNIST
 Methods SAE+HKM DCN+HKM SAE+EKM DCN+EKM NMI 0.725 0.798 0.711 0.813 ARI 0.667 0.744 0.642 0.731 ACC 0.795 0.837 0.782 0.808
-TABLE II: Evaluation on imbalanced MNIST
+TABLE VII: Evaluation on imbalanced MNIST
 Methods SAE+HKM DCN+HKM SAE+EKM DCN+EKM NMI 0.551 0.584 0.583 0.701 ARI 0.317 0.325 0.396 0.826 ACC 0.413 0.434 0.497 0.784
-Figure 12: Visualization of DCN representations by t-SNE.
-(a) Representations obtained by DCN+HKM, colored by true labels.
+Figure 8: The t-SNE visualization of DCN representations learned on the full MNIST dataset.
+(a) Representations obtained by DCN+HKM, colored by reference labels.
 (b) Representations obtained by DCN+HKM, colored by labels generated by HKM.
-(c) Representations obtained by DCN+EKM, colored by true labels.
+(c) Representations obtained by DCN+EKM, colored by reference labels.
+(d) Representations obtained by DCN+EKM, colored by labels generated by EKM.
+Figure 9: The t-SNE visualization of DCN representations learned on the imbalanced MNIST dataset.
+(a) Representations obtained by DCN+HKM, colored by reference labels.
+(b) Representations obtained by DCN+HKM, colored by labels generated by HKM.
+(c) Representations obtained by DCN+EKM, colored by reference labels.
 (d) Representations obtained by DCN+EKM, colored by labels generated by EKM.
 VII Conclusion
-This paper presents equilibrium K-means (EKM), a novel clustering algorithm effective for both balanced and imbalanced data.
-EKM is simple, interpretable, scalable to large datasets, and capable of creating unequal-sized clusters.
-Experimental results on real datasets from various domains show that EKM outperforms HKM, FKM, and other popular centroid-based algorithms on datasets with imbalanced data and performs comparably on datasets with balanced data.
-In the same deep clustering framework, compared with HKM and FKM, the clustering accuracy of imbalanced data using EKM has been noticeably improved.
+This paper presents equilibrium K-means (EKM), a novel clustering algorithm with a new centroid repulsion mechanism effective for imbalanced data.
+EKM is simple, interpretable, and scalable to large datasets.
+Experimental results on datasets from various domains show that EKM outperforms HKM, FKM, and other state-of-the-art centroid-based algorithms on imbalanced datasets and performs comparably on balanced datasets.
+We also demonstrate that EKM is a better alternative to HKM and FKM in deep clustering when dealing with imbalanced data.
 Furthermore, we reformulate HKM, FKM, and EKM in a general form of gradient descent.
-We encourage readers to study more properties of this general form to develop new K-means algorithms with non-heuristic memberships to address different issues.
+We encourage readers to study more properties of this general form.
+Combining EKM with kernels, multi-prototype mechanisms, and other techniques to deal with more complex data structures is also an important research direction.
+We believe that these explorations will promote the development of EKM and further solve the data imbalance issue.
 Appendix A Proof of Theorem 1
-For concise and tidy, we denote the partial derivative $\frac{\partial h}{\partial d_{k}}(d_{1n}^{(\tau)},\cdots,d_{Kn}^{(\tau)})$ as $\frac{\partial h}{\partial d_{k}}|_{n,{\tau}}$.
+For concise and tidy, we denote the partial derivative $\frac{\partial h}{\partial d_{k}}(d_{1n}^{(\tau)},\cdots,d_{Kn}^{(\tau)})$ as $\frac{\partial h}{\partial d_{k}}|_{n,{\tau}}$, $h(d_{1n},\cdots,d_{Kn})$ as $h$, and $h(d_{1n}^{(\tau)},\cdots,d_{Kn}^{(\tau)})$ as $h^{(\tau)}$.
 Since $h$ is a concave function at $[0,\,+\infty)^{K}$, we have
-$h(d_{1n},\cdots,d_{Kn})$ (49)
-$\leq h(d_{1n}^{(\tau)},\cdots,d_{Kn}^{(\tau)})+\sum_{k=1}^{K}\frac{\partial h}{\partial d_{k}}\bigg{|}_{n,{\tau}}\cdot(d_{kn}-d_{kn}^{(\tau)}),$
+$h\leq h^{(\tau)}+\sum_{k=1}^{K}\frac{\partial h}{\partial d_{k}}\bigg{|}_{n,{\tau}}\cdot(d_{kn}-d_{kn}^{(\tau)}),$ (51)
 which holds for any $n\in\{1,\cdots,N\}$ and $d_{1n},\cdots,d_{Kn}\in[0,\,+\infty)$. Summing over $n$, it follows that
-$J(\mathbf{c}_{1},\cdots,\mathbf{c}_{K})\\ \leq\sum_{n=1}^{N}\bigg{[}h(d_{1n}^{(\tau)},\cdots,d_{Kn}^{(\tau)})+\sum_{k=1}^{K}\frac{\partial h}{\partial d_{k}}\bigg{|}_{n,{\tau}}\cdot(d_{kn}-d_{kn}^{(\tau)})\bigg{]}\\ =J(\mathbf{c}_{1}^{(\tau)},\cdots,\mathbf{c}_{K}^{(\tau)})+\frac{1}{2}\sum_{k=1}^{K}\sum_{n=1}^{N}\bigg{(}\frac{\partial h}{\partial d_{k}}\bigg{|}_{n,{\tau}}\cdot\big{(}\|\mathbf{x}_{n}-\mathbf{c}_{k}\|_{2}^{2}\\ -\|\mathbf{x}_{n}-\mathbf{c}_{k}^{(\tau)}\|_{2}^{2}\big{)}\bigg{)}.$ (50)
+$J(\mathbf{c}_{1},\cdots,\mathbf{c}_{K})\leq\sum_{n=1}^{N}\bigg{[}h^{(\tau)}+\sum_{k=1}^{K}\frac{\partial h}{\partial d_{k}}\bigg{|}_{n,{\tau}}\cdot(d_{kn}-d_{kn}^{(\tau)})\bigg{]}\\ =J(\mathbf{c}_{1}^{(\tau)},\cdots,\mathbf{c}_{K}^{(\tau)})+\frac{1}{2}\sum_{k=1}^{K}\sum_{n=1}^{N}\bigg{(}\frac{\partial h}{\partial d_{k}}\bigg{|}_{n,{\tau}}\cdot\big{(}\|\mathbf{x}_{n}-\mathbf{c}_{k}\|_{2}^{2}\\ -\|\mathbf{x}_{n}-\mathbf{c}_{k}^{(\tau)}\|_{2}^{2}\big{)}\bigg{)}.$ (52)
 Denote the function on the right side of the inequality as $M(\mathbf{c}_{1},\cdots,\mathbf{c}_{K})$.
@@
 Because $(J^{(1)},J^{(2)},\cdots)$ is non-increasing, only (local) minimizers or saddle points appear as limit points.
-Supplemental Materials: Imbalanced Data Clustering using Equilibrium K-Means
-supplement
-We evaluate the clustering performance of the proposed equilibrium K-means (EKM) on ten real datasets.
-Four datasets (Iris, Imbalanced Iris, Wisconsin Diagnostic Breast Cancer (WDBC), and Imbalanced WDBC) are introduced in detail in the main text.
-Herein we supplement the results for the remaining six datasets.
-They are from the UCI repository (available at https://archive.ics.uci.edu/datasets), which are: Wine, Ecoli, Image Segmentation (IS), Htru2, Rice, and Dry Bean.
-A detailed description of all ten datasets are summarized in TABLE SI, ordered according to the imbalance level measured by CV0, which we will define later.
-The dataset with a lower order has a higher CV0 value, indicating a higher imbalance level.
-In the “Feature” column of TABLE SI, the format “x (y)” indicates that the dataset has y features, of which x features are used for clustering.
-For instance, the Iris dataset has four features, of which two are used for clustering.
-Partial features are used for the sake of visualizing clustering results.
-In the “CV0” column, we calculate the coefficient of variation (CV) of the dataset to measure the level of data imbalance.
-CV is defined as the ratio of the standard deviation to the mean.
-Mathematically, given the number of instances in each class as $N_{1},\cdots,N_{K}$, we have
-$\text{CV}=\text{std}(\mathbf{N})/\text{mean}(\mathbf{N}),$
-where
-$\text{mean}(\mathbf{N})=\frac{\sum_{k=1}^{K}N_{k}}{K},$
-and
-$\text{std}(\mathbf{N})=\sqrt{\frac{\sum_{k=1}^{K}(N_{k}-\text{mean}(\mathbf{N}))^{2}}{K-1}}.$
-In general, the larger the CV value, the greater the imbalance of the data.
-Each dataset is normalized so that each feature has zero mean and unit variance.
-The number of clusters is assumed to be equal to the number of classes for all datasets.
-For the supplementary six datasets, the smoothness parameter $\alpha$ is chosen adaptively as follows:
-$\alpha=2/\bar{d}_{0n},$
-where $\bar{d}_{0n}=\frac{1}{2}\sum_{n=1}^{N}\|\mathbf{x}_{n}\|_{2}^{2}/N$, $\mathbf{x}_{n}$ is the $n$-th data point, and $N$ is the total number of data points.
-The value $\bar{d}_{0n}$ can be interpreted as a data variance after the normalization of zero data mean.
-The normalized mutual information (NMI), adjusted Rand index (ARI), and clustering accuracy (ACC) of each algorithm on the ten datasets are presented in TABLE SII.
-Bold values indicate the best results.
-We also include three metrics: “CV1”, ”DCV” and ”Running time” for readers’ interest.
-“CV1” is the CV value of the clustering result.
-“DCV” is the difference between CV0 and CV1, i.e., $\text{DCV}=\text{CV}0-\text{CV}1$.
-A positive DCV value indicates that the data distribution of the true classes is more imbalanced than that of the generated clusters, while a negative DCV value means the data distribution of the true classes is more uniform than that of the generated clusters.
-“Running time” is the sum of the running time of ten replicates.
-We observe that the proposed EKM generally outperforms the other six algorithms on the last four datasets with CV values greater than 0.5.
-EKM outperforms other algorithms on two of the top six datasets with CV values less than 0.5.
-On the rest of the datasets, EKM still has a competitive performance.
-Overall, these results demonstrate that EKM is an efficient clustering algorithm for both balanced and imbalanced data.
-TABLE SI: Detailed description of REAL-WORLD Datasets
-Dataset Instance Feature Class CV0
-Iris 150 2 (4) 3 0
-IS 2310 19 (19) 7 0
-Wine 178 13 (13) 3 0.1939
-Rice 3810 7 (7) 2 0.2042
-WDBC 569 3 (30) 3 0.3604
-Dry Bean 13611 16 (16) 7 0.4951
-Imbalanced Iris 120 2 (4) 2 0.9428
-Ecoli 336 7 (7) 8 1.1604
-Imbalanced WDBC 369 3 (30) 2 1.3222
-Htru2 17898 8 (8) 2 1.4142
-TABLE SII: Experimental results of different algorithms on real-world datasets
-Dataset Measurement HKM FKM MEFC csiFKM siibFKM GMM EKM
-Iris NMI 0.5945 0.5671 0.5194 0.5671 0.5675 0.6775 0.5457
-ARI 0.5462 0.5328 0.5291 0.5328 0.5281 0.5354 0.5134
-ACC 0.7867 0.7800 0.7933 0.7800 0.7800 0.7000 0.7733
-CV1 0.0346 0.0200 0.1217 0.0200 0.1442 0.8902 0.0346
-DCV -0.0346 -$0$0.1217 -$0$0.1442 -0.8902 -0.0346
-Running time 0.0113 0.0111 0.0305 0.0142 0.1844 0.0715 0.0328
-IS NMI 0.5864 0.5953 0.6222 0.5971 0.5997 0.5657 0.6463
-ARI 0.4605 0.4973 0.4995 0.4962 0.3214 0.4456 0.5161
-ACC 0.5455 0.6455 0.5944 0.6433 0.5861 0.5641 0.5944
-CV1 0.8246 0.2450 0.6433 0.4617 1.0377 0.5037 0.6582
-DCV -0.8246 -0.2450 -0.6433 -0.4617 -1.0377 -0.5037 -0.6582
-Running time 0.0321 0.4714 0.2616 0.7913 49.5369 1.0463 0.7205
-Wine NMI 0.8759 0.8759 0.8759 0.8610 0.3361 0.8466 0.8920
-ARI 0.8975 0.8975 0.8975 0.8804 0.1355 0.8649 0.9134
-ACC 0.9663 0.9663 0.9663 0.9607 0.5562 0.9551 0.9719
-CV1 0.1242 0.1242 0.1242 0.1219 1.1513 0.1070 0.1403
-DCV 0.0696 0.0696 0.0696 0.0720 -0.9574 0.0868 0.0535
-Running time 0.0075 0.0153 0.0088 0.0126 1.6469 0.0515 0.0094
-Rice NMI 0.5685 0.5688 0.5699 0.5554 0.5415 0.4891 0.5653
-ARI 0.6815 0.6824 0.6833 0.6642 0.6394 0.5649 0.6772
-ACC 0.9129 0.9131 0.9134 0.9076 0.9000 0.8759 0.9115
-CV1 0.2442 0.2183 0.2309 0.2977 0.3623 0.0163 0.2643
-DCV -0.0401 -0.0141 -0.0267 -0.0935 -0.1581 0.1878 -0.0601
-Running time 0.0200 0.0438 0.0357 0.1032 2.1275 0.2303 0.0397
-WBDC NMI 0.4920 0.4870 0.5005 0.4423 0.2951 0.4618 0.4906
-ARI 0.5844 0.5963 0.6257 0.4402 0.2797 0.5722 0.5340
-ACC 0.8840 0.8875 0.8963 0.8366 0.7750 0.8787 0.8682
-CV1 0.5990 0.5393 0.4051 0.8227 0.9768 0.2461 0.7133
-DCV -0.2386 -0.1790 -0.0447 -0.4623 -0.6164 0.1143 -0.3529
-Running time 0.0156 0.0226 0.0171 0.0385 1.7632 0.0894 0.0160
-Dry Bean NMI 0.7138 0.7055 0.7050 0.6990 0.6673 0.7429 0.6859
-ARI 0.6687 0.6676 0.6565 0.6607 0.6170 0.6914 0.5792
-ACC 0.7865 0.8373 0.7834 0.8001 0.8032 0.8160 0.7478
-CV1 0.5616 0.4697 0.5441 0.5154 0.5394 0.5480 0.6385
-DCV -0.0666 0.0254 -0.0491 -0.0203 -0.0444 -0.0529 -0.1434
-Running time 0.2186 1.3271 0.6290 1.8517 139.5876 3.4712 0.8499
-Imbalanced Iris NMI 0.0001 0.0247 0.0061 0.6414 0.6870 0.9101 0.9101
-ARI -0.0053 0.0049 -0.0023 0.7421 0.7865 0.9582 0.9582
-ACC 0.5250 0.5500 0.5250 0.9500 0.9583 0.9917 0.9917
-CV1 0.1179 0.0471 0.0236 1.0842 1.0607 0.9664 0.9664
-DCV 0.8250 0.8957 0.9192 -0.1414 -0.1179 -0.0236 -0.0236
-Running time 0.0057 0.0121 0.0142 0.0257 0.1474 0.0146 0.0115
-Ecoli NMI 0.6370 0.5837 0.6087 0.5999 0.6026 0.6015 0.6530
-ARI 0.4953 0.4113 0.4704 0.4176 0.4468 0.6149 0.5202
-ACC 0.6310 0.5863 0.6190 0.6012 0.6012 0.6964 0.6458
-CV1 0.6871 0.4671 0.7517 0.4482 0.7770 1.0893 0.7001
-DCV 0.4733 0.6933 0.4087 0.7123 0.3835 0.0711 0.4604
-Running time 0.0255 0.2672 0.0760 0.1770 5.3315 0.1939 0.0886
-Imbalanced WBDC NMI 0.0966 0.0848 0.0828 0.6846 0.3893 0.3038 0.6907
-ARI 0.0337 0.0184 0.0158 0.7855 0.5156 0.3323 0.8308
-ACC 0.6260 0.5827 0.5745 0.9892 0.9539 0.9051 0.9892
-CV1 0.2644 0.1418 0.1188 1.3529 1.2073 1.0540 1.3069
-DCV 1.0578 1.1804 1.2034 -0.0307 0.1150 0.2683 0.0153
-Runing time 0.0141 0.0234 0.0094 0.0959 1.2488 0.0870 0.0535
-Htru2 NMI 0.4068 0.4101 0.4076 0.0547 0.3854 0.2611 0.5869
-ARI 0.6071 0.5798 0.6076 0.0217 0.5967 0.3504 0.7331
-ACC 0.9366 0.9253 0.9366 0.9097 0.9428 0.8473 0.9660
-CV1 1.0891 1.0125 1.0882 1.4107 1.1813 0.7805 1.2342
-DCV 0.3251 0.4017 0.3260 0.0035 0.2329 0.6337 0.1800
-Running time 0.0825 0.2913 0.1094 2.0558 68.3254 0.4956 0.1119